### v1

# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

Abstract
Recent advancements in Large Language Models (LLMs) have enhanced the efficacy of agent communication and social interactions. Despite these advancements, building LLM-based agents for reasoning in dynamic environments involving competition and collaboration remains challenging due to the limitations of informed graph-based search methods. We propose PLAYER*, a novel framework based on an anytime sampling-based planner, which utilises sensors and pruners to enable a purely question-driven searching framework for complex reasoning tasks. We also introduce a quantifiable evaluation method using multiple-choice questions and construct the WellPlay dataset with 1,482 QA pairs. Experiments demonstrate PLAYER*’s efficiency and performance enhancements compared to existing methods in complex, dynamic environments with quantifiable results.
**footnotetext: Equal contribution.
1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-as-Agent. Building upon this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite notable progress in enabling LLM-based agents to mimic human language, building agents for reasoning in dynamic environments, especially in scenarios considering competition and collaboration with other agents, remains a challenge, for example, the Murder Mystery Games (MMGs), a strategic game requires both cooperation and competition among 4-12 players through negotiation and tactical coordination (Figure 1). We argue that the primary reason is that most existing LLM-based agents involve the informed graph-based search method with a heuristics term to guide the reasoning progress. While this approach might be efficient in well-defined problems, where constructing a relation graph and producing a clear reasoning path as output is feasible, in complex problems requiring both cooperation and competition through natural language negotiation and tactical coordination, defining such a clear graph of potential actions becomes challenging. The main difficulty of applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios arises from the challenges in defining state spaces, action spaces, and rewards, where the reward associated with the final decision often represents the only clear certainty (FAIR et al., 2022; Shi et al., 2024).
Therefore, we propose a new framework based on an anytime sampling-based planner. Unlike informed graph-based search methods, such as A* (Dechter & Pearl, 1985), the anytime sampling-based planner, such as RRT* (Karaman & Frazzoli, 2011) or BIT* (Gammell et al., 2020a), tackles optimal motion planning problems in a more complex scenario. That is, by sampling possible states with detective sensors in real-time, the tree-style search path is constructed and pruned dynamically until the target is reached. Similarly, in MMGs, by proposing a set of sensors considering relations and emotions between players during interaction, along with a pruner that targets extracting highly suspicious murders, we construct a purely questioning-driven search framework. Our main contribution is that we proposed an approach which constructs a reasoning process without the need for pre-defined questions or prompt templates. Instead, it relies on a generator regulated by sensors and a pruner to efficiently generate high-quality questions for the agent’s reasoning task.
Moreover, the current evaluation is difficult to quantify. These methods rely heavily on manual evaluation for in-depth understanding but are considerably influenced by subjective interpretation (Xu et al., 2023a; Wang et al., 2023).To address this, we propose a quantifiable and reproducible evaluation criterion by devising a series of multiple-choice questions. This approach encompasses three types of questions: fact-based questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c), questions about the shared and individual objectives of each character set within the game, and questions probing the rationale behind the given responses.
In summary, we made the following contributions: (1) We propose PLAYER*, a framework aimed at efficiently optimising path planning in MMGs. (2) We conduct experiments to assess the efficiency and performance enhancements brought by PLAYER*, demonstrating its superiority over existing multi-agent methods. (3) We propose a quantifiable evaluation method using multiple-choice questions focusing on facts, character objectives, and reasoning, mitigating the subjectivity in current evaluation practices with the construction and annotation of the corresponding dataset, WellPlay, which includes 1,482 QA pairs.  [^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
Figure 1: The gameplay comprises three stages: introduction, discussion, and decision. During the introduction stage, players familiarise themselves with their characters’ background information and objectives, then briefly introduce themselves in character. Subsequently, they engage in interactive discussions to deduce the identity of the murderer.
2 PLAYER*
2.1 Problem Setting & Preliminary: Search by LLMs
In response to the complexities of social interactions such as MMGs, we have developed an innovative interactive framework tailored for such scenarios. As illustrated in Figure 1, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game featuring $N_{a}$ playable characters. $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$ represents the set of victims in the game. Each agent $a_{i}$ is assigned to a corresponding character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $o_{i}$. Here, $C_{i}$ is designed from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective. The suspicious state $s_{ik}$ is a vector representing the suspicions of killing victim $v_{k}$ from the perspective of agent $a_{i}$, which is an $(N_{a}-1)$-dim vector, where the entries can be discrete or continuous based on suspicion strength. The complete game rules and examples can be found in the appendix A.
Unlike transitional planning task in which the state space can be accessed directly, $s_{i}$ in our task is generated from a language model. Therefore, searching in state space is done by adjusting the prompting. We define the possible prompts as an action space, allowing the states to be obtained through an LLM-based search: $s_{i}=P(C^{\prime}_{i},T)$, where $P$ is the LLM, $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts, and $T$ is the prompt instruction. The preliminary planner contains two main components:
  • Explorer: By choosing a different prompt template $T$, which is usually from a pre-defined prompt template set, explore the corresponding $s_{i}$.
  • Predictor: a self-reflection estimator to evaluate $s_{i}$. collaborate with other heuristic strategies (like A* or MC tree search) to determine if it is necessary to explore another $s_{i}$ by switching $T$, or using this $s_{i}$ and a clue to continuously explore in the next iteration.
In general, the preliminary method only construct a prompting based planner for a give script. However, in the complex scenario like MMGs, the objective can be multiple, noted as $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$, including both personal and shared objectives. Subsequently, the agents requires to engage in a simulated discussion by asking questions and expressing opinions. To ensure fairness, the game imposes a limit of $m$ questions per agent per round, spanning a total of $n$ rounds. Following these rounds, agents are required to make decisions and cast their votes to identify the suspected murderer. Therefore, we need to construct a more complicated strategies considering the simulation among agents.
2.2 Construction of Agents
The strength of the preliminary method is that the tree structure and the potential reasoning path is clearly defined by the $T$. However, the lack of flexibility and the over simplified explorer may lead to a over consistency result and limit the reasoning capability of LLM. But the entirely free generation, especially allowing the communication between agents may lead to an uncontrollable results. Therefore, in our framework, we propose another two components to regularise the generation of $T$ before and after prompting respectively:
  • Sensors: inspired by sociology, which claims that the interpersonal relationships are often conceptualised within a multidimensional space encompassing factors(Latané, 1981; Zhao et al., 2021; Trope & Liberman, 2010), we propose employing a set of sensors associated with each task. For example, the emotion, motivation, and suspicion, are considered as the hints to generate prompt for state searching.
  • Pruner: After prompting, we design a pruner to reduce the state space to only few high suspicions in the next round of reasoning.
Besides, To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement a Memory Retrieval module to store and retrieve narrative scripts and dialog logs generated during gameplay. Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent. When an agent encounters a new event requiring action, we use the Faiss library (Douze et al., 2024) to retrieve relevant memories. This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length. It ensures consistent and coherent interactions among agents while assisting in their strategic planning.
Figure 2: PLAYER*’s planning strategy begins with an introduction round, during which each agent provides a brief self-introduction. Agents then initialise scenario understanding and begin questioning other agents. Subsequently, they start searching for further information by questioning other agents. Agents then adapt their level of suspicion and intended search area based on their received response, mirroring human gameplay strategies. This iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
2.3 PLAYER* Planning Strategy
To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximate the search domain through sampling, and planning the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 2, this framework is fundamentally composed of two key components:
Questioning Strategy in Searching based on Sensors
Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components. Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023b), we generate actions directed towards the targets based on the values detected by those sensors. The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors. For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Suspicion}]$: Emotion, which assesses one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible; Motivation, evaluating from one’s perspective whether the individual possesses a motive to be the perpetrator; and Suspicion, objectively determining whether they had the opportunity to commit the crime.
Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration, while the reward for such a cost is the information that can be obtained from the response. Under the assumption that increased inquiries to an agent corresponds to the heightened exploration within the world space related to that agent, the expected reward associated with questioning the same agent decreases in proportion to the number of inquiries posed. Therefore, it is necessary to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if this person might be highly suspicious. We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ about the agent to be questioned, as well as the number of inquiries that have already posed $|Q_{i}|$. Specifically, for agent $a_{i}$ to be questioned, PLAYER* would generate the question $q$ using the values of sensors $S_{i}$ in conjunction with the expected reward $P_{informative}$ and other related information $(C^{\prime}_{i},D_{i})$, that is $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, Suspicious States $s$, Max Round $n$
Output: Evaluation of Results
1 current_round = 0
2 while current_round < $n$ do
       // Action Generation
3       for i = 1 to $N_{a}$ do
4             for k = 1 to $N_{v}$ do
5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
6                   for $a_{j}$ in $suspect\_list_{ik}$ do
7                         question = Action_Generation($a_{i},a_{j},v_{i}$)
8                         answer = Reply($a_{i}$)
9
10
11
      // Constrain action space
12       for i = 1 to $N_{a}$ do
13             for k = 1 to $N_{v}$ do
14                   $s_{ik}$ = Refinement_Action_Space($a_{i},a_{j},v_{i}$)
15
16
17      current_round += 1
// Evaluation
18 for i = 1 to $N_{a}$ do
19       Result = Evaluate($a_{i}$)
20
Algorithm 1 Anytime Sampling-based Planning in Multi-agent Interaction
Action Space Refinement with a Pruner
Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents. Each agent first updates its comprehension of the interpersonal relationships, thereby adjusting sensor values: $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$. Subsequently, it also updated $P_{informative}$ based on the updated information $(S_{i},D_{i})$ and the changes in sensor values $\Delta S_{i}$.
In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals. This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Emulating the cognitive pattern of human beings, we introduce an Action Space Refinement strategy to refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets. We did not impose a hard restriction on the refinement process and allowed for flexible adjustment of the action space. For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked. This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
Algorithm 1 outlines a detailed procedure of the game process. While initially designed for MMGs, the framework’s adaptable nature allows it to be readily applied to various games by adjusting the game rules and task-specific sensors, thereby enhancing the gaming experience and offering a structured way for agents to handle complex social dynamics efficiently.
3 Experiments
3.1 Experimental Setup
Base Models
We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 and text-embedding-ada-002. To minimise randomness, we conducted the evaluation experiments 3 times and report the average and variance. The detailed introduction to the experiment settings is in Appendix C.1.
Baselines
For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games, such as Werewolf (Xu et al., 2023a), Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b), and ThinkThrice (Wu et al., 2024). Additionally, we assess the performance of agents that do not actively participate in the game but have direct access to either their own script (Single Script Access) or the scripts of all agents (Full Script Access). In werewolf, questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario. In O-CoT, agent interactions are driven by their set objectives. ThinkThrice has its agents crafting questions from retrieved memory and the current scenario. However, the efficiency of exploration with the questioning strategies is low, as evidenced by their performance, which falls significantly short of full script access. In contrast, our method leverages sensor data to identify the optimal domain to question and guide the generation of questions. The detailed comparison is in Appendix C.2.
WellPlay Dataset
We create an evaluation dataset, called the WellPlay Dataset, based on the existing dataset Conan (Zhao et al., 2024c), originated from background narratives created for MMGs, which provides annotated relationships between characters. We aim to concentrate on the most challenging and significant relationships that assess the agents’ comprehension of the scenario. Our approach is grounded in the original game’s objectives and informed by a trial round involving human players. This method enabled us to formulate the following types of evaluation questions in line with the human players:
1. Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
2. Reasoning. This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
3. Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information. This dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions. On average, each game features 5.67 agents and 1.75 victims. More detailed statistics are available in the appendix B.1.
Evaluation Metrics
We adopted the scoring system used in the game to evaluate agents’ performance: Awarding 10 points for achieving an objective, 5 points for correct reasoning, and 2 points for providing additional information based on the players’ understanding of the current situation. However, instead of using open-ended questions with human players, we assess each agent’s performance using our constructed dataset, WellPlay, to facilitate replication and comparison for future research. WellPlay comprises three types of questions: objective, reasoning, and relations, which are assigned 10, 5, and 2 points, respectively. We calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance. Additionally, we assess the accuracy of each question type. For the computation of the overall score for performance assessment, we have utilized weighted mean and weighted standard deviation methodologies. The detailed algorithm for these calculations is in Appendix B.2.
3.2 Results
Performance Evaluation Results
Table 1 presents the evaluation of agents’ performance across 12 unique games of varying complexity and settings. The number of evaluation questions (“#QA”) varies based on script complexity, with more complex scripts generating a larger volume of questions. The SSA (Single Script Access) measures agents’ performance with access to only their own script. This setup is designed to represent the starting point for searching. The FSA (Full Script Access) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint. Ideally, FSA should aim to achieve an accuracy of 1. However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model. Despite this, FSA still scores significantly higher than SSA. In addition, FSA also serves as an indicator of the complexity of the script.
We can observe that PLAYER* exhibits superior performance other baselinse across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents. Additionally, PLAYER* significantly outperforms others in objective questions, even surpassing FSA, which has access to all information. This showcases the effectiveness of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective. Despite some baselines exhibiting competitive performance in reasoning or relation questions, their significant drop in performance concerning objectives indicates a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals.
We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4. The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Death Wears White (9 players, 1 victim) Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
Reasoning $102$$454_{\pm.012}.$$503_{\pm.015}.$$350_{\pm.021}.$$395_{\pm.031}.$$392_{\pm.010}.$$408_{\pm.030}.$
Relations $72$$431_{\pm.036}.$$425_{\pm.031}.$$454_{\pm.035}.$$472_{\pm.037}.$$398_{\pm.032}.$$500_{\pm.048}.$
Overall $184$$420_{\pm.007}.$$480_{\pm.011}.$$367_{\pm.015}.$$393_{\pm.009}.$$377_{\pm.013}.$$407_{\pm.007}.$
Ghost Revenge (7 players, 3 victims) Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
Reasoning $152$$390_{\pm.020}.$$533_{\pm.030}.$$417_{\pm.023}.$$410_{\pm.019}.$$450_{\pm.056}.$$399_{\pm.040}.$
Relations $69$$246_{\pm.014}.$$411_{\pm.103}.$$348_{\pm.014}.$$309_{\pm.047}.$$420_{\pm.038}.$$386_{\pm.030}.$
Overall $240$$362_{\pm.016}.$$483_{\pm.020}.$$399_{\pm.009}.$$381_{\pm.049}.$$438_{\pm.030}.$$417_{\pm.032}.$
Danshui Villa (7 players, 2 victims) Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
Reasoning $128$$315_{\pm.020}.$$422_{\pm.016}.$$307_{\pm.044}.$$352_{\pm.008}.$$347_{\pm.005}.$$367_{\pm.008}.$
Relations $63$$323_{\pm.040}.$$476_{\pm.028}.$$344_{\pm.051}.$$381_{\pm.069}.$$333_{\pm.016}.$$370_{\pm.056}.$
Overall $203$$326_{\pm.014}.$$403_{\pm.028}.$$293_{\pm.040}.$$349_{\pm.012}.$$339_{\pm.013}.$$367_{\pm.011}.$
Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
Reasoning $61$$552_{\pm.025}.$$563_{\pm.025}.$$563_{\pm.010}.$$486_{\pm.025}.$$552_{\pm.019}.$$601_{\pm.019}.$
Relations $72$$500_{\pm.028}.$$546_{\pm.021}.$$523_{\pm.071}.$$583_{\pm.028}.$$556_{\pm.042}.$$616_{\pm.016}.$
Overall $145$$457_{\pm.017}.$$475_{\pm.033}.$$469_{\pm.018}.$$467_{\pm.004}.$$460_{\pm.030}.$$525_{\pm.029}.$
Cruise Incident (5 players, 1 victim) Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
Reasoning $24$$653_{\pm.064}.$$472_{\pm.064}.$$542_{\pm.042}.$$528_{\pm.024}.$$528_{\pm.105}.$$514_{\pm.024}.$
Relations $30$$267_{\pm.058}.$$411_{\pm.102}.$$456_{\pm.020}.$$456_{\pm.051}.$$422_{\pm.039}.$$434_{\pm.058}.$
Overall $58$$459_{\pm.012}.$$415_{\pm.079}.$$510_{\pm.037}.$$503_{\pm.017}.$$509_{\pm.107}.$$520_{\pm.025}.$
Sin (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
Reasoning $20$$483_{\pm.058}.$$567_{\pm.029}.$$767_{\pm.058}.$$600_{\pm.100}.$$717_{\pm.058}.$$600_{\pm.000}.$
Relations $21$$476_{\pm.082}.$$587_{\pm.099}.$$429_{\pm.000}.$$699_{\pm.055}.$$524_{\pm.048}.$$698_{\pm.027}.$
Overall $44$$397_{\pm.029}.$$531_{\pm.053}.$$570_{\pm.000}.$$539_{\pm.029}.$$564_{\pm.012}.$$577_{\pm.053}.$
Deadly Fountain (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
Reasoning $21$$413_{\pm.153}.$$667_{\pm.126}.$$460_{\pm.027}.$$476_{\pm.082}.$$539_{\pm.055}.$$587_{\pm.028}.$
Relations $12$$305_{\pm.048}.$$389_{\pm.048}.$$333_{\pm.084}.$$222_{\pm.096}.$$222_{\pm.048}.$$361_{\pm.127}.$
Overall $36$$319_{\pm.097}.$$604_{\pm.104}.$$354_{\pm.022}.$$369_{\pm.053}.$$390_{\pm.043}.$$463_{\pm.041}.$
Unbelievable Incident (5 players, 1 victim) Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
Reasoning $24$$389_{\pm.105}.$$528_{\pm.064}.$$431_{\pm.064}.$$361_{\pm.064}.$$264_{\pm.127}.$$583_{\pm.042}.$
Relations $15$$422_{\pm.102}.$$822_{\pm.102}.$$644_{\pm.077}.$$756_{\pm.102}.$$289_{\pm.154}.$$556_{\pm.077}.$
Overall $43$$330_{\pm.081}.$$481_{\pm.020}.$$444_{\pm.037}.$$382_{\pm.035}.$$230_{\pm.034}.$$509_{\pm.038}.$
Desperate Sunshine (4 players, 1 victim) Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
Reasoning $18$$556_{\pm.056}.$$778_{\pm.096}.$$648_{\pm.032}.$$704_{\pm.032}.$$759_{\pm.064}.$$741_{\pm.085}.$
Relations $36$$537_{\pm.064}.$$611_{\pm.048}.$$556_{\pm.028}.$$630_{\pm.043}.$$500_{\pm.056}.$$565_{\pm.070}.$
Overall $57$$514_{\pm.013}.$$680_{\pm.047}.$$599_{\pm.023}.$$618_{\pm.020}.$$647_{\pm.075}.$$680_{\pm.047}.$
Riverside Inn (4 players, 1 victim) Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
Reasoning $18$$519_{\pm.032}.$$667_{\pm.056}.$$463_{\pm.064}.$$500_{\pm.056}.$$685_{\pm.032}.$$667_{\pm.056}.$
Relations $18$$407_{\pm.085}.$$500_{\pm.096}.$$463_{\pm.085}.$$426_{\pm.064}.$$389_{\pm.000}.$$426_{\pm.032}.$
Overall $39$$479_{\pm.032}.$$671_{\pm.036}.$$459_{\pm.073}.$$387_{\pm.020}.$$614_{\pm.018}.$$590_{\pm.023}.$
Solitary Boat Firefly (6 players, 4 victims) Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
Reasoning $109$$318_{\pm.014}.$$569_{\pm.018}.$$370_{\pm.014}.$$425_{\pm.046}.$$401_{\pm.014}.$$379_{\pm.005}.$
Relations $69$$483_{\pm.017}.$$507_{\pm.014}.$$536_{\pm.014}.$$565_{\pm.038}.$$503_{\pm.044}.$$594_{\pm.087}.$
Overall $198$$332_{\pm.015}.$$544_{\pm.024}.$$354_{\pm.023}.$$430_{\pm.016}.$$375_{\pm.015}.$$444_{\pm.030}.$
Manna (6 players, 3 victims) Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
Reasoning $123$$458_{\pm.005}.$$393_{\pm.040}.$$569_{\pm.028}.$$499_{\pm.033}.$$526_{\pm.017}.$$539_{\pm.021}.$
Relations $88$$447_{\pm.058}.$$640_{\pm.036}.$$538_{\pm.043}.$$614_{\pm.030}.$$511_{\pm.034}.$$580_{\pm.046}.$
Overall $235$$408_{\pm.017}.$$434_{\pm.002}.$$525_{\pm.017}.$$505_{\pm.004}.$$501_{\pm.025}.$$553_{\pm.045}.$
Overall Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
Reasoning $800$$417_{\pm.095}.$$508_{\pm.094}.$$437_{\pm.104}.$$430_{\pm.077}.$$449_{\pm.113}.$$465_{\pm.106}.$
Relations $565$$411_{\pm.102}.$$513_{\pm.110}.$$466_{\pm.090}.$$495_{\pm.133}.$$444_{\pm.089}.$$500_{\pm.109}.$
Overall $1482$$386_{\pm.057}.$$484_{\pm.072}.$$415_{\pm.074}.$$424_{\pm.067}.$$426_{\pm.087}.$$461_{\pm.086}.$
Table 1: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
Efficiency and Cost Comparison
As shown in Table 2, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table. The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages  [^2] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved. This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling. The costs are divided into Gameplay and Evaluation phases. During Gameplay, agents actively participate in the game, generating actions and responses. In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset. Notably, SSA and FSA methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner.
PLAYER* stands out for its cost-efficiency and performance. By employing the Action Space Refinement strategy, PLAYER* minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters. This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations. This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase. Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in SSA and FSA costs arising from their differing levels of script access.
Script #Tokens Stage SSA FSA Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $3.363$
Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $5.947$
Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $4.603$
Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $3.505$
Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.846$
Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.596$
Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $1.150$
Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.701$
Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.561$
Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $6.800$
Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $5.028$
Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
Table 2: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. #$\#Tokens$ represent the average length of each character’s script. Costs are reported for one complete gameplay and one evaluation process for each script.
3.3 Ablation Studies
Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions. This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of information gained from conversations tends to converge. These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases. For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
Figure 3: Comparison of agents’ behaviour across different numbers of rounds, where each agent can ask a specific number of questions (denoted as $m$).
4 Related Works
Multi-Agent Interaction
Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023). However, these methods often require extensive time and computational resources and lack linguistic communication capabilities. With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024). Exemplified by AlphaGo (Silver et al., 2017), which demonstrated exceptional performance against human competitors, self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance. This idea has also been adapted for evaluation through negotiation games (Davidson et al., 2024). Compared to classic methods (Wang & Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states. However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024). Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
Optimisation for Complicated Tasks
Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive. To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced. These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang & Li, 2023; Ismail et al., 2024). Adapting LLM-as-a-Judge prompting to evaluate performance (Yuan et al., 2024) or select agents (Liu et al., 2023c) has become a popular approach, but negative results on self-reflection were also investigated (Huang et al., 2024), leaving LLM’s role as a self-reflective agent as an unresolved question. We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020b; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
5 Conclusion
PLAYER* addresses limitations of LLM-based agents in complex reasoning using an anytime sampling-based planner with sensors and pruners. Our approach enables efficient, question-driven searching. We propose a quantifiable evaluation method, contribute the WellPlay dataset, and demonstrate PLAYER*’s superiority, advancing effective reasoning agents in complex environments.
6 Ethics Statement
Please note that MMGs and our WellPlay dataset, created to assess agents’ behaviors in MMGs, may include descriptions of violent events, actions, or characters. This content is included solely for academic, research, and narrative analysis purposes. It is not meant to glorify or trivialise violence in any form. We have informed annotators about the potential exposure to violent content. Users or researchers intending to use this dataset should be aware of this potential exposure and are advised to engage with the dataset in a professional and responsible manner. This dataset is unsuitable for minors and those sensitive to such content.
Acknowledgments
This work was supported in part by the UK Engineering and Physical Sciences Research Council through a New Horizons grant (EP/X019063/1) and a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2).
Appendix A MMGs Rules and Procedure
A.1 Detailed Rules
Rule 1:
The total number of players participating in the game depends on the script. There may be one or more players who are the murderer(s), while the rest are civilians.
Rule 2:
The goal of the game is for civilian players to collaborate and face a meticulously planned murder case together, collecting evidence and reasoning to identify the real murderer among the suspects, all the while ensuring they are not mistaken for the murderer; murderer players must concoct lies to hide their identity and avoid detection, while also achieving other objectives in the game.
Rule 3:
Throughout the game, only murderer players are allowed to lie. To conceal their identity, murderers may choose to frame others to absolve themselves of guilt; non-murderer players (civilians) must answer questions from other players and the host honestly and provide as much information as they know about the case to help uncover the truth.
Rule 4:
At the start of the game, each player receives their character script from the host, which contains information about their role and identity.
Rule 5:
Other players cannot see the content of each player’s character script, so players must and can only collect information about other players through interaction after the game starts.
Rule 6:
In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with the most votes is the murderer, the civilian players win. Otherwise, the murderer players win.
A.2 Procedure
Stage 1: Distribution of Character Scripts
The host distributes character scripts to each player. These scripts contain the player’s name, role (murderer or civilian), and a brief character backstory.
Stage 2: Self-Introduction Session
Players introduce their characters to the group, laying the groundwork for the game’s interactions.
Stage 3: Rounds of Open Questioning
The game progresses through three rounds of open questioning. Players take turns to ask and answer questions, aiming to gather information about others.
Stage 4: Voting
In this stage, players vote anonymously to determine their suspicious regarding the identity of the murderer. Each player has one vote.
Stage 5: Outcome Reveal
The game concludes with the announcement of the voting results, revealing whether the civilian players successfully identified the murderer or not.
A.3 Example: Solitary Boat Firefly Script
As an illustrative example, we examine the Solitary Boat Firefly murder mystery script, involving six players: [ “Tian Chou", “Zhou Lianyi", “Xi Yan", “Yu Sunian", Yannan", and “Zhou Chitong" ], along with four victims: [ “Zhou Mengdang", “Bao Liu", “Cui Shouheng", and “Wang Xi Rong" ]. We model these characters through a set of agents, denoted as $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, where $N_{a}=6$ corresponds to the number of players. In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
Taking a closer look at “Tian Chou", depicted as the murderer player $a_{1}$, responsible for the demise of two among the four victims, her characterization unfolds as follows:
  • Character Background ($C_{1}$): "Originating from the Sun lineage, you are endearingly called ’Tian’er’. Your birth year was the twelfth of Guangxu’s reign during the Qing Dynasty (1886), marking the beginning of a life filled with extraordinary episodes…"
  • Suspicion State ($s_{1}$): The script features four victims, each harboring suspicions towards the remaining five players. Consequently, $s_{1}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players. For example, a first row of $[1,0,0,1,0]$ signifies that victim “Cui Shouheng" suspects both “Zhou Lianyi" and “Yannan". So we can get $suspect\_list$ of victim “Cui Shouheng" is [“Zhou Lianyi" ,“Yannan"]
  • Individual Objectives ($o_{i}$): The personal objectives for Tian Chou include:
    1. 1.
      Concealing that you killed “Zhou Mengdang".
    2. 2.
      Concealing that you killed “Bao Liu".
    3. 3.
      Find out the truth about “Taitai’s death".
    4. 4.
      Conceal your relationship with “Zhou Chitong".
    5. 5.
During the game, the dialogue history is recorded in $D$, and as the game progresses, inquiries and responses are conducted in accordance with Algorithm 1, leading to updates in $s_{i}$.
Appendix B The WellPlay Dataset
B.1 Details of Dataset
Script Agents Victims #token(CN) #token(EN) Question
avg overall avg overall Objective Reasoning Relations overall
Death Wears White 9 1 3190.67 28716 1742.33 15681 10 102 72 184
Ghost Revenge 7 3 5487.86 38415 3960.43 27723 19 152 69 240
Danshui Villa 7 2 5111.29 35779 3338.57 23370 12 128 63 203
Unfinished Love 7 2 2501.00 17507 1651.71 11562 12 61 72 145
Cruise Incident 5 1 1262.60 6313 808.00 4040 4 24 30 58
Sin 4 1 2121.25 8485 1378.00 5512 3 20 21 44
Deadly Fountain 4 1 1852.50 7410 1193.75 4775 3 21 12 36
Unbelievable Incident 5 1 3182.40 15912 2012.40 10062 4 24 15 43
Desperate Sunshine 4 1 3370.25 13481 2218.50 8874 3 18 36 57
Riverside Inn 4 1 1909.50 7638 1257.00 5028 3 18 18 39
Solitary Boat Firefly 6 4 8893.67 53362 6874.00 41244 20 109 69 198
Manna 6 3 9028.17 54169 6492.33 38954 24 123 88 235
Avg 5.67 1.75 3992.60 23932.25 2743.92 16402.08 9.75 66.67 47.08 125.50
Sum 68 21 47911.16 287187 32927.02 196825 117 800 565 1482
Table A1: Dataset Statistics. Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively. Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types.
In this work, we leverage the Conan dataset, originally constructed by Zhao et al. (2024c), which comprises scripts from MMGs, including detailed annotations of character relationships. Our bilingual dataset, encompassing both Chinese and English versions, is derived from this foundational work. Our dataset consists of two main components: scripts and evaluation questions.
Each script features distinct narratives for individual characters. We have rephrased the original script from the Conan dataset into two parts:
  1. 1.
    Background Script. For each character, there is corresponding background script includes all the information from their perspective. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
    You are the head nurse of the emergency ward.You climbed to this position for your hard work and were proud. You are a very professional person and highly appreciated by colleagues. This job is perfect for you, except for a problem - you have not made enough money. Your salary is actually not enough for you to live a decent life, far from paying enough to take care of Mother’s overhead …
  2. 2.
    Personal Objectives. For each character, there are corresponding objectives that guide their actions. For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
    1. Ensure that no one will discover your illegal organ trafficking activities;
    2. Ensure that the kidnappers leave without being injured or killing anyone - you want to ensure that the police do not investigate deeply enough to discover your organ trafficking. The further away you are from the police, the safer you feel;
    3. …
Type Aspects Examples (The correct answer has been highlighted in bold.)
A
(score 10)
Who
Who killed Hans Li Morette?
A. Gale Li Morette
B.Nurse head [Sylvia Costa]
C. Drake Li Morette
D. Frank Bijeli
B (score 5) How
How did Hans Li Morette die?
A. Shot to death
B. Beaten to death
C. Poisoned to death by poison
D.Drowned by water
Why
What was the motive behind the killer killing Hans Li Morette?
A. Love killing
B. Vendetta
C. Interest
D. Accidental killing
Relationship
What is the relationship between Murderer and Victim Hans Li Morette?
A. Enemies
B. Colleague
C. Friend
D. Wife
Where
Where was Hans Li Morette killed?
A. Emergency room
B. Johnson’s House
C. Laboratory
D. Dressing room
When
When was Hans Li Morette killed?
A. This afternoon from 5:00 to 5:30
B. This afternoon from 6:30 to 7:00
C. Tonight from 7:00 to 7:30
D. This morning from 6:30 to 7:00
Suspect
Please select the two people you most suspect of killing Hans Li Morette
A. Gale Li Morette
B. Nurse head [Sylvia Costa]
C. Drake Li Morette
D. Frank Bijeli
C (score 2)
Three
relationships
What is the non-existent relationship between Hans Li Morette and Andrew Paloski?
A. Andrew Paloski is colleague of Hans Li Morette,
B. Andrew Paloski is mentor of Hans Li Morette,
C. Andrew Paloski is jealous of Hans Li Morette,
D. Hans Li Morette is future daughter in law of Andrew Paloski
Two
relationships
What is the relationship between Father Tom and Tony?
A. Tony is manipulated by x and deceived by x of Father Tom
B. Father Tom is authority over x and student of Tony
C. Father Tom is student and ex-girlfriend of Tony
D. Father Tom is ex-girlfriend and admired by x of Tony
One
relationships
What is the relationship between Father Tom and Drake Li Morette?
A. Drake Li Morette is doctor of Father Tom
B. Father Tom is helped by Drake Li Morette
C. Father Tom is step-brother of Drake Li Morette
D. Father Tom is hate of Drake Li Morette
Table A2: Examples of Each Type of Our Evaluation Questions
Index Question
1 What was your timeline on the day of the incident?
2 How would you describe your relationship with the victim?
3 When was the last time you saw the victim?
4 Do you know if the victim had any enemies or conflicts with anyone?
5 What details or anomalies did you notice at the scene of the crime?
6 Did the victim mention anything to you or others that made them worried or fearful recently?
7 Did you notice any unusual people or behaviors on the day of the incident?
8 How much do you know about the victim’s secrets or personal life?
9 Were there any items or remains found at the crime scene that could be related to the crime?
10 Do you have any personal opinions or theories about the case?
Table A3: Predefined questions for Werewolf method.
The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts. These questions are segmented into three categories, with each category of questions and examples shown in Table A2.
  1. 1.
    Objective: This category includes common objectives shared among agents, such as finding the perpetrator(s).
  2. 2.
    Reasoning: Comprising six types of questions, this section tests the models’ reasoning capabilities across various aspects:
    • How: How the murder was committed.
    • Why: The motive behind the murder.
    • Relationship: The relationship between the murderer and the victim.
    • Where: The location where the murder took place.
    • When: The time at which the murder occurred.
    • Suspect: Identify the two most suspicious individuals. Analogous to human reasoning, it may be challenging to definitively pinpoint the suspect, yet it is possible to determine those who seem most suspect.
  3. 3.
    Relations: This segment examines the model’s capacity to comprehend complex narratives by querying about the relationships between characters. Utilizing the relationship annotations from the Conan dataset(Zhao et al., 2024c), we approach this analysis through various questioning strategies based on the number of relationships between two characters:
    • For three relationships, we employ the elimination method, asking the model to identify the incorrect relation among the given options.
    • For two relationships, we connect the correct two relationships with “and" and introduce a distractor from other relationship categories.
    • For single relationships, we list the correct option alongside distractors randomly selected from other categories.
For each individual mentioned in the purpose section, we extract three significant relationships associated with them. In this context, “significant" is defined as the characters who have the most complex connections with the individual, that is, the top three characters who share the greatest variety of relationships with them. For instance, in the provided example, Andrew Paloski is Hans Li Morette’s colleague, mentor, and is also jealous of him.
For various types of questions, we assign different weights based on the original scoring system of the script. Specifically, Type A questions are valued at 10 points, Type B questions at 5 points, and Type C questions at 2 points.
Detailed script statistics and evaluation question metrics are presented in TableA1.
B.2 Overall Performance Computing
In calculating the overall score for performance, we have employed both the weighted mean and the weighted standard deviation. The weighted mean is computed by considering the count of questions for a specific category across various scripts as the weight. For the overall score, the total possible score for each script serves as the weight. This method allows us to adjust the influence of each category and script based on its significance and scale, thus providing a more nuanced and accurate reflection of performance.
The weighted mean is calculated as:
$\bar{x}_{w}=\frac{\sum_{i=1}^{n}(w_{i}\cdot x_{i})}{\sum_{i=1}^{n}w_{i}}$
The weighted standard deviation, which measures the spread of the scores, is calculated using the weighted variance:
$s_{w}^{2}=\frac{\sum_{i=1}^{n}w_{i}\cdot(x_{i}-\bar{x}_{w})^{2}}{\sum_{i=1}^{n}w_{i}}$
And the weighted standard deviation is the square root of the weighted variance:
$s_{w}=\sqrt{s_{w}^{2}}$
Appendix C Implementation
C.1 Implementation Details
RAG (Retrieval-Augmented Generation)
For retrieval enhancement, we utilised the FAISS [^3] https://github.com/facebookresearch/faiss library to build a vector database, creating FAISS indices using the L2 distance. Embeddings were obtained via the Azure API’s text-embedding-ada-002 service. Scripts were stored in segments, with each segment having a maximum length of 50 tokens. For dialog records, a question-and-answer pair was stored as a single segment. During retrieval, the maximum script length and dialog length inserted into the prompt were both set to 4000 tokens. For evaluation, these maximum lengths were increased to 5000 tokens.
Experiment
Following the results of our ablation studies, the gameplay phase was structured to ask one question per round over three rounds. After the game concluded, the evaluation phase consisted of three separate evaluations, with the final results being the average of these evaluations.
C.2 Comparison Models
In this section, we provide a detailed overview of the methodologies compared in our study. Given the game’s rules and questioning sequence in our comparisons, we segment the discussion into two phases: the questioning phase and the answering phase.
Considering the strategic core of the game, a direct confession from the murderer would compromise its competitive nature. Hence, in all the methodologies we explore, we maintain that the murderer’s responses follow a predefined template, detailed in C.6.2. This method preserves the game’s integrity and level of challenge.
  1. 1.
    Werewolf(Xu et al., 2023a)
    Question Generation:
    1. a.
      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table A3.
    2. b.
      Formulating questions by the selected predefined questions and the ongoing dialogue and script.
    Answering Questions:
    1. a.
      Responding based on the current script and dialog history.
    2. b.
      Reflecting on the initial response in light of the dialog history.
    3. c.
      Generating the final answer after reflection.
  2. 2.
    Objective-Guided Chain of Thought (O-CoT)(Park et al., 2023; Zhao et al., 2024b)
    Question Generation: Entails two critical steps:
    1. a.
      Sequentially reflecting on whether current objectives have been met, with considerations spanning multiple goals such as identifying the murderer, uncovering hidden relationships, or concealing facts. For example, Who is the murderer?
    2. b.
      Crafting questions based on these reflections and the current narrative and dialogue.
    Answering Questions:
    1. a.
      Answers are formulated leveraging the narrative and dialog history.
  3. 3.
    ThinkThrice(Wu et al., 2024)
    Question Generation:
    1. a.
      Questions are generated based on the script and dialog history.
    Answering Questions:
    1. a.
      Extracting timelines relevant to the victim.
    2. b.
      Evaluating each timeline’s relevance to answering the question.
    3. c.
      Responding based on the dialog history, character relationships, and the relative script.
    4. d.
      Ensuring timelines that aid in answering the question are included in the response. If these are initially missed, the model will later augment and clarify the answer with the required timelines.
  4. 4.
    The SSA (Single Script Access)
    The SSA assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
  5. 5.
    The FSA (Full Script Access)
    The FSA evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
C.3 Performance Evaluation Results for English Dataset
Due to budgetary constraints, we only evaluated four scripts from the English dataset, with the performance results reported in Table A4. For the same scripts, we found that the experiments conducted on the English corpus corroborate the results obtained from the Chinese corpus. Additionally, we observed that the performance of agents based on the English corpus surpassed those based on the Chinese corpus, showing the differential inferencing abilities of LLMs across languages. This discrepancy could be attributed to language biases inherent in the training data utilised for these models.
Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Death Wears White (9 players, 1 victim) Objective $10$$200_{\pm.173}.$$900_{\pm.100}.$$300_{\pm.100}.$$267_{\pm.058}.$$300_{\pm.000}.$$267_{\pm.058}.$
Reasoning $102$$350_{\pm.044}.$$520_{\pm.026}.$$356_{\pm.037}.$$363_{\pm.049}.$$399_{\pm.006}.$$441_{\pm.030}.$
Relations $72$$547_{\pm.008}.$$445_{\pm.026}.$$495_{\pm.016}.$$495_{\pm.057}.$$398_{\pm.032}.$$491_{\pm.021}.$
Overall $184$$367_{\pm.040}.$$549_{\pm.012}.$$375_{\pm.033}.$$375_{\pm.026}.$$385_{\pm.010}.$$427_{\pm.029}.$
Ghost Revenge (7 players, 3 victims) Objective $19$$403_{\pm.061}.$$561_{\pm.132}.$$439_{\pm.109}.$$211_{\pm.106}.$$211_{\pm.106}.$$526_{\pm.106}.$
Reasoning $152$$456_{\pm.004}.$$507_{\pm.007}.$$439_{\pm.031}.$$539_{\pm.023}.$$441_{\pm.017}.$$423_{\pm.031}.$
Relations $69$$309_{\pm.009}.$$261_{\pm.025}.$$328_{\pm.059}.$$304_{\pm.015}.$$280_{\pm.017}.$$353_{\pm.059}.$
Overall $240$$428_{\pm.010}.$$485_{\pm.024}.$$425_{\pm.002}.$$452_{\pm.023}.$$380_{\pm.010}.$$432_{\pm.013}.$
Danshui Villa (7 players, 2 victims) Objective $12$$167_{\pm.083}.$$278_{\pm.048}.$$472_{\pm.048}.$$333_{\pm.000}.$$305_{\pm.048}.$$389_{\pm.096}.$
Reasoning $128$$325_{\pm.031}.$$427_{\pm.009}.$$401_{\pm.020}.$$373_{\pm.020}.$$344_{\pm.014}.$$357_{\pm.033}.$
Relations $63$$328_{\pm.056}.$$460_{\pm.028}.$$391_{\pm.066}.$$312_{\pm.106}.$$376_{\pm.046}.$$407_{\pm.075}.$
Overall $203$$292_{\pm.012}.$$412_{\pm.017}.$$409_{\pm.009}.$$359_{\pm.010}.$$343_{\pm.010}.$$369_{\pm.033}.$
Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$195_{\pm.048}.$$389_{\pm.048}.$$417_{\pm.000}.$$361_{\pm.048}.$$528_{\pm.127}.$
Reasoning $61$$536_{\pm.018}.$$557_{\pm.028}.$$650_{\pm.038}.$$612_{\pm.009}.$$634_{\pm.038}.$$656_{\pm.000}.$
Relations $72$$491_{\pm.016}.$$551_{\pm.016}.$$509_{\pm.032}.$$481_{\pm.021}.$$592_{\pm.008}.$$500_{\pm.037}.$
Overall $145$$446_{\pm.006}.$$479_{\pm.029}.$$560_{\pm.011}.$$538_{\pm.010}.$$566_{\pm.009}.$$589_{\pm.018}.$
Table A4: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
C.4 Experiments with Open-Source LLMs
Models Llama2 70b Llama2 13b Llama2 7b gemma 7b
Overall Score 0.312 0.281 0.267 0.273
Table A5: Compare the performance of PLAYER* method with different open-source LLMS.
In addition to experimenting with GPT-3.5-turbo-16k 0613, we also explored Llama2 (70b, 13b, 7b) and Gemma 7b. These models were tested using the scenario “Solitary Boat Firefly", and the overall results are reported in Table A5. The findings indicate that all four models scored significantly lower than GPT-3.5, presumably due to the limitations imposed by a 4k context window. This limitation likely hindered the models’ ability to encapsulate sufficient relevant information within such a constrained window. After thorough testing, we observed that despite being evaluated on a Chinese dataset, Llama2 70b primarily resorted to English conversations due to its limited proficiency in Chinese. It responded in English even to Chinese prompts. Other models struggled even more with executing the prompts as required. This limitation greatly hindered their performance in complex gaming situations, such as MMGs. Therefore, we decided to exclusively use GPT-3.5 for future experiments, given its ability to navigate these intricate scenarios.
C.5 Sensors
This section provides a detailed explanation of the sensors employed in the 2.3, which are essential for both the Search by Questioning and Action Space Refinement components.
  • Emotion Sensor: Assesses emotional inclination towards a character. Used in both Search by Questioning and Action Space Refinement. It categorises emotional inclination into “Positive", “Neutral", or “Negative".
  • Motivation Sensor: Evaluates the character’s relationship with the victim and the presence of a motive for the crime. It is active in both phases, with choices being “Yes" or “No".
  • Suspicion Sensor: Determines if a character is a suspect by analysing their opportunity to commit the crime. It applies to both stages, with responses “Yes" or “No".
  • Information Value Sensor: Exclusive to Action Space Refinement, it estimates the probability of obtaining valuable information from further questioning. The choices are “High", “Medium", or “Low".
"name": "emotion",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
"sensor_prompt": "What is your emotional inclination towards the character mentioned above?",
"choices": ["Positive", "Natural", "Negative"],
"name": "motivation",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
[…]
"choices": ["Yes", "No"],
"name": "suspicion",
"for_Search_by_Questioning": True,
"for_Action_Space_Refinement": True,
[…]
"choices": ["Yes", "No"],
"name": "information value",
"for_Search_by_Questioning": False,
"for_Action_Space_Refinement": True,
"sensor_prompt": "What do you think is the probability of obtaining valuable information by continuing to question the character mentioned above?",
"choices": ["High", "Medium", "Low"],
C.6 Prompt
C.6.1 System Prompt
System Prompt
designed to introduce the gameplay of the MMG, along with providing essential information about the agents involved. The prompts dynamically adapt to include {character_name}, representing the agent’s character in the game, and {character_name_list}, listing the characters played by other participants.
For Civilian Players:
You are playing a game called "Murder Mystery" with other players, which is based on textual interaction. Here are the game rules:
Rule 1: The total number of players participating in the game depends on the script. There may be one or more players who are the murderer(s), while the rest are civilians.
Rule 2: The goal of the game is for civilian players to collaborate and face a meticulously planned murder case together, collecting evidence and reasoning to identify the real murderer among the suspects, all the while ensuring they are not mistaken for the murderer; murderer players must concoct lies to hide their identity and avoid detection, while also achieving other objectives in the game.
Rule 3 Throughout the game, only murderer players are allowed to lie. To conceal their identity, murderers may choose to frame others to absolve themselves of guilt; non-murderer players (civilians) must answer questions from other players and the host honestly and provide as much information as they know about the case to help uncover the truth.
Rule 4: At the start of the game, each player receives their character script from the host, which contains information about their role and identity.
Rule 5: Other players cannot see the content of each players character script, so players must and can only collect information about other players through interaction after the game starts.
Rule 6: In the voting phase, each player needs to cast their vote for who they think is the murderer in each case. If the player with the most votes is the murderer, the civilian players win. Otherwise, the murderer players win.
Gameplay:
The game has one or more acts. At the beginning of the game, players introduce themselves according to the script, and in each act, you will receive more plot information. In each act, you can ask questions, share your observations, or make deductions to help solve the murder case.
The goal is to identify the true murderer and explain their motive. If you are the true murderer, you must hide your identity and avoid detection.
Now, you are playing the role of {character_name}, and the other players are {character_name_list}.
You are not the murderer. Please collaborate with the other civilian players to achieve your personal objective while finding the true culprit!
For Killer Players, we replace the last paragraph with:
You are the murderer who kills {victims}, and you havent killed anyone else. Please hide the fact that you committed the murder by fabricating lies and other information, and accomplish your personal objective!
C.6.2 Gameplay Prompt
Self-Introduction Prompt
designed to facilitate introductions based on the agent’s script and objectives. Within this prompt, {current_script} represents the character’s script, and {goal} represents the objectives of the character within that script.
For Civilian Players:
Your Script is {current_script}.
Your goal is {goal}.
You are not a murderer, so tell more details to help find the murderer.
If you have something to hide, then be sure not to divulge the relevant information! Dont reveal your goals.
Please introduce yourself.
For Killer Players:
Your Script is {current_script}.
Your goal is {goal}.
If you have something to hide, then be sure not to divulge the relevant information!
Please introduce yourself.
You are a murderer, so You can lie to cover yourself!
Sensor Prompt
As outlined in sections 2.3 and C.5, the Sensor Prompt is crafted to collect a wide array of crucial information. Within this framework, the placeholders {victim} and {character} are employed to reference the victim’s name and the agent being questioned, respectively. Meanwhile, {current_script} and {dialog_history} retrieve the script and dialog logs relevant to both the deceased and the agent under scrutiny, utilising the RAG technique. The constructs {sensor_Prompt} and {choices} have been formulated in C.5 within the Sensor Prompt segment.
For Civilian Players:
{victim} was murdered, you are not the murderer, you need to try to find the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
Be warned, if its a murderers word it might deceive you.
Based on the information above, {sensor}
Please answer {choices} and explain your reasoning in one or two sentences.
For Killer Players:
{victim} was murdered, you are the murderer, but you need to hide yourself, and pretend youre not the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
Based on the information above, {sensor}
Please answer {choices} and explain your reasoning in one or two sentences.
Search by Questioning Prompt
As outlined in sections 2.3, the Search by Questioning Prompt is developed based on data acquired from sensors. It includes variables like {victim}, {character}, {current_script}, {dialog_history}. Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data. The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 3.3 chapter.
For Civilian Players:
{victim} was murdered, you are not the murderer, you need to try to find the murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
{summary}
You can ask {character} {question_number} questions. What would you ask? Please include the victims name in your question when asking,
Since the murderer will lie, you can ask questions based on the loopholes and contradictions in what they have previously said.
Please respond in the JSON format without any additional comments.
For example,
Question1’: Your question’,
Question2’: Your question
For Killer Players:
{victim} was murdered, you are the murderer, But you need to hide yourself, pretend youre not a murderer, and ask questions of other people pretending you suspect the other person is a murderer.
Your Script is about {character} is {current_script}.
The dialog history about {character} is {dialog_history}.
{summary}
You can ask {character} {question_number} questions. What would you ask? Please include the victims name in your question when asking.
Please respond in the JSON format without any additional comments.
For example,
Question1’: Your question’,
Question2’: Your question
Action Space Refinement Prompt
As outlined in sections 2.3, the Action Space Refinement prompt serves as a strategic tool for narrowing down the suspect list, effectively reducing the search domain to augment the efficiency and performance of the algorithm. Within this setup, the term {victim} refers to the individual who has been harmed, {summary} synthesised from the sensor-collected data, and {character_suspect} identifies the roster of individuals under suspicion. Initially, this roster includes all participating agents, excluding the itself.
{victim} was murdered.
You think {character_suspect} are suspected of killing {victim}, and your reasons for suspecting them are respectively:
{summary}
Please select several people you think are the most suspicious. You can choose one or more, Please try to reduce the number of suspects.
Please respond in the JSON format without any additional comments.
For example,
suspicion’: ["character_name1", "character_name2"]
Question Reply Prompt
The Question Reply Prompt is designed for responding to inquiries, where {character} denotes the name of the character asking you a question, and {question} is the question itself. {current_script} refers to the script extracted using RAG that pertains to both the deceased and the agent being questioned; {dialog_history} is the dialog log related to the deceased and the interrogated agent, also extracted using RAG.
For Civilian Players:
{character} ask you a question: {question}
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
What you need to pay attention to is {goal}.
Be warned, in the dialog history, if its a murderers word it might deceive you.
Please answer the question: "{question}" based on the information above.
You are not the murderer, and you need to work hard to find the murderer. Therefore, provide as much information as possible, such as clues related to the timeline, emotional information, etc.
Please answer the questions from a first-person perspective, rather than saying what someone else said.
For Killer Players:
{character} ask you a question : {question}
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
What you need to pay attention to is {goal}.
Be warned, in the dialog history, if its a murderers word it might deceive you.
Please answer the questions: "{question}" based on the information above.
You are the murderer. Please hide the fact that you killed {victim}. You can fabricate lies.
Please answer the question from a first-person perspective, rather than saying what someone else said.
C.6.3 Evaluation Prompt
Single-Choice Question Template for the Evaluation Stage Prompt
This template is utilised during the evaluation stage for answering single-choice questions. It incorporates {current_script}, the script related to the question extracted using RAG, and {dialog_history}, the dialog log relevant to the question, also extracted via RAG. {question} is the inquiry presented, and {choices} are the available options for that question.
Please answer the questions based on the information in your script and the content of the dialog
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
The question is {question}, the choices is {choices}
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number, e.g., a’)
You must choose one from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a"
Multiple-Choice Question Template for the Evaluation Stage Prompt
Similar to the single-choice template, this format is designed for responding to multiple-choice questions during the evaluation stage.
Please answer the questions based on the information in your script and the content of the dialog
Your Script relative to the question is {current_script}.
The dialog history relative to the question is {dialog_history}.
The question is {question}, the choices is {choices}
That is a multiple-choice question.
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must make a choice from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"
Single-Choice Question Template for FSA Prompt
This template is specifically designed for Full Script Access (FSA), distinguishing it from conventional templates by granting access to the scripts of all players. Unlike standard procedures, it leverages the RAG technique to extract pertinent information from each player’s script for comprehensive analysis.
Please answer the questions based on the information in each characters script:
{current_script}
The question is {question}, the choices is {choices}
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must choose one from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"
Multiple-Choice Question Template for FSA Prompt
Similar to the single-choice version, this template is tailored for answering multiple-choice questions under FSA conditions.
Please answer the questions based on the information in each characters script:
{current_script}
The question is {question}, the choices is {choices}
That is a multiple-choice question.
Lets think about this problem step by step, please provide your reasoning and your choice (only the option number)
You must make a choice from the options.
Please respond in the JSON format without any additional comments.
For example,
"reason": "Your reason",
"answer": "a,b"

### v2

@@
 Abstract
-Recent advancements in Large Language Models (LLMs) have enhanced the efficacy of agent communication and social interactions.
-Despite these advancements, building LLM-based agents for reasoning in dynamic environments involving competition and collaboration remains challenging due to the limitations of informed graph-based search methods.
-We propose PLAYER*, a novel framework based on an anytime sampling-based planner, which utilises sensors and pruners to enable a purely question-driven searching framework for complex reasoning tasks.
-We also introduce a quantifiable evaluation method using multiple-choice questions and construct the WellPlay dataset with 1,482 QA pairs.
-Experiments demonstrate PLAYER*’s efficiency and performance enhancements compared to existing methods in complex, dynamic environments with quantifiable results.
+We propose PLAYER*, a novel framework that addresses the limitations of existing agent-based approaches built on Large Language Models (LLMs) in handling complex questions and understanding interpersonal relationships in dynamic environments.
+PLAYER* enhances path planning in Murder Mystery Games (MMGs) using an anytime sampling-based planner and a questioning-driven search framework.
+By equipping agents with a set of sensors, PLAYER* eliminates the need for pre-defined questions and enables agents to navigate complex social interactions.
+We additionally make a contribution by introducing a quantifiable evaluation method using multiple-choice questions and present WellPlay, a dataset containing 1,482 question-answer pairs.
+Experimental results demonstrate PLAYER*’s superiority over existing multi-agent methods, enhancing the generalisability and adaptability of agents in MMGs and paving the way for more effective multi-agent interactions.
+PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games
+Qinglin Zhu1∗, Runcong Zhao1∗, Jinhua Du2, Lin Gui1, Yulan He1,3 1King’s College London, 2Huawei London Research Centre, 3The Alan Turing Institute {qinglin.1.zhu, runcong.zhao}@kcl.ac.uk{jinhua.du}@huawei.com {lin.1.gui,yulan.he}@kcl.ac.uk
 **footnotetext: Equal contribution.
 1 Introduction
-Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-as-Agent.
-Building upon this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023).
-Despite notable progress in enabling LLM-based agents to mimic human language, building agents for reasoning in dynamic environments, especially in scenarios considering competition and collaboration with other agents, remains a challenge, for example, the Murder Mystery Games (MMGs), a strategic game requires both cooperation and competition among 4-12 players through negotiation and tactical coordination (Figure 1).
-We argue that the primary reason is that most existing LLM-based agents involve the informed graph-based search method with a heuristics term to guide the reasoning progress.
-While this approach might be efficient in well-defined problems, where constructing a relation graph and producing a clear reasoning path as output is feasible, in complex problems requiring both cooperation and competition through natural language negotiation and tactical coordination, defining such a clear graph of potential actions becomes challenging.
-The main difficulty of applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios arises from the challenges in defining state spaces, action spaces, and rewards, where the reward associated with the final decision often represents the only clear certainty (FAIR et al., 2022; Shi et al., 2024).
-Therefore, we propose a new framework based on an anytime sampling-based planner.
-Unlike informed graph-based search methods, such as A* (Dechter & Pearl, 1985), the anytime sampling-based planner, such as RRT* (Karaman & Frazzoli, 2011) or BIT* (Gammell et al., 2020a), tackles optimal motion planning problems in a more complex scenario.
-That is, by sampling possible states with detective sensors in real-time, the tree-style search path is constructed and pruned dynamically until the target is reached.
-Similarly, in MMGs, by proposing a set of sensors considering relations and emotions between players during interaction, along with a pruner that targets extracting highly suspicious murders, we construct a purely questioning-driven search framework.
-Our main contribution is that we proposed an approach which constructs a reasoning process without the need for pre-defined questions or prompt templates.
-Instead, it relies on a generator regulated by sensors and a pruner to efficiently generate high-quality questions for the agent’s reasoning task.
-Moreover, the current evaluation is difficult to quantify.
-These methods rely heavily on manual evaluation for in-depth understanding but are considerably influenced by subjective interpretation (Xu et al., 2023a; Wang et al., 2023).To address this, we propose a quantifiable and reproducible evaluation criterion by devising a series of multiple-choice questions.
-This approach encompasses three types of questions: fact-based questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c), questions about the shared and individual objectives of each character set within the game, and questions probing the rationale behind the given responses.
-In summary, we made the following contributions: (1) We propose PLAYER*, a framework aimed at efficiently optimising path planning in MMGs.
-(2) We conduct experiments to assess the efficiency and performance enhancements brought by PLAYER*, demonstrating its superiority over existing multi-agent methods.
-(3) We propose a quantifiable evaluation method using multiple-choice questions focusing on facts, character objectives, and reasoning, mitigating the subjectivity in current evaluation practices with the construction and annotation of the corresponding dataset, WellPlay, which includes 1,482 QA pairs.
+Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023).
+Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023), deception (Wang et al., 2023), and the spread of information (Park et al., 2023).
+Despite these advances, building agents that intentionally interact with humans in dynamic environments remains challenging due to several limitations.
+First, current state-of-the-art approaches struggle with complex questions that lack straightforward answers and have difficulty in understanding interpersonal relationships.
+As a result, in addition to the game rules, agents often require additional human-written guides tailored to specific scenarios, relying on human context understanding (Xu et al., 2023; Wang et al., 2023).
+This limitation highlights the need for significant manual efforts and reduces the generalisability of these approaches across different domains and situations.
+Second, applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios is challenging because of the difficulties in defining state spaces, action spaces, and rewards.
+The reward associated with the final decision (e.g., identifying the murderer) often represents the only clear certainty (win or lose) (FAIR et al., 2022; Shi et al., 2024).
+Third, current evaluation methods are difficult to quantify.
+These methods rely heavily on manual evaluation to gain a deep understanding, but they are considerably influenced by subjectivity (Xu et al., 2023; Wang et al., 2023).
+Figure 1: The murder mystery game unfolds as players take on roles, question suspects based on their character’s objectives, and use deduction to identify the killer, while achieving their ideal role-specific ending.
+The complete game rules can be found in the appendix A.
+To address the limitation of lacking generalisability, we propose a novel framework that efficiently optimises path planning in Murder Mystery Games (MMGs), a strategic game involving both cooperation and competition among 4-12 players through natural language negotiation and tactical coordination (as illustrated in Figure 1).
+Our approach eliminates the need for pre-defined questions or instructions, allowing agents to play various scripts without extra human-written guides, just like human players.
+This framework enables agents to navigate complex social interactions and make decisions by considering the types of sensors they need.
+In the case of MMGs, this involves a set of sensors that detect emotions between players, motivations, and opportunities for actions like murder.
+To address the difficulty in defining state spaces and rewards, we designed an anytime sampling-based planner instead of using MARL.
+Unlike informed graph-based search methods, such as A* (Dechter and Pearl, 1985), our planner, inspired by algorithms like RRT* (Karaman and Frazzoli, 2011) and BIT* (Gammell et al., 2020), samples possible states with detective sensors in real-time.
+The tree-style search path is constructed and pruned dynamically until the target is reached.
+For MMGs, we combine sensors with a pruner that focuses on extracting highly suspicious murders to construct a purely questioning-driven search framework.
+PLAYER* is guided by sensors to efficiently generate high-quality questions to uncover more clues, and uses a pruner to more precisely identify the killer.
+To address the evaluation problem, we propose more quantifiable and reproducible evaluation criteria by devising a series of multiple-choice questions.
+This approach encompasses three types of questions: (1) Objective questions about the shared and individual objectives of each character in the game, (2) Reasoning questions probing the reasoning behind the given responses, and (3) Relations questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c).
+In summary, we have made the following contributions:
+(1) We propose PLAYER*, a framework that enhances the generalisability and adaptability of agents by efficiently optimising path planning in MMGs.
+(2) We designed an anytime sampling-based planner to dynamically construct and prune the search path, enhancing performance and efficiency.
+(3) We introduce a quantifiable evaluation method using multiple-choice questions focusing on character objectives, reasoning, and facts, thereby reducing the subjectivity in current evaluation practices.
+We also constructed and annotated a corresponding dataset, WellPlay, which includes 1,482 QA pairs.
 [^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
-Figure 1: The gameplay comprises three stages: introduction, discussion, and decision.
-During the introduction stage, players familiarise themselves with their characters’ background information and objectives, then briefly introduce themselves in character.
-Subsequently, they engage in interactive discussions to deduce the identity of the murderer.
 2 PLAYER*
-2.1 Problem Setting & Preliminary: Search by LLMs
-In response to the complexities of social interactions such as MMGs, we have developed an innovative interactive framework tailored for such scenarios.
-As illustrated in Figure 1, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game featuring $N_{a}$ playable characters.
-$\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$ represents the set of victims in the game.
-Each agent $a_{i}$ is assigned to a corresponding character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $o_{i}$.
-Here, $C_{i}$ is designed from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective.
-The suspicious state $s_{ik}$ is a vector representing the suspicions of killing victim $v_{k}$ from the perspective of agent $a_{i}$, which is an $(N_{a}-1)$-dim vector, where the entries can be discrete or continuous based on suspicion strength.
-The complete game rules and examples can be found in the appendix A.
-Unlike transitional planning task in which the state space can be accessed directly, $s_{i}$ in our task is generated from a language model.
-Therefore, searching in state space is done by adjusting the prompting.
-We define the possible prompts as an action space, allowing the states to be obtained through an LLM-based search: $s_{i}=P(C^{\prime}_{i},T)$, where $P$ is the LLM, $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts, and $T$ is the prompt instruction.
-The preliminary planner contains two main components:
-  • Explorer: By choosing a different prompt template $T$, which is usually from a pre-defined prompt template set, explore the corresponding $s_{i}$.
-  • Predictor: a self-reflection estimator to evaluate $s_{i}$.
-collaborate with other heuristic strategies (like A* or MC tree search) to determine if it is necessary to explore another $s_{i}$ by switching $T$, or using this $s_{i}$ and a clue to continuously explore in the next iteration.
-In general, the preliminary method only construct a prompting based planner for a give script.
-However, in the complex scenario like MMGs, the objective can be multiple, noted as $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$, including both personal and shared objectives.
-Subsequently, the agents requires to engage in a simulated discussion by asking questions and expressing opinions.
-To ensure fairness, the game imposes a limit of $m$ questions per agent per round, spanning a total of $n$ rounds.
-Following these rounds, agents are required to make decisions and cast their votes to identify the suspected murderer.
-Therefore, we need to construct a more complicated strategies considering the simulation among agents.
+In this section, we describe the MMG settings and provide an overview of how PLAYER* works.
+2.1 Problem Setting
+Figure 2: Problem Setting.
+Figure 3: PLAYER* begins with an introduction round, where each agent provides a brief self-introduction.
+Following this, agents initialise scenario understanding and begin to search for further information by questioning other agents.
+We use arrows to identify their questioning actions, with blue arrows representing questions asked by the agent itself and yellow arrows representing questions asked by other agents, which are also remembered by the given agent.
+Based on the responses they receive, agents update their suspicion state, identified by the black arrow (which decides who to ask), and the corresponding action space, identified by the pink area (which decides what questions to ask).
+The iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
+In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored for such scenarios.
+As illustrated in Figure 2, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game with $N_{a}$ playable characters and a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$.
+Each agent $a_{i}$ is assigned to a character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$.
+The script $C_{i}$ is crafted from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective.
+The suspicious state $s_{ik}$ is represented as an $(N_{a}-1)$-dimensional vector that encodes agent $a_{i}$’s beliefs about other players’ suspicions regarding the murder of victim $v_{k}$.
+The vector entries can be either discrete ($\in{\{0,1}\}$) or continuous values ($\in[0,1]$), reflecting the strength of the suspicions.
+Unlike traditional planning tasks in which the state space can be accessed directly, $s_{i}$ in our scenario is generated from a language model.
+We define the possible questions as an action space, allowing the derivation of suspicious states through an LLM-based search.
+For an illustrative example, refer to Appendix A.3.
+As shown in Figure 3, each agent is provided with character-specific information, enabling them to interact with others based on their scripts and historical interactions within the game.
+Subsequently, agents engage in simulated discussions by asking questions and expressing opinions.
+To ensure fairness, each agent is limited to asking $m$ questions per round over $n$ rounds.
+After the discussion, agents must collectively decide and vote on the suspected killer.
 2.2 Construction of Agents
-The strength of the preliminary method is that the tree structure and the potential reasoning path is clearly defined by the $T$.
-However, the lack of flexibility and the over simplified explorer may lead to a over consistency result and limit the reasoning capability of LLM.
-But the entirely free generation, especially allowing the communication between agents may lead to an uncontrollable results.
-Therefore, in our framework, we propose another two components to regularise the generation of $T$ before and after prompting respectively:
-  • Sensors: inspired by sociology, which claims that the interpersonal relationships are often conceptualised within a multidimensional space encompassing factors(Latané, 1981; Zhao et al., 2021; Trope & Liberman, 2010), we propose employing a set of sensors associated with each task.
-For example, the emotion, motivation, and suspicion, are considered as the hints to generate prompt for state searching.
-  • Pruner: After prompting, we design a pruner to reduce the state space to only few high suspicions in the next round of reasoning.
-Besides, To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement a Memory Retrieval module to store and retrieve narrative scripts and dialog logs generated during gameplay.
+To enable agents to think and act like human players, we equip them with the following components: (1) Game Rules.
+The essential information for both human and AI players.
+(2) Chain-of-Thought.
+Facilitating a sequential reasoning process to aid in problem-solving.
+(3) Sensors.
+In sociology, interpersonal relationships are often conceptualised within a multidimensional space encompassing factors like time, distance, and various other elements (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010).
+Drawing inspiration from this concept, we propose employing a set of sensors associated with each task.
+For example, in a murder mystery game, these factors could include emotion, motivation, and suspicion.
+(4) Memory Retrieval.
+To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement this module to store and retrieve narrative scripts and dialog logs generated during gameplay.
 Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent.
@@
 This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length.
-It ensures consistent and coherent interactions among agents while assisting in their strategic planning.
-Figure 2: PLAYER*’s planning strategy begins with an introduction round, during which each agent provides a brief self-introduction.
-Agents then initialise scenario understanding and begin questioning other agents.
-Subsequently, they start searching for further information by questioning other agents.
-Agents then adapt their level of suspicion and intended search area based on their received response, mirroring human gameplay strategies.
-This iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
+It ensures consistent and coherent interactions among agents while assisting in their strategic planning throughout the game.
+(5) Self-Reflection.
+Preserving historical experience by reflecting on the current situation to provide a more detailed, accurate, and profound answer.
 2.3 PLAYER* Planning Strategy
-To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximate the search domain through sampling, and planning the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions.
+To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions.
 As illustrated in Figure 4, this framework is fundamentally composed of two key components:
+Figure 4: Search and Approximate.
+PLAYER* generates questions based on sensor-detected values and identifies the killer more precisely with a prunner.
+This dynamic question generation allows agents to adjust to the specific circumstances of each game session, ensuring a more focused and effective investigation.
 Questioning Strategy in Searching based on Sensors
 Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components.
-Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023b), we generate actions directed towards the targets based on the values detected by those sensors.
+Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023), we generate actions directed towards the targets based on the values detected by those sensors.
 The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors.
-For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Suspicion}]$: Emotion, which assesses one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible; Motivation, evaluating from one’s perspective whether the individual possesses a motive to be the perpetrator; and Suspicion, objectively determining whether they had the opportunity to commit the crime.
-Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration, while the reward for such a cost is the information that can be obtained from the response.
-Under the assumption that increased inquiries to an agent corresponds to the heightened exploration within the world space related to that agent, the expected reward associated with questioning the same agent decreases in proportion to the number of inquiries posed.
-Therefore, it is necessary to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if this person might be highly suspicious.
-We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ about the agent to be questioned, as well as the number of inquiries that have already posed $|Q_{i}|$.
-Specifically, for agent $a_{i}$ to be questioned, PLAYER* would generate the question $q$ using the values of sensors $S_{i}$ in conjunction with the expected reward $P_{informative}$ and other related information $(C^{\prime}_{i},D_{i})$, that is $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
+For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Opportunity}]$: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible.
+Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator.
+Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime.
+The sensor values are obtained based on the related information retrieved from the script and dialog, represented by the function $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$, where $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts and $D_{i}$ denotes the associated dialogue history.
+Then, for agent $a_{i}$ to be questioned, PLAYER* uses LLMs to sample the question $q$ from the action space given the values of sensors $S_{i}$ and related information $(C^{\prime}_{i},D_{i})$, represented by the function $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
+Action Space Refinement with a Pruner
+Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents.
+Each agent first updates its understanding of interpersonal relationships, thereby adjusting sensor values $S$.
+Subsequently, it also updated $P_{informative}$ based on the revised information and changes in sensor values.
+Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration.
+The reward for questioning is the information that can be obtained from the anticipated response.
+As the number of inquiries directed at an agent increases, the assumed exploration within the world space relevant to that agent also rises.
+Consequently, the expected reward associated with questioning the same agent diminishes proportionally to the number of inquiries already made.
+Therefore, it becomes essential to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if they appear highly suspicious.
+We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ concerning the agent to be questioned, and the number of inquiries already posed $|Q_{i}|$.
+In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals.
+This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning.
+Emulating the cognitive pattern of human beings, we refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets (See Lower Part of Figure 4).
+We did not impose a hard restriction on the refinement process and allow for flexible adjustment of the suspicious state.
+For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked.
+This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility.
+Then we update the action space based on the suspicious state.
+This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads.
+By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
 Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, Suspicious States $s$, Max Round $n$
@@
 11
-      // Constrain action space
+      // Update the suspect list
 12       for i = 1 to $N_{a}$ do
 13             for k = 1 to $N_{v}$ do
-14                   $s_{ik}$ = Refinement_Action_Space($a_{i},a_{j},v_{i}$)
+14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{i}$)
 15
@@
 Algorithm 1 Anytime Sampling-based Planning in Multi-agent Interaction
-Action Space Refinement with a Pruner
-Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents.
-Each agent first updates its comprehension of the interpersonal relationships, thereby adjusting sensor values: $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$.
-Subsequently, it also updated $P_{informative}$ based on the updated information $(S_{i},D_{i})$ and the changes in sensor values $\Delta S_{i}$.
-In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals.
-This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning.
-Emulating the cognitive pattern of human beings, we introduce an Action Space Refinement strategy to refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets.
-We did not impose a hard restriction on the refinement process and allowed for flexible adjustment of the action space.
-For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked.
-This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility.
-This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads.
-By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
 Algorithm 1 outlines a detailed procedure of the game process.
 While initially designed for MMGs, the framework’s adaptable nature allows it to be readily applied to various games by adjusting the game rules and task-specific sensors, thereby enhancing the gaming experience and offering a structured way for agents to handle complex social dynamics efficiently.
+3 The WellPlay Dataset
+Script Agents Victims #token(CN) #token(EN) Question
+avg overall avg overall Objective Reasoning Relations overall
+Death Wears White 9 1 3,191 28,716 1,742 15,681 10 102 72 184
+Ghost Revenge 7 3 5,488 38,415 3,960 27,723 19 152 69 240
+Danshui Villa 7 2 5,111 35,779 3,339 23,370 12 128 63 203
+Unfinished Love 7 2 2,501 17,507 1,652 11,562 12 61 72 145
+Cruise Incident 5 1 1,263 6,313 808 4,040 4 24 30 58
+Sin 4 1 2,121 8,485 1,378 5,512 3 20 21 44
+Deadly Fountain 4 1 1,852 7,410 1,194 4,775 3 21 12 36
+Unbelievable Incident 5 1 3,182 15,912 2,012 10,062 4 24 15 43
+Desperate Sunshine 4 1 3,370 13,481 2,219 8,874 3 18 36 57
+Riverside Inn 4 1 1,910 7,638 1,257 5,028 3 18 18 39
+Solitary Boat Firefly 6 4 8,894 53,362 6,874 41,244 20 109 69 198
+Manna 6 3 9,028 54,169 6,492 38,954 24 123 88 235
+Avg 5.67 1.75 3,993 23,932 2,744 16,402 9.75 66.67 47.08 125.50
+Sum 68 21 47,911 287,187 32,927 196,825 117 800 565 1482
+Table 1: Dataset Statistics.
+Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively.
+Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types.
+The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
+WellPlay Dataset
+We built an evaluation dataset, called WellPlay, from the existing Conan dataset (Zhao et al., 2024c), derived from background narratives created for MMGs.
+Conan labels relationships from each character’s perspective, along with ground truth relationships, consolidating all available information.
+To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers.
+Besides existing labels of interpersonal relationships that gauge agents’ understanding of the situation, we also recruited four annotators to label character objectives and reasoning.
+The dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions.
+On average, each game features 5.67 agents and 1.75 victims as detailed in Table 1 .
+Our evaluation questions are grounded in characters’ objectives in the game and are crafted with input from human players, focusing on the following question types:
+(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
+(2) Reasoning.
+This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
+(3) Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
+Evaluation Metrics
+We report the performance of each type of question and also calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance.
+The awarded points follow the scoring system used in MMGs, which awards 10 points for objective questions, 5 points for reasoning questions, and 2 points for relation questions.
+For these calculations, we utilise weighted mean and weighted standard deviation methodologies.
+In addition to assessing the accuracy of the three types of questions and the overall score, we also evaluate the Win Rates of players for performance assessment.
+The game rules are as follows: every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) will be eliminated.
+Additionally, we assume that murderers would always vote for someone else instead of voting for themselves.
+If the true murderer is successfully voted out, the game is considered a victory for the players seeking to identify the culprit.
+The detailed about the dataset and calculations is provided in Appendix B.
 4 Experiments
@@
 Base Models
-We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 and text-embedding-ada-002.
-To minimise randomness, we conducted the evaluation experiments 3 times and report the average and variance.
+We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 (Default) and text-embedding-ada-002.
+To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and variance.
 The detailed introduction to the experiment settings is in Appendix C.1.
 Baselines
-For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games, such as Werewolf (Xu et al., 2023a), Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b), and ThinkThrice (Wu et al., 2024).
-Additionally, we assess the performance of agents that do not actively participate in the game but have direct access to either their own script (Single Script Access) or the scripts of all agents (Full Script Access).
-In werewolf, questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario.
-In O-CoT, agent interactions are driven by their set objectives.
-ThinkThrice has its agents crafting questions from retrieved memory and the current scenario.
-However, the efficiency of exploration with the questioning strategies is low, as evidenced by their performance, which falls significantly short of full script access.
-In contrast, our method leverages sensor data to identify the optimal domain to question and guide the generation of questions.
-The detailed comparison is in Appendix C.2.
-WellPlay Dataset
-We create an evaluation dataset, called the WellPlay Dataset, based on the existing dataset Conan (Zhao et al., 2024c), originated from background narratives created for MMGs, which provides annotated relationships between characters.
-We aim to concentrate on the most challenging and significant relationships that assess the agents’ comprehension of the scenario.
-Our approach is grounded in the original game’s objectives and informed by a trial round involving human players.
-This method enabled us to formulate the following types of evaluation questions in line with the human players:
-1. Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
-2.
-Reasoning.
-This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
-3. Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
-To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information.
-This dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions.
-On average, each game features 5.67 agents and 1.75 victims.
-More detailed statistics are available in the appendix B.1.
-Evaluation Metrics
-We adopted the scoring system used in the game to evaluate agents’ performance: Awarding 10 points for achieving an objective, 5 points for correct reasoning, and 2 points for providing additional information based on the players’ understanding of the current situation.
-However, instead of using open-ended questions with human players, we assess each agent’s performance using our constructed dataset, WellPlay, to facilitate replication and comparison for future research.
-WellPlay comprises three types of questions: objective, reasoning, and relations, which are assigned 10, 5, and 2 points, respectively.
-We calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance.
-Additionally, we assess the accuracy of each question type.
-For the computation of the overall score for performance assessment, we have utilized weighted mean and weighted standard deviation methodologies.
-The detailed algorithm for these calculations is in Appendix B.2.
+For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games.
+Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games.
+We adapted their methodology to MMGs for a fair comparison, with implementation details in Appendix C.2.
+(1) Werewolf (Xu et al., 2023).
+Werewolf is another multi-agent game, where players identify werewolves through group discussion.
+Questions are chosen from a role-specific predefined list to facilitate game progression, alongside questions generated based on the current scenario.
+We adapt its pre-set game instructions and role-specific information to MMGs settings.
+(2) Objective-Guided Chain of Thought (O-CoT) (Park et al., 2023; Zhao et al., 2024b).
+Agents think, reflect, and choose who and what to ask based on their objectives.
+We use the framework from previous works, only replacing the agents’ objectives with those set in MMGs.
+(3) ThinkThrice (Wu et al., 2024).
+Designed for MMGs, agents craft questions from retrieved memory and the current scenario.
+(4) Personal Perspective.
+For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script.
+(5) Omniscient Perspective.
+Agents do not actively participate but make their final decisions based on all agents’ scripts.
+To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and PLAYER*).
+This allows us to fairly compare the performance of the side trying to find the murderer, who face the same level of opposition from the murderer’s side.
 4.2 Results
 Performance Evaluation Results
-Table 1 presents the evaluation of agents’ performance across 12 unique games of varying complexity and settings.
-The number of evaluation questions (“#QA”) varies based on script complexity, with more complex scripts generating a larger volume of questions.
-The SSA (Single Script Access) measures agents’ performance with access to only their own script.
-This setup is designed to represent the starting point for searching.
-The FSA (Full Script Access) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint.
-Ideally, FSA should aim to achieve an accuracy of 1.
-However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model.
-Despite this, FSA still scores significantly higher than SSA.
-In addition, FSA also serves as an indicator of the complexity of the script.
-We can observe that PLAYER* exhibits superior performance other baselinse across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents.
-Additionally, PLAYER* significantly outperforms others in objective questions, even surpassing FSA, which has access to all information.
-This showcases the effectiveness of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective.
-Despite some baselines exhibiting competitive performance in reasoning or relation questions, their significant drop in performance concerning objectives indicates a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals.
+Figure 5: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
+The Personal Perspective (PP) is designed to represent the starting point for searching.
+The Omniscient Perspective (OP) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint.
+Figure 5 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table A3).
+PLAYER* shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents.
+Notably, PLAYER* significantly outperforms others in objective questions, even surpassing OP, which has access to all information.
+The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies.
+It represents the ideal “search” endpoint.
+However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model.
+This showcases the “approximate” ability of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective.
+Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
+The results demonstrate that PLAYER* outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability.
+The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals.
+This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
+Our findings reveal a significant discrepancy between the win rate and the objective metric (the average of non-murderer players correctly identifying the murderer) for agents, with OP exhibiting a small difference and PP showing the largest difference.
+The case analysis unveils that in OP games, agents make similar decisions, resulting in voting outcomes that are either mostly correct or mostly wrong, leading to consistency between individual and collective voting results.
+Conversely, in PP games, agents base their decisions on their individual scripts, leading to independent actions and difficulty in reaching a consensus to vote out the murderer.
+The performance of the agents falls between that of OP and PP, with vote distributions more dispersed than OP but more concentrated than PP.
+O-CoT suffered from unfortunate circumstances, falling short by a single vote in numerous games.
+This resulted in O-CoT being on the verge of victory in many instances but ultimately yielding poor overall results.
 We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4.
 The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
-Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
-Werewolf O-CoT ThinkThrice PLAYER*
-Death Wears White (9 players, 1 victim) Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
-Reasoning $102$$454_{\pm.012}.$$503_{\pm.015}.$$350_{\pm.021}.$$395_{\pm.031}.$$392_{\pm.010}.$$408_{\pm.030}.$
-Relations $72$$431_{\pm.036}.$$425_{\pm.031}.$$454_{\pm.035}.$$472_{\pm.037}.$$398_{\pm.032}.$$500_{\pm.048}.$
-Overall $184$$420_{\pm.007}.$$480_{\pm.011}.$$367_{\pm.015}.$$393_{\pm.009}.$$377_{\pm.013}.$$407_{\pm.007}.$
-Ghost Revenge (7 players, 3 victims) Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
-Reasoning $152$$390_{\pm.020}.$$533_{\pm.030}.$$417_{\pm.023}.$$410_{\pm.019}.$$450_{\pm.056}.$$399_{\pm.040}.$
-Relations $69$$246_{\pm.014}.$$411_{\pm.103}.$$348_{\pm.014}.$$309_{\pm.047}.$$420_{\pm.038}.$$386_{\pm.030}.$
-Overall $240$$362_{\pm.016}.$$483_{\pm.020}.$$399_{\pm.009}.$$381_{\pm.049}.$$438_{\pm.030}.$$417_{\pm.032}.$
-Danshui Villa (7 players, 2 victims) Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
-Reasoning $128$$315_{\pm.020}.$$422_{\pm.016}.$$307_{\pm.044}.$$352_{\pm.008}.$$347_{\pm.005}.$$367_{\pm.008}.$
-Relations $63$$323_{\pm.040}.$$476_{\pm.028}.$$344_{\pm.051}.$$381_{\pm.069}.$$333_{\pm.016}.$$370_{\pm.056}.$
-Overall $203$$326_{\pm.014}.$$403_{\pm.028}.$$293_{\pm.040}.$$349_{\pm.012}.$$339_{\pm.013}.$$367_{\pm.011}.$
-Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
-Reasoning $61$$552_{\pm.025}.$$563_{\pm.025}.$$563_{\pm.010}.$$486_{\pm.025}.$$552_{\pm.019}.$$601_{\pm.019}.$
-Relations $72$$500_{\pm.028}.$$546_{\pm.021}.$$523_{\pm.071}.$$583_{\pm.028}.$$556_{\pm.042}.$$616_{\pm.016}.$
-Overall $145$$457_{\pm.017}.$$475_{\pm.033}.$$469_{\pm.018}.$$467_{\pm.004}.$$460_{\pm.030}.$$525_{\pm.029}.$
-Cruise Incident (5 players, 1 victim) Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
-Reasoning $24$$653_{\pm.064}.$$472_{\pm.064}.$$542_{\pm.042}.$$528_{\pm.024}.$$528_{\pm.105}.$$514_{\pm.024}.$
-Relations $30$$267_{\pm.058}.$$411_{\pm.102}.$$456_{\pm.020}.$$456_{\pm.051}.$$422_{\pm.039}.$$434_{\pm.058}.$
-Overall $58$$459_{\pm.012}.$$415_{\pm.079}.$$510_{\pm.037}.$$503_{\pm.017}.$$509_{\pm.107}.$$520_{\pm.025}.$
-Sin (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
-Reasoning $20$$483_{\pm.058}.$$567_{\pm.029}.$$767_{\pm.058}.$$600_{\pm.100}.$$717_{\pm.058}.$$600_{\pm.000}.$
-Relations $21$$476_{\pm.082}.$$587_{\pm.099}.$$429_{\pm.000}.$$699_{\pm.055}.$$524_{\pm.048}.$$698_{\pm.027}.$
-Overall $44$$397_{\pm.029}.$$531_{\pm.053}.$$570_{\pm.000}.$$539_{\pm.029}.$$564_{\pm.012}.$$577_{\pm.053}.$
-Deadly Fountain (4 players, 1 victim) Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
-Reasoning $21$$413_{\pm.153}.$$667_{\pm.126}.$$460_{\pm.027}.$$476_{\pm.082}.$$539_{\pm.055}.$$587_{\pm.028}.$
-Relations $12$$305_{\pm.048}.$$389_{\pm.048}.$$333_{\pm.084}.$$222_{\pm.096}.$$222_{\pm.048}.$$361_{\pm.127}.$
-Overall $36$$319_{\pm.097}.$$604_{\pm.104}.$$354_{\pm.022}.$$369_{\pm.053}.$$390_{\pm.043}.$$463_{\pm.041}.$
-Unbelievable Incident (5 players, 1 victim) Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
-Reasoning $24$$389_{\pm.105}.$$528_{\pm.064}.$$431_{\pm.064}.$$361_{\pm.064}.$$264_{\pm.127}.$$583_{\pm.042}.$
-Relations $15$$422_{\pm.102}.$$822_{\pm.102}.$$644_{\pm.077}.$$756_{\pm.102}.$$289_{\pm.154}.$$556_{\pm.077}.$
-Overall $43$$330_{\pm.081}.$$481_{\pm.020}.$$444_{\pm.037}.$$382_{\pm.035}.$$230_{\pm.034}.$$509_{\pm.038}.$
-Desperate Sunshine (4 players, 1 victim) Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
-Reasoning $18$$556_{\pm.056}.$$778_{\pm.096}.$$648_{\pm.032}.$$704_{\pm.032}.$$759_{\pm.064}.$$741_{\pm.085}.$
-Relations $36$$537_{\pm.064}.$$611_{\pm.048}.$$556_{\pm.028}.$$630_{\pm.043}.$$500_{\pm.056}.$$565_{\pm.070}.$
-Overall $57$$514_{\pm.013}.$$680_{\pm.047}.$$599_{\pm.023}.$$618_{\pm.020}.$$647_{\pm.075}.$$680_{\pm.047}.$
-Riverside Inn (4 players, 1 victim) Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
-Reasoning $18$$519_{\pm.032}.$$667_{\pm.056}.$$463_{\pm.064}.$$500_{\pm.056}.$$685_{\pm.032}.$$667_{\pm.056}.$
-Relations $18$$407_{\pm.085}.$$500_{\pm.096}.$$463_{\pm.085}.$$426_{\pm.064}.$$389_{\pm.000}.$$426_{\pm.032}.$
-Overall $39$$479_{\pm.032}.$$671_{\pm.036}.$$459_{\pm.073}.$$387_{\pm.020}.$$614_{\pm.018}.$$590_{\pm.023}.$
-Solitary Boat Firefly (6 players, 4 victims) Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
-Reasoning $109$$318_{\pm.014}.$$569_{\pm.018}.$$370_{\pm.014}.$$425_{\pm.046}.$$401_{\pm.014}.$$379_{\pm.005}.$
-Relations $69$$483_{\pm.017}.$$507_{\pm.014}.$$536_{\pm.014}.$$565_{\pm.038}.$$503_{\pm.044}.$$594_{\pm.087}.$
-Overall $198$$332_{\pm.015}.$$544_{\pm.024}.$$354_{\pm.023}.$$430_{\pm.016}.$$375_{\pm.015}.$$444_{\pm.030}.$
-Manna (6 players, 3 victims) Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
-Reasoning $123$$458_{\pm.005}.$$393_{\pm.040}.$$569_{\pm.028}.$$499_{\pm.033}.$$526_{\pm.017}.$$539_{\pm.021}.$
-Relations $88$$447_{\pm.058}.$$640_{\pm.036}.$$538_{\pm.043}.$$614_{\pm.030}.$$511_{\pm.034}.$$580_{\pm.046}.$
-Overall $235$$408_{\pm.017}.$$434_{\pm.002}.$$525_{\pm.017}.$$505_{\pm.004}.$$501_{\pm.025}.$$553_{\pm.045}.$
-Overall Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
-Reasoning $800$$417_{\pm.095}.$$508_{\pm.094}.$$437_{\pm.104}.$$430_{\pm.077}.$$449_{\pm.113}.$$465_{\pm.106}.$
-Relations $565$$411_{\pm.102}.$$513_{\pm.110}.$$466_{\pm.090}.$$495_{\pm.133}.$$444_{\pm.089}.$$500_{\pm.109}.$
-Overall $1482$$386_{\pm.057}.$$484_{\pm.072}.$$415_{\pm.074}.$$424_{\pm.067}.$$426_{\pm.087}.$$461_{\pm.086}.$
-Table 1: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
-SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
 Efficiency and Cost Comparison
-As shown in Table 2, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table.
-The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages  [^2] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
-Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved.
-This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
+As shown in Figure 6, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table (Table A5).
+The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages.
+[^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
+Figure 6: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
+Costs are reported for one complete gameplay and one evaluation process for each script.
 The costs are divided into Gameplay and Evaluation phases.
-During Gameplay, agents actively participate in the game, generating actions and responses.
-In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset.
-Notably, SSA and FSA methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner.
+Notably, PP and OP methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner.
+We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps.
+In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives.
 PLAYER* stands out for its cost-efficiency and performance.
@@
 This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase.
-Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in SSA and FSA costs arising from their differing levels of script access.
-Script #Tokens Stage SSA FSA Agent’s Response After Playing the Game
-Werewolf O-CoT ThinkThrice PLAYER*
-Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $3.363$
-Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
-Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $5.947$
-Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
-Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $4.603$
-Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
-Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $3.505$
-Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
-Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.846$
-Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
-Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
-Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
-Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.596$
-Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
-Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $1.150$
-Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
-Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.701$
-Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
-Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.561$
-Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
-Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $6.800$
-Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
-Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $5.028$
-Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
-Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
-Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
-Table 2: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
-#$\#Tokens$ represent the average length of each character’s script.
-Costs are reported for one complete gameplay and one evaluation process for each script.
+In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset.
+Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in PP and OP costs arising from their differing levels of script access.
 4.3 Ablation Studies
+We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7.
 Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions.
@@
 However, these methods often require extensive time and computational resources and lack linguistic communication capabilities.
-With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024).
-Exemplified by AlphaGo (Silver et al., 2017), which demonstrated exceptional performance against human competitors, self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance.
-This idea has also been adapted for evaluation through negotiation games (Davidson et al., 2024).
-Compared to classic methods (Wang & Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states.
+With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024).
+Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance.
+Compared to classic methods (Wang and Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states.
 However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024).
@@
 To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced.
-These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang & Li, 2023; Ismail et al., 2024).
-Adapting LLM-as-a-Judge prompting to evaluate performance (Yuan et al., 2024) or select agents (Liu et al., 2023c) has become a popular approach, but negative results on self-reflection were also investigated (Huang et al., 2024), leaving LLM’s role as a self-reflective agent as an unresolved question.
-We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020b; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
+These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024).
+We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
 6 Conclusion
@@
 We propose a quantifiable evaluation method, contribute the WellPlay dataset, and demonstrate PLAYER*’s superiority, advancing effective reasoning agents in complex environments.
-6 Ethics Statement
+Limitations
+We identify some of the limitations below:
+Limited Annotation Scope
+Due to the cost of manual annotation and running experiments, we were unable to annotate all 100 games in the Conan dataset.
+However, we believe that the current scope of our study is sufficient to demonstrate the efficiency of our methodology.
+Future research can expand upon our work by annotating additional games and applying our approach to MMGs on a larger scale.
+Limited Evaluation of LLMs
+Due to the high cost associated with running experiments on LLMs, our experiments were primarily conducted using GPT-3.5.
+To further support our findings, we also performed smaller-scale experiments on LLaMa, which yielded results consistent with those obtained from GPT-3.5.
+However, we acknowledge that our inability to run more extensive evaluations across a wider range of popular LLMs, such as GPT-4 or lesser-known models, may have limited our ability to uncover additional insights.
+The cost of API usage for models like GPT-4 and the substantial running time required for models like LLaMa posed significant constraints on our experimental scope.
+These limitations highlight the need for future research to explore more cost-effective and efficient methods for evaluating and comparing the performance of various LLMs in the context of complex social deduction games.
+Ethics Statement
 Please note that MMGs and our WellPlay dataset, created to assess agents’ behaviors in MMGs, may include descriptions of violent events, actions, or characters.
@@
 Appendix B The WellPlay Dataset
-B.1 Details of Dataset
-Script Agents Victims #token(CN) #token(EN) Question
-avg overall avg overall Objective Reasoning Relations overall
-Death Wears White 9 1 3190.67 28716 1742.33 15681 10 102 72 184
-Ghost Revenge 7 3 5487.86 38415 3960.43 27723 19 152 69 240
-Danshui Villa 7 2 5111.29 35779 3338.57 23370 12 128 63 203
-Unfinished Love 7 2 2501.00 17507 1651.71 11562 12 61 72 145
-Cruise Incident 5 1 1262.60 6313 808.00 4040 4 24 30 58
-Sin 4 1 2121.25 8485 1378.00 5512 3 20 21 44
-Deadly Fountain 4 1 1852.50 7410 1193.75 4775 3 21 12 36
-Unbelievable Incident 5 1 3182.40 15912 2012.40 10062 4 24 15 43
-Desperate Sunshine 4 1 3370.25 13481 2218.50 8874 3 18 36 57
-Riverside Inn 4 1 1909.50 7638 1257.00 5028 3 18 18 39
-Solitary Boat Firefly 6 4 8893.67 53362 6874.00 41244 20 109 69 198
-Manna 6 3 9028.17 54169 6492.33 38954 24 123 88 235
-Avg 5.67 1.75 3992.60 23932.25 2743.92 16402.08 9.75 66.67 47.08 125.50
-Sum 68 21 47911.16 287187 32927.02 196825 117 800 565 1482
-Table A1: Dataset Statistics.
-Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively.
-Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types.
 In this work, we leverage the Conan dataset, originally constructed by Zhao et al.
@@
 Our dataset consists of two main components: Scripts and Evaluation questions.
+B.1 Scripts
 Each script features distinct narratives for individual characters. We have rephrased the original script from the Conan dataset into two parts:
@@
 Table A2: Predefined questions for Werewolf method.
+B.2 Evaluation questions
 The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts.
@@
 Specifically, Type A questions are valued at 10 points, Type B questions at 5 points, and Type C questions at 2 points.
-Detailed script statistics and evaluation question metrics are presented in TableA1.
-B.2 Overall Performance Computing
+Detailed script statistics and evaluation question metrics are presented in Table1.
+B.3 Overall Performance Computing
 In calculating the overall score for performance, we have employed both the weighted mean and the weighted standard deviation.
@@
 $s_{w}=\sqrt{s_{w}^{2}}$
+B.4 Annotation Details
+We recruited four human experts to read all the scripts and write the questions and answers.
+We conducted training sessions for them and organised game sessions to let them play and become familiar with the game flow.
+They were compensated at an hourly rate of \$31.92, with each narrative estimated to take about 5-20 hours to complete, depending on the complexity of the game.
+B.5 Disclaimers of Risks
+Consider that the detective narrative dataset may contain descriptions of violent events, actions, or characters, we warn annotators of such content before the annotation process:
+Please be advised that the detective narrative dataset may contain descriptions of violent events, actions, or characters.
+This content is included solely for academic, research, and narrative analysis purposes.
+It is not meant to glorify or trivialise violence in any form.
+Annotators should be aware of potential exposure to violent content and engage with the dataset professionally and responsibly.
 Appendix C Implementation
@@
 Given the game’s rules and questioning sequence in our comparisons, we segment the discussion into two phases: the questioning phase and the answering phase.
-Considering the strategic core of the game, a direct confession from the murderer would compromise its competitive nature.
-Hence, in all the methodologies we explore, we maintain that the murderer’s responses follow a predefined template, detailed in C.6.2.
-This method preserves the game’s integrity and level of challenge.
   1. 1.
-    Werewolf(Xu et al., 2023a)
+    Werewolf(Xu et al., 2023)
     Question Generation:
@@
   4. 4.
-    The SSA (Single Script Access)
-    The SSA assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
+    The PP (Personal Perspective)
+    The PP assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
   5. 5.
-    The FSA (Full Script Access)
-    The FSA evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
-C.3 Performance Evaluation Results for English Dataset
+    The OP (Omniscient Perspective)
+    The OP evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
+C.3 Performance Evaluation Results
+Script Evaluation #QA PP OP Agent’s Response After Playing the Game
+Werewolf O-CoT ThinkThrice PLAYER*
+Death Wears White (9 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
+Reasoning $102$$454_{\pm.012}.$$503_{\pm.015}.$$350_{\pm.021}.$$395_{\pm.031}.$$392_{\pm.010}.$$408_{\pm.030}.$
+Relations $72$$431_{\pm.036}.$$425_{\pm.031}.$$454_{\pm.035}.$$472_{\pm.037}.$$398_{\pm.032}.$$500_{\pm.048}.$
+Overall $184$$420_{\pm.007}.$$480_{\pm.011}.$$367_{\pm.015}.$$393_{\pm.009}.$$377_{\pm.013}.$$407_{\pm.007}.$
+Ghost Revenge (7 players, 3 victims) Voting -- .$222_{\pm.416}.$$333_{\pm.471}.$$222_{\pm.416}.$$111_{\pm.314}.$$333_{\pm.471}.$$333_{\pm.471}.$
+Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
+Reasoning $152$$390_{\pm.020}.$$533_{\pm.030}.$$417_{\pm.023}.$$410_{\pm.019}.$$450_{\pm.056}.$$399_{\pm.040}.$
+Relations $69$$246_{\pm.014}.$$411_{\pm.103}.$$348_{\pm.014}.$$309_{\pm.047}.$$420_{\pm.038}.$$386_{\pm.030}.$
+Overall $240$$362_{\pm.016}.$$483_{\pm.020}.$$399_{\pm.009}.$$381_{\pm.049}.$$438_{\pm.030}.$$417_{\pm.032}.$
+Danshui Villa (7 players, 2 victims) Voting -- .$167_{\pm.373}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
+Reasoning $128$$315_{\pm.020}.$$422_{\pm.016}.$$307_{\pm.044}.$$352_{\pm.008}.$$347_{\pm.005}.$$367_{\pm.008}.$
+Relations $63$$323_{\pm.040}.$$476_{\pm.028}.$$344_{\pm.051}.$$381_{\pm.069}.$$333_{\pm.016}.$$370_{\pm.056}.$
+Overall $203$$326_{\pm.014}.$$403_{\pm.028}.$$293_{\pm.040}.$$349_{\pm.012}.$$339_{\pm.013}.$$367_{\pm.011}.$
+Unfinished Love (7 players, 2 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
+Reasoning $61$$552_{\pm.025}.$$563_{\pm.025}.$$563_{\pm.010}.$$486_{\pm.025}.$$552_{\pm.019}.$$601_{\pm.019}.$
+Relations $72$$500_{\pm.028}.$$546_{\pm.021}.$$523_{\pm.071}.$$583_{\pm.028}.$$556_{\pm.042}.$$616_{\pm.016}.$
+Overall $145$$457_{\pm.017}.$$475_{\pm.033}.$$469_{\pm.018}.$$467_{\pm.004}.$$460_{\pm.030}.$$525_{\pm.029}.$
+Cruise Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$333_{\pm.471}.$$333_{\pm.471}.$
+Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
+Reasoning $24$$653_{\pm.064}.$$472_{\pm.064}.$$542_{\pm.042}.$$528_{\pm.024}.$$528_{\pm.105}.$$514_{\pm.024}.$
+Relations $30$$267_{\pm.058}.$$411_{\pm.102}.$$456_{\pm.020}.$$456_{\pm.051}.$$422_{\pm.039}.$$434_{\pm.058}.$
+Overall $58$$459_{\pm.012}.$$415_{\pm.079}.$$510_{\pm.037}.$$503_{\pm.017}.$$509_{\pm.107}.$$520_{\pm.025}.$
+Sin (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$333_{\pm.471}.$
+Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
+Reasoning $20$$483_{\pm.058}.$$567_{\pm.029}.$$767_{\pm.058}.$$600_{\pm.100}.$$717_{\pm.058}.$$600_{\pm.000}.$
+Relations $21$$476_{\pm.082}.$$587_{\pm.099}.$$429_{\pm.000}.$$699_{\pm.055}.$$524_{\pm.048}.$$698_{\pm.027}.$
+Overall $44$$397_{\pm.029}.$$531_{\pm.053}.$$570_{\pm.000}.$$539_{\pm.029}.$$564_{\pm.012}.$$577_{\pm.053}.$
+Deadly Fountain (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
+Reasoning $21$$413_{\pm.153}.$$667_{\pm.126}.$$460_{\pm.027}.$$476_{\pm.082}.$$539_{\pm.055}.$$587_{\pm.028}.$
+Relations $12$$305_{\pm.048}.$$389_{\pm.048}.$$333_{\pm.084}.$$222_{\pm.096}.$$222_{\pm.048}.$$361_{\pm.127}.$
+Overall $36$$319_{\pm.097}.$$604_{\pm.104}.$$354_{\pm.022}.$$369_{\pm.053}.$$390_{\pm.043}.$$463_{\pm.041}.$
+Unbelievable Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
+Reasoning $24$$389_{\pm.105}.$$528_{\pm.064}.$$431_{\pm.064}.$$361_{\pm.064}.$$264_{\pm.127}.$$583_{\pm.042}.$
+Relations $15$$422_{\pm.102}.$$822_{\pm.102}.$$644_{\pm.077}.$$756_{\pm.102}.$$289_{\pm.154}.$$556_{\pm.077}.$
+Overall $43$$330_{\pm.081}.$$481_{\pm.020}.$$444_{\pm.037}.$$382_{\pm.035}.$$230_{\pm.034}.$$509_{\pm.038}.$
+Desperate Sunshine (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$667_{\pm.471}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
+Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
+Reasoning $18$$556_{\pm.056}.$$778_{\pm.096}.$$648_{\pm.032}.$$704_{\pm.032}.$$759_{\pm.064}.$$741_{\pm.085}.$
+Relations $36$$537_{\pm.064}.$$611_{\pm.048}.$$556_{\pm.028}.$$630_{\pm.043}.$$500_{\pm.056}.$$565_{\pm.070}.$
+Overall $57$$514_{\pm.013}.$$680_{\pm.047}.$$599_{\pm.023}.$$618_{\pm.020}.$$647_{\pm.075}.$$680_{\pm.047}.$
+Riverside Inn (4 players, 1 victim) Voting -- .$333_{\pm.471}.$ $1.000_{\pm.000}$$333_{\pm.471}.$$000_{\pm.000}.$ $1.000_{\pm.000}$$333_{\pm.471}.$
+Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
+Reasoning $18$$519_{\pm.032}.$$667_{\pm.056}.$$463_{\pm.064}.$$500_{\pm.056}.$$685_{\pm.032}.$$667_{\pm.056}.$
+Relations $18$$407_{\pm.085}.$$500_{\pm.096}.$$463_{\pm.085}.$$426_{\pm.064}.$$389_{\pm.000}.$$426_{\pm.032}.$
+Overall $39$$479_{\pm.032}.$$671_{\pm.036}.$$459_{\pm.073}.$$387_{\pm.020}.$$614_{\pm.018}.$$590_{\pm.023}.$
+Solitary Boat Firefly (6 players, 4 victims) Voting -- .$167_{\pm.373}.$$583_{\pm.493}.$$000_{\pm.000}.$$250_{\pm.433}.$$167_{\pm.373}.$$500_{\pm.500}.$
+Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
+Reasoning $109$$318_{\pm.014}.$$569_{\pm.018}.$$370_{\pm.014}.$$425_{\pm.046}.$$401_{\pm.014}.$$379_{\pm.005}.$
+Relations $69$$483_{\pm.017}.$$507_{\pm.014}.$$536_{\pm.014}.$$565_{\pm.038}.$$503_{\pm.044}.$$594_{\pm.087}.$
+Overall $198$$332_{\pm.015}.$$544_{\pm.024}.$$354_{\pm.023}.$$430_{\pm.016}.$$375_{\pm.015}.$$444_{\pm.030}.$
+Manna (6 players, 3 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$222_{\pm.416}.$$333_{\pm.471}.$$444_{\pm.497}.$$556_{\pm.497}.$
+Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
+Reasoning $123$$458_{\pm.005}.$$393_{\pm.040}.$$569_{\pm.028}.$$499_{\pm.033}.$$526_{\pm.017}.$$539_{\pm.021}.$
+Relations $88$$447_{\pm.058}.$$640_{\pm.036}.$$538_{\pm.043}.$$614_{\pm.030}.$$511_{\pm.034}.$$580_{\pm.046}.$
+Overall $235$$408_{\pm.017}.$$434_{\pm.002}.$$525_{\pm.017}.$$505_{\pm.004}.$$501_{\pm.025}.$$553_{\pm.045}.$
+Overall Voting -- .$074_{\pm.117}.$$326_{\pm.342}.$$176_{\pm.257}.$$058_{\pm.115}.$$245_{\pm.326}.$$282_{\pm.308}.$
+Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
+Reasoning $800$$417_{\pm.095}.$$508_{\pm.094}.$$437_{\pm.104}.$$430_{\pm.077}.$$449_{\pm.113}.$$465_{\pm.106}.$
+Relations $565$$411_{\pm.102}.$$513_{\pm.110}.$$466_{\pm.090}.$$495_{\pm.133}.$$444_{\pm.089}.$$500_{\pm.109}.$
+Overall $1482$$386_{\pm.057}.$$484_{\pm.072}.$$415_{\pm.074}.$$424_{\pm.067}.$$426_{\pm.087}.$$461_{\pm.086}.$
+Table A3: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
+PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
+For English Dataset
 Due to budgetary constraints, we only evaluated four scripts from the English dataset, with the performance results reported in Table A4.
@@
 This discrepancy could be attributed to language biases inherent in the training data utilised for these models.
-Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
+Script Evaluation #QA PP OP Agent’s Response After Playing the Game
 Werewolf O-CoT ThinkThrice PLAYER*
@@
 Table A4: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
-SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
+PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
+Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved.
+This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
+Script #Tokens Stage PP OP Agent’s Response After Playing the Game
+Werewolf O-CoT ThinkThrice PLAYER*
+Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $3.363$
+Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
+Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $5.947$
+Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
+Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $4.603$
+Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
+Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $3.505$
+Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
+Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.846$
+Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
+Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
+Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
+Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.596$
+Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
+Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $1.150$
+Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
+Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.701$
+Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
+Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.561$
+Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
+Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $6.800$
+Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
+Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $5.028$
+Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
+Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
+Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
+Table A5: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
+#$\#Tokens$ represent the average length of each character’s script.
+Costs are reported for one complete gameplay and one evaluation process for each script.
 C.4 Experiments with Open-Source LLMs
@@
 Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data.
-The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 3.3 chapter.
+The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 4.3 chapter.
 For Civilian Players:
@@
 "answer": "a,b"
-Single-Choice Question Template for FSA Prompt
-This template is specifically designed for Full Script Access (FSA), distinguishing it from conventional templates by granting access to the scripts of all players.
+Single-Choice Question Template for OP Prompt
+This template is specifically designed for Omniscient Perspective (OP), distinguishing it from conventional templates by granting access to the scripts of all players.
 Unlike standard procedures, it leverages the RAG technique to extract pertinent information from each player’s script for comprehensive analysis.
@@
 "answer": "a,b"
-Multiple-Choice Question Template for FSA Prompt
-Similar to the single-choice version, this template is tailored for answering multiple-choice questions under FSA conditions.
+Multiple-Choice Question Template for OP Prompt
+Similar to the single-choice version, this template is tailored for answering multiple-choice questions under OP conditions.
 Please answer the questions based on the information in each characters script:

### v3

@@
-PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games
+Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games
 Abstract
-We propose PLAYER*, a novel framework that addresses the limitations of existing agent-based approaches built on Large Language Models (LLMs) in handling complex questions and understanding interpersonal relationships in dynamic environments.
-PLAYER* enhances path planning in Murder Mystery Games (MMGs) using an anytime sampling-based planner and a questioning-driven search framework.
-By equipping agents with a set of sensors, PLAYER* eliminates the need for pre-defined questions and enables agents to navigate complex social interactions.
-We additionally make a contribution by introducing a quantifiable evaluation method using multiple-choice questions and present WellPlay, a dataset containing 1,482 question-answer pairs.
-Experimental results demonstrate PLAYER*’s superiority over existing multi-agent methods, enhancing the generalisability and adaptability of agents in MMGs and paving the way for more effective multi-agent interactions.
-PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games
+We present Questum, a novel framework for Large Language Model (LLM)-based agents in Murder Mystery Games (MMGs).
+MMGs pose unique challenges, including undefined state spaces, absent intermediate rewards, and the need for strategic interaction in a continuous language domain.
+Questum addresses these complexities through a sensor-based representation of agent states, a question-targeting mechanism guided by information gain, and a pruning strategy to refine suspect lists and enhance decision-making efficiency.
+To enable systematic evaluation, we propose WellPlay, a dataset comprising 1,482 inferential questions across 12 games, categorised into objectives, reasoning, and relationships.
+Experiments demonstrate Questum’s capacity to achieve superior performance in reasoning accuracy and efficiency compared to existing approaches, while also significantly improving the quality of agent-human interactions in MMGs.
+This study advances the development of reasoning agents for complex social and interactive scenarios.
+Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games
 Qinglin Zhu1∗, Runcong Zhao1∗, Jinhua Du2, Lin Gui1, Yulan He1,3 1King’s College London, 2Huawei London Research Centre, 3The Alan Turing Institute {qinglin.1.zhu, runcong.zhao}@kcl.ac.uk{jinhua.du}@huawei.com {lin.1.gui,yulan.he}@kcl.ac.uk
@@
 Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023).
-Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023), deception (Wang et al., 2023), and the spread of information (Park et al., 2023).
-Despite these advances, building agents that intentionally interact with humans in dynamic environments remains challenging due to several limitations.
-First, current state-of-the-art approaches struggle with complex questions that lack straightforward answers and have difficulty in understanding interpersonal relationships.
-As a result, in addition to the game rules, agents often require additional human-written guides tailored to specific scenarios, relying on human context understanding (Xu et al., 2023; Wang et al., 2023).
-This limitation highlights the need for significant manual efforts and reduces the generalisability of these approaches across different domains and situations.
-Second, applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios is challenging because of the difficulties in defining state spaces, action spaces, and rewards.
-The reward associated with the final decision (e.g., identifying the murderer) often represents the only clear certainty (win or lose) (FAIR et al., 2022; Shi et al., 2024).
-Third, current evaluation methods are difficult to quantify.
-These methods rely heavily on manual evaluation to gain a deep understanding, but they are considerably influenced by subjectivity (Xu et al., 2023; Wang et al., 2023).
-Figure 1: The murder mystery game unfolds as players take on roles, question suspects based on their character’s objectives, and use deduction to identify the killer, while achieving their ideal role-specific ending.
-The complete game rules can be found in the appendix A.
-To address the limitation of lacking generalisability, we propose a novel framework that efficiently optimises path planning in Murder Mystery Games (MMGs), a strategic game involving both cooperation and competition among 4-12 players through natural language negotiation and tactical coordination (as illustrated in Figure 1).
-Our approach eliminates the need for pre-defined questions or instructions, allowing agents to play various scripts without extra human-written guides, just like human players.
-This framework enables agents to navigate complex social interactions and make decisions by considering the types of sensors they need.
-In the case of MMGs, this involves a set of sensors that detect emotions between players, motivations, and opportunities for actions like murder.
-To address the difficulty in defining state spaces and rewards, we designed an anytime sampling-based planner instead of using MARL.
-Unlike informed graph-based search methods, such as A* (Dechter and Pearl, 1985), our planner, inspired by algorithms like RRT* (Karaman and Frazzoli, 2011) and BIT* (Gammell et al., 2020), samples possible states with detective sensors in real-time.
-The tree-style search path is constructed and pruned dynamically until the target is reached.
-For MMGs, we combine sensors with a pruner that focuses on extracting highly suspicious murders to construct a purely questioning-driven search framework.
-PLAYER* is guided by sensors to efficiently generate high-quality questions to uncover more clues, and uses a pruner to more precisely identify the killer.
-To address the evaluation problem, we propose more quantifiable and reproducible evaluation criteria by devising a series of multiple-choice questions.
-This approach encompasses three types of questions: (1) Objective questions about the shared and individual objectives of each character in the game, (2) Reasoning questions probing the reasoning behind the given responses, and (3) Relations questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c).
-In summary, we have made the following contributions:
-(1) We propose PLAYER*, a framework that enhances the generalisability and adaptability of agents by efficiently optimising path planning in MMGs.
-(2) We designed an anytime sampling-based planner to dynamically construct and prune the search path, enhancing performance and efficiency.
-(3) We introduce a quantifiable evaluation method using multiple-choice questions focusing on character objectives, reasoning, and facts, thereby reducing the subjectivity in current evaluation practices.
-We also constructed and annotated a corresponding dataset, WellPlay, which includes 1,482 QA pairs.
-[^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
-2 PLAYER*
-In this section, we describe the MMG settings and provide an overview of how PLAYER* works.
+Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023).
+Despite these advances, building agents for Murder Mystery Games (MMGs) that intentionally interact with humans in dynamic environments remains challenging.
+Figure 1: The murder mystery game involves players taking on roles, questioning suspects, and using deduction to identify the killer, all while pursuing their character-specific objectives.
+As shown in Figure 1, MMGs are strategic games that involve 4-12 players assuming character roles with specific scripts and objectives.
+Gameplay unfolds through language-based negotiation and tactical coordination over $n$ rounds, with players limited to $m$ questions per round.
+The ultimate goal is to uncover the truth collaboratively while advancing individual objectives.
+At the end of the game, players vote on the suspected killer and evaluate their objective completion.
+MMGs are script-based games in which each script can only be played once.
+Unlike traditional games such as Go, which have fixed settings and well-defined action spaces for searching, or Werewolves (e.g., voting results) and Avalon (e.g., mission success), which have intermediate rewards, MMGs lack both well-defined searching spaces and intermediate rewards in decision making.
+To address these challenges, we first propose using a discrete space, which is called "sensor" to model each character to character states.
+Since it can be captured or updated by prompting the natural language description, we are able to explore this search space using generative Question-Answering (what to ask) between agents in real time as a regular planning problem (Karaman and Frazzoli, 2011; Gammell et al., 2020).
+For the issue of intermediate reward in searching, we design an Information Gain (IG) based heuristic objective which aims to balance existing searching results and the expected reward for further questions.
+After each action, the agent evaluates whether the current information can update the character states and narrow down the range of suspects.
+Meanwhile, our method focuses on extracting highly suspicious characters, creating a question-target selection framework guided by IG (who to ask).
+Compared to previous work, this framework optimises the information collection path, significantly improving system efficiency.
+Furthermore, current evaluation methods either focus solely on the final win/loss outcome or rely heavily on manual evaluation, which significantly limits the scope of analysis due to the high-cost (Xu et al., 2023a; Wang et al., 2023; Wu et al., 2024).
+For instance, Wu et al.
+(2024) manually designed only 56 inferential questions for 4 selected MMGs.
+To address the evaluation problem, we constructed a dataset, WellPlay, containing 1,482 inferential questions across 12 MMGs.
+These questions are categorised into three types: objective, reasoning, and relation.
+Including the win rate metric used in previous works, this provides a more diverse evaluation framework.
+We also recruited human players to interact with the agent.
+The results revealed that previous agents often overfit to agent-vs-agent interaction environments, exhibiting verbose and repetitive dialogues that detracted from the player experience.
+By integrating our goal-oriented pruner, these issues were significantly mitigated, resulting in more human-like agent behaviour and greatly enhancing player satisfaction.
+In summary, we have made the following contributions: (1) We propose Questum, a framework using detective sensors to define and optimise the search in continuous language space.
+(2) We introduce an IG-based reward to address the lack of intermediate rewards, improving agent behaviour and efficiency.
+(3) We construct WellPlay, a dataset with 1,482 inferential questions across diverse categories, offering a more robust evaluation framework.
+(4) By integrating a goal-oriented pruner, we enhance agent-human interaction, reducing verbosity and improving player experience.
+[^1] Our code and dataset are available at https://github.com/alickzhu/Questum.
+2 Questum
+In this section, we describe the MMG settings and provide an overview of how Questum works.
 2.1 Problem Setting
-Figure 2: Problem Setting.
-Figure 3: PLAYER* begins with an introduction round, where each agent provides a brief self-introduction.
-Following this, agents initialise scenario understanding and begin to search for further information by questioning other agents.
-We use arrows to identify their questioning actions, with blue arrows representing questions asked by the agent itself and yellow arrows representing questions asked by other agents, which are also remembered by the given agent.
-Based on the responses they receive, agents update their suspicion state, identified by the black arrow (which decides who to ask), and the corresponding action space, identified by the pink area (which decides what questions to ask).
-The iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
-In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored for such scenarios.
-As illustrated in Figure 2, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game with $N_{a}$ playable characters and a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$.
-Each agent $a_{i}$ is assigned to a character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$.
-The script $C_{i}$ is crafted from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective.
-The suspicious state $s_{ik}$ is represented as an $(N_{a}-1)$-dimensional vector that encodes agent $a_{i}$’s beliefs about other players’ suspicions regarding the murder of victim $v_{k}$.
-The vector entries can be either discrete ($\in{\{0,1}\}$) or continuous values ($\in[0,1]$), reflecting the strength of the suspicions.
-Unlike traditional planning tasks in which the state space can be accessed directly, $s_{i}$ in our scenario is generated from a language model.
-We define the possible questions as an action space, allowing the derivation of suspicious states through an LLM-based search.
-For an illustrative example, refer to Appendix A.3.
-As shown in Figure 3, each agent is provided with character-specific information, enabling them to interact with others based on their scripts and historical interactions within the game.
-Subsequently, agents engage in simulated discussions by asking questions and expressing opinions.
-To ensure fairness, each agent is limited to asking $m$ questions per round over $n$ rounds.
-After the discussion, agents must collectively decide and vote on the suspected killer.
-2.2 Construction of Agents
-To enable agents to think and act like human players, we equip them with the following components: (1) Game Rules.
-The essential information for both human and AI players.
-(2) Chain-of-Thought.
-Facilitating a sequential reasoning process to aid in problem-solving.
-(3) Sensors.
-In sociology, interpersonal relationships are often conceptualised within a multidimensional space encompassing factors like time, distance, and various other elements (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010).
-Drawing inspiration from this concept, we propose employing a set of sensors associated with each task.
-For example, in a murder mystery game, these factors could include emotion, motivation, and suspicion.
-(4) Memory Retrieval.
-To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement this module to store and retrieve narrative scripts and dialog logs generated during gameplay.
-Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent.
-When an agent encounters a new event requiring action, we use the Faiss library (Douze et al., 2024) to retrieve relevant memories.
-This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length.
-It ensures consistent and coherent interactions among agents while assisting in their strategic planning throughout the game.
-(5) Self-Reflection.
-Preserving historical experience by reflecting on the current situation to provide a more detailed, accurate, and profound answer.
-2.3 PLAYER* Planning Strategy
-To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions.
-As illustrated in Figure 4, this framework is fundamentally composed of two key components:
+In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored to such scenarios.
+This framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ and a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{a}$ and $N_{v}$ denote the numbers of playable characters and victims, respectively.
+Each agent $a_{i}$ is assigned to a character and initialised with the following:
+(1) Role $\mathbf{r}_{i}=\{r_{ik}\}_{k=1}^{N_{v}}$: Whether or not they are the murderer of each victim;
+(2) Role background script $C_{i}$: Crafted from the unique viewpoint of $a_{i}$;
+(3) Objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{o_{i}}}$: A set of $N_{o_{i}}$ goals for agent $a_{i}$.
+Game Rules are also provided to Agents as essential information. The detailed rules and procedures are provided in Appendix A.
+2.2 Questum  Planning Strategy
 Figure 2: Search and Approximate.
-PLAYER* generates questions based on sensor-detected values and identifies the killer more precisely with a prunner.
-This dynamic question generation allows agents to adjust to the specific circumstances of each game session, ensuring a more focused and effective investigation.
-Questioning Strategy in Searching based on Sensors
-Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components.
-Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023), we generate actions directed towards the targets based on the values detected by those sensors.
-The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors.
-For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Opportunity}]$: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible.
-Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator.
-Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime.
-The sensor values are obtained based on the related information retrieved from the script and dialog, represented by the function $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$, where $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts and $D_{i}$ denotes the associated dialogue history.
-Then, for agent $a_{i}$ to be questioned, PLAYER* uses LLMs to sample the question $q$ from the action space given the values of sensors $S_{i}$ and related information $(C^{\prime}_{i},D_{i})$, represented by the function $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
-Action Space Refinement with a Pruner
-Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents.
-Each agent first updates its understanding of interpersonal relationships, thereby adjusting sensor values $S$.
-Subsequently, it also updated $P_{informative}$ based on the revised information and changes in sensor values.
-Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration.
-The reward for questioning is the information that can be obtained from the anticipated response.
-As the number of inquiries directed at an agent increases, the assumed exploration within the world space relevant to that agent also rises.
-Consequently, the expected reward associated with questioning the same agent diminishes proportionally to the number of inquiries already made.
-Therefore, it becomes essential to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if they appear highly suspicious.
-We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ concerning the agent to be questioned, and the number of inquiries already posed $|Q_{i}|$.
-In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals.
-This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning.
-Emulating the cognitive pattern of human beings, we refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets (See Lower Part of Figure 4).
-We did not impose a hard restriction on the refinement process and allow for flexible adjustment of the suspicious state.
-For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked.
-This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility.
-Then we update the action space based on the suspicious state.
+Questum  generates questions based on character states, selecting agents to question based on past observations of critical information and the likelihood of uncovering more.
+The goal is to minimise the suspect list.
+Questum approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions.
+As illustrated in Figure 2, this framework is fundamentally composed of two key components: (1) Search based on sensors: Questum  searches for the murderer based on their proximity to the ideal murderer in the language space.
+(2) Approximation with a pruner: Questum  focuses on a subset of characters that are highly suspicious of being a murderer, and decides whom to query based on this suspect list.
+In MMGs, the suspect status is $\mathbf{s_{i}}=\left[s_{ijk}\right]_{j=1,k=1,j\neq i}^{N_{a},N_{v}}$, where $s_{ijk}\in[0,1]$.
+Here, $s_{ijk}$ encodes agent $a_{i}$’s beliefs about player $a_{j}$’s likelihood of being victim $v_{k}$’s murderer.
+In our setting, $s_{ijk}$ is set to be discrete ($\in\{0,1\}$), representing suspect and non-suspect.
+While our list of suspects is finite, our actual search space $L\in\mathbb{R}^{d}$ (the language space) is continuous and high-dimensional, where $d$ is potentially very large.
+It encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets.
+Search based on sensors
+We map each suspect to a point in the state space $L$.
+Our search process operates continuously in the state space $L$, where character attributes such as emotions, motivations, and opportunities are represented as vectors $u_{i}=(e_{i},m_{i},p_{i},...)$ in this multi-dimensional space.
+Character attributes are chosen with domain knowledge, drawing on insights from sociology (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010).
+We use sensors to sample the values of these attributes for two main purposes:
+1) Estimating the distance between a given suspect and the hypothetical murderer.
+The hypothetical murderer is assumed to be a point in the search space.
+At the beginning of the game, when participants lack additional information, the hypothetical murderer refers to an idealised profile that meets specific criteria across all sensor dimensions: someone with motivation, negative emotions, and the opportunity to commit the crime.
+As the game progresses and more available information, agents update the criteria defining the hypothetical murderer.
+For instance, if it is suddenly revealed that the victim’s death was accidental, the murderer may no longer be expected to have a motive.
+2) Exploration by generating questions.
+It is modelled as $q=g(u_{i},X)$, where $q$ is a newly generated question by language model $g$.
+In the inputs, $X$ is the relevant information retrieved from the context by Retrieval-Augmented Generation (RAG) and $u_{i}$ is the state representation of suspect $i$.
+Approximation with a pruner
+Following the questioning step, agents receive responses from the questioned agent as well as conversational exchanges among other agents.
+Each agent then updates its understanding of interpersonal relationships, thereby updating character states.
+In human gameplay, players often form judgments regarding the scenario and focus on a subset of highly suspicious individuals.
+This strategic refinement process is crucial for conserving cognitive resources and enhancing the efficacy of inferential reasoning.
+Inspired by the human thinking process, we also enable agents to consider the suspect list.
+In an MMG with $N_{a}$ players, to find out who is the murderer $Y$ of a given victim $v_{k}$, we define the entropy to be $H=-ln(\frac{1}{n})$, where $n$ is the number of suspects in the agent’s suspect list.
+Based on the size of the suspect list for each round of the game, the agent can calculate the corresponding entropy.
+The target is to maximise this entropy.
+The initial entropy is $H_{0}=-ln(\frac{1}{N_{a}-1})$, where everyone except the agent itself, who aims to find the murderers, are all suspects, representing the uncertainty about who the murderer is at the start of the game.
+For each character, we have its historical IG being $IG_{i}=H_{i-1}-H_{i}$ for round $i$, which represents the reduction in uncertainty about the murderer’s identity after receiving the sensor reading and the corresponding answer.
+$IG_{1:i}=[IG_{1},IG_{2},...,IG_{i}]$.
+The weighted historical IG is:
+$IG_{weighted}[i]=\frac{\sum_{k}w_{k}\times IG_{k}}{\sum_{k}w_{k}}$ (1)
+where $w_{k}$ are weights, set to $e^{-(i-k)}$.
+After the search step, the agent acquires more information.
+However, accessing the conditional probability to update the conditional entropy $H_{i+1}=H(Y|u_{i},X)$ might be challenging in closed-source platforms like ChatGPT, and open-source projects such as Llama could introduce strong bias as they usually focus on one suspect Xu et al.
+(2023b).
+Hence, we propose a heuristic estimator based on prompting.
+First, we ask the LLMs to estimate the expected IG.
+We then convert the discrete value from the prompt to a numerical value:
+$E[IG_{i+1}]=\begin{cases}1,&\text{if prompt returns ``high''}\\ 0,&\text{if prompt returns ``medium''}\\ -1,&\text{if prompt returns ``low''}\end{cases}$ (2)
+Thus, even closed-source platforms are able to estimate conditional entropy, and the bias in open-source models is reduced by smoothing in the form of multiple-choice questions.
+We then combine historical and expected gain to determine the highest-scoring character to ask:
+$Score_{i}=\beta IG_{weighted}[i]+(1-\beta)E[IG_{i+1}]$ (3)
+In addition, we also implement an $\epsilon$-greedy strategy to choose a random character with probability $\epsilon$, and the highest-scoring character with probability $1-\epsilon$ to balance exploration and exploitation.
 This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads.
-By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
-Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, Suspicious States $s$, Max Round $n$
+Thus, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and increasing the likelihood of identifying the perpetrator within the designated number of discussion rounds.
+The detailed procedure is in Algorithm 1.
+Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, Suspicious States $s$, Max round $n$.
 Output: Evaluation of Results
 1 current_round = 0
-2 while current_round < $n$ do
-       // Action Generation
+2 while current_round $\leq$ $n$ do
+       // Search based on sensors
 3       for i = 1 to $N_{a}$ do
@@
 6                   for $a_{j}$ in $suspect\_list_{ik}$ do
-7                         question = Action_Generation($a_{i},a_{j},v_{i}$)
+7                         question = Action_Generation($a_{i},a_{j},v_{k}$)
 8                         answer = Reply($a_{i}$)
@@
 11
-      // Update the suspect list
+      // Approximation with a pruner
 12       for i = 1 to $N_{a}$ do
 13             for k = 1 to $N_{v}$ do
-14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{i}$)
+14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{k}$)
 15
@@
 17      current_round += 1
-// Evaluation
-18 for i = 1 to $N_{a}$ do
-19       Result = Evaluate($a_{i}$)
-20
-Algorithm 1 Anytime Sampling-based Planning in Multi-agent Interaction
-Algorithm 1 outlines a detailed procedure of the game process.
-While initially designed for MMGs, the framework’s adaptable nature allows it to be readily applied to various games by adjusting the game rules and task-specific sensors, thereby enhancing the gaming experience and offering a structured way for agents to handle complex social dynamics efficiently.
+Algorithm 1 Questum Framework
 3 The WellPlay Dataset
-Script Agents Victims #token(CN) #token(EN) Question
+MMG #Agents #Victims #token(CN) #token(EN) Question
 avg overall avg overall Objective Reasoning Relations overall
@@
 Table 1: Dataset Statistics.
-Agents is the count of players, Victims is the number of script victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively.
+Agents is the count of players, Victims is the number of victims, #token(CN) and #token(EN) are the token counts in the Chinese and English dataset versions, respectively.
 Avg shows the average script length per character, Overall is the total script token count, and Question enumerates the number of questions by types.
 The number of evaluation questions varies based on script complexity, with more complex scripts generating a larger volume of questions.
-WellPlay Dataset
-We built an evaluation dataset, called WellPlay, from the existing Conan dataset (Zhao et al., 2024c), derived from background narratives created for MMGs.
-Conan labels relationships from each character’s perspective, along with ground truth relationships, consolidating all available information.
+Dataset
+We built an evaluation dataset, WellPlay, derived from background narratives created for MMGs (Zhao et al., 2024c).
 To establish a quantifiable evaluation method, we employ multiple-choice questions focusing on factual information to reduce controversial answers.
-Besides existing labels of interpersonal relationships that gauge agents’ understanding of the situation, we also recruited four annotators to label character objectives and reasoning.
-The dataset encompasses 12 MMGs, comprising a total of 1,482 evaluation questions.
-On average, each game features 5.67 agents and 1.75 victims as detailed in Table 1 .
-Our evaluation questions are grounded in characters’ objectives in the game and are crafted with input from human players, focusing on the following question types:
-(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles wallet, for each character in the game.
+We have recruited four annotators to label inferential questions on:
+(1) Objective. Including shared objectives, such as identifying the perpetrator(s), and individual objectives, such as determining who stoles the wallet, for each character in the game.
 (2) Reasoning.
-This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, disappearance, explosion); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, financial conflicts, manslaughter).
-(3) Relations. This includes interpersonal relationships between victims and other characters, as well as relationships among suspects.
-Evaluation Metrics
-We report the performance of each type of question and also calculate the agent’s final score as Awarded PointsTotal $\frac{\text{Awarded Points}}{\text{Total Points}}$ to determine its overall performance.
-The awarded points follow the scoring system used in MMGs, which awards 10 points for objective questions, 5 points for reasoning questions, and 2 points for relation questions.
-For these calculations, we utilise weighted mean and weighted standard deviation methodologies.
-In addition to assessing the accuracy of the three types of questions and the overall score, we also evaluate the Win Rates of players for performance assessment.
-The game rules are as follows: every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) will be eliminated.
-Additionally, we assume that murderers would always vote for someone else instead of voting for themselves.
-If the true murderer is successfully voted out, the game is considered a victory for the players seeking to identify the culprit.
-The detailed about the dataset and calculations is provided in Appendix B.
+This entails questions that delve into the reasoning behind provided answers, relating to agents’ objectives, including: Who; What (the nature of the incident, such as murder, theft, or disappearance); When (the time of the incident); Where(the location of the incident); Cause (e.g., shooting, poisoning, stabbing); Motive (e.g., crime of passion, vendetta, or manslaughter).
+(3) Relations. This includes interpersonal relationships between victims and others, as well as relationships among suspects, with labels adapted from the Conan dataset (Zhao et al., 2024c).
+WellPlay encompasses 12 MMGs, comprising a total of 1,482 evaluation questions.
+On average, each game features 5.67 agents and 1.75 victims (see Table 1).
+More details of the dataset and annotating schema are in Appendix B.
 4 Experiments
+Script Evaluation Werewolf O-CoT ThinkThrice Questum
+Death Wears White (9 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective .$033_{\pm.047}.$$067_{\pm.047}.$$067_{\pm.047}.$$033_{\pm.047}.$
+Reasoning .$258_{\pm.024}.$$324_{\pm.014}.$$327_{\pm.012}.$$356_{\pm.030}.$
+Relations .$421_{\pm.024}.$$458_{\pm.041}.$$384_{\pm.026}.$$435_{\pm.017}.$
+Ghost Revenge (7 players, 3 victims) Win Rate .$222_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$333_{\pm.000}.$
+Objective .$193_{\pm.066}.$$158_{\pm.043}.$$193_{\pm.025}.$$333_{\pm.066}.$
+Reasoning .$307_{\pm.008}.$$322_{\pm.019}.$$377_{\pm.022}.$$353_{\pm.016}.$
+Relations .$314_{\pm.018}.$$295_{\pm.018}.$$353_{\pm.025}.$$353_{\pm.018}.$
+Danshui Villa (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective .$083_{\pm.000}.$$111_{\pm.039}.$$194_{\pm.039}.$$111_{\pm.079}.$
+Reasoning .$286_{\pm.027}.$$286_{\pm.019}.$$310_{\pm.004}.$$286_{\pm.004}.$
+Relations .$312_{\pm.027}.$$365_{\pm.045}.$$259_{\pm.015}.$$296_{\pm.075}.$
+Unfinished Love (7 players, 2 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$500_{\pm.000}.$
+Objective .$083_{\pm.118}.$$028_{\pm.039}.$$000_{\pm.000}.$$333_{\pm.068}.$
+Reasoning .$443_{\pm.023}.$$426_{\pm.035}.$$481_{\pm.039}.$$536_{\pm.028}.$
+Relations .$472_{\pm.034}.$$519_{\pm.036}.$$514_{\pm.020}.$$560_{\pm.007}.$
+Cruise Incident (5 players, 1 victim) Win Rate .$667_{\pm.471}.$ $1.000_{\pm.000}$$667_{\pm.471}.$ $1.000_{\pm.000}$
+Objective .$417_{\pm.118}.$$500_{\pm.000}.$$583_{\pm.312}.$$667_{\pm.236}.$
+Reasoning .$458_{\pm.068}.$$444_{\pm.086}.$$458_{\pm.059}.$$528_{\pm.052}.$
+Relations .$367_{\pm.082}.$$422_{\pm.042}.$$411_{\pm.016}.$$422_{\pm.016}.$
+Sin (4 players, 1 victim) Win Rate .$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$
+Objective .$333_{\pm.272}.$$000_{\pm.000}.$$000_{\pm.000}.$$444_{\pm.314}.$
+Reasoning .$650_{\pm.108}.$$467_{\pm.024}.$$533_{\pm.047}.$$550_{\pm.041}.$
+Relations .$333_{\pm.067}.$$571_{\pm.067}.$$413_{\pm.090}.$$492_{\pm.022}.$
+Deadly Fountain (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Reasoning .$381_{\pm.000}.$$444_{\pm.098}.$$508_{\pm.022}.$$587_{\pm.022}.$
+Relations .$250_{\pm.068}.$$194_{\pm.039}.$$222_{\pm.039}.$$333_{\pm.068}.$
+Unbelievable Incident (5 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective .$083_{\pm.118}.$$000_{\pm.000}.$$000_{\pm.000}.$$083_{\pm.118}.$
+Reasoning .$431_{\pm.052}.$$278_{\pm.071}.$$194_{\pm.071}.$$472_{\pm.020}.$
+Relations .$622_{\pm.083}.$$733_{\pm.109}.$$289_{\pm.126}.$$533_{\pm.054}.$
+Desperate Sunshine (4 players, 1 victim) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
+Objective .$333_{\pm.000}.$$111_{\pm.157}.$$556_{\pm.157}.$$778_{\pm.157}.$
+Reasoning .$537_{\pm.026}.$$630_{\pm.094}.$$759_{\pm.052}.$$741_{\pm.069}.$
+Relations .$491_{\pm.094}.$$574_{\pm.035}.$$491_{\pm.047}.$$556_{\pm.068}.$
+Riverside Inn (4 players, 1 victim) Win Rate .$083_{\pm.118}.$$000_{\pm.000}.$$083_{\pm.118}.$$250_{\pm.204}.$
+Objective .$117_{\pm.024}.$$117_{\pm.062}.$$117_{\pm.094}.$$383_{\pm.062}.$
+Reasoning .$248_{\pm.013}.$$312_{\pm.022}.$$281_{\pm.011}.$$373_{\pm.017}.$
+Relations .$531_{\pm.014}.$$541_{\pm.038}.$$483_{\pm.030}.$$589_{\pm.071}.$
+Solitary Boat Firefly (6 players, 4 victims) Win Rate .$222_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective .$250_{\pm.090}.$$181_{\pm.039}.$$181_{\pm.039}.$$250_{\pm.034}.$
+Reasoning .$409_{\pm.004}.$$369_{\pm.044}.$$453_{\pm.008}.$$539_{\pm.027}.$
+Relations .$473_{\pm.028}.$$568_{\pm.048}.$$439_{\pm.033}.$$542_{\pm.027}.$
+Manna (6 players, 3 victims) Win Rate .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$333_{\pm.471}.$
+Objective .$111_{\pm.157}.$$000_{\pm.000}.$$556_{\pm.157}.$$444_{\pm.157}.$
+Reasoning .$463_{\pm.052}.$$426_{\pm.026}.$$593_{\pm.026}.$$648_{\pm.069}.$
+Relations .$444_{\pm.045}.$$407_{\pm.069}.$$333_{\pm.045}.$$444_{\pm.045}.$
+Overall Win Rate .$127_{\pm.059}.$$063_{\pm.022}.$$111_{\pm.045}.$$222^{*}_{\pm.081}.$
+Objective .$160_{\pm.020}.$$123_{\pm.008}.$$162_{\pm.024}.$$288^{*}_{\pm.021}.$
+Reasoning .$343_{\pm.006}.$$349_{\pm.012}.$$384_{\pm.002}.$$423_{\pm.009}.$
+Relations .$425_{\pm.008}.$$473_{\pm.010}.$$405_{\pm.006}.$$471_{\pm.009}.$
+Overall .$324_{\pm.004}.$$329_{\pm.008}.$$347_{\pm.006}.$$407^{*}_{\pm.006}.$
+Table 2: Results of Agent-vs-Agent Evaluation.
+An “*” indicates statistical significance under the two-sample t-test with a significance level of $\alpha=0.05$, comparing it with the second-best model.
+More detailed data is provided in Table 8.
 4.1 Experimental Setup
-Base Models
-We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models in Jan-Mar 2024, using the default model versions gpt-35-turbo-16k 0613 (Default) and text-embedding-ada-002.
-To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and variance.
-The detailed introduction to the experiment settings is in Appendix C.1.
-Baselines
+We conducted experiments with GPT-3.5 for conversation and the GPT Embedding Model for memory retrieval via the Azure API  [^2] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/models , using gpt-35-turbo-16k 0613 and text-embedding-ada-002.
+To minimise randomness, we conducted the evaluation experiments 3 times and reported the average and standard deviation.
+In this paper, we focus on the strategy for the good camp.
+The murderer agent’s questioning strategy mirrors that of other agents, with an added prompt instructing it to act as if it is not the murderer while questioning others.
+4.2 Evaluation Metrics
+For evaluation, we adopt the following metrics:
+Win Rate:
+Every player, including the murderer, can vote, and the player with the majority number of votes ($\geq 50\%$) is eliminated.
+Murderers always vote for others instead of themselves.
+If the true murderer is voted out, the game is considered a victory for the identifying players.
+Question Accuracy:
+The accuracy of agents’ answers to three types of questions—Objective, Reasoning, and Relations—and Overall, where Overall represents the combined accuracy across all question types to provide a comprehensive evaluation of the agents’ performance.
+This metric is used in both Agent-vs-Agent and Human-vs-Agent evaluations to measure agents’ ability to answer inferential questions correctly.
+4.3 Baselines
 For baselines, we compare our approach with other multi-agent algorithms designed for multiplayer deduction games.
 Although some methods were not designed for MMGs, they are the most relevant and adaptable frameworks in this relatively new area of LLM-based multi-agent games.
-We adapted their methodology to MMGs for a fair comparison, with implementation details in Appendix C.2.
-(1) Werewolf (Xu et al., 2023).
+(1) Werewolf (Xu et al., 2023a).
 Werewolf is another multi-agent game, where players identify werewolves through group discussion.
@@
 Designed for MMGs, agents craft questions from retrieved memory and the current scenario.
-(4) Personal Perspective.
+(4) Personal Perspective (PP).
 For a more comprehensive comparison, we also assess the performance of agents who do not actively participate in the game but make their final decisions only based on their script.
-(5) Omniscient Perspective.
+(5) Omniscient Perspective (OP).
 Agents do not actively participate but make their final decisions based on all agents’ scripts.
-To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and PLAYER*).
-This allows us to fairly compare the performance of the side trying to find the murderer, who face the same level of opposition from the murderer’s side.
-4.2 Results
-Performance Evaluation Results
+To account for the zero-sum nature of MMGs, where it would be hard to tell if agents can identify the murderer due to their good performance or the poor performance of their competitor, we fix the murderer’s framework in all experiments (including baselines and Questum).
+The implementation details, model comparison, information on sensors, method to calculate the Overall, and prompts are provided in Appendix C.
+4.4 Agent-vs-Agent Evaluation
 Figure 3: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
@@
 The Omniscient Perspective (OP) measures performance when agents have access to all agents’ scripts, representing the ideal search endpoint.
-Figure 5 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table A3).
-PLAYER* shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents.
-Notably, PLAYER* significantly outperforms others in objective questions, even surpassing OP, which has access to all information.
+In this section, we evaluate agent performance in an Agent-vs-Agent setting to compare their decision-making processes against competitive baselines.
+Figure 3 presents the average of agents’ performance across 12 unique games of varying complexity and settings as detailed in our results table (Table 2).
+Questum  shows superior performance compared to other baselines across all evaluation questions, demonstrating its enhanced understanding of the search space through interactions with other agents.
+Notably, Questum  significantly outperforms others in Objective Questions.
 The OP setting, where agents have access to the scripts of all agents without interacting with them, generally yields better performance compared to all other methodologies.
@@
 However, in practice, its effectiveness is often limited by the deductive capabilities of the underlying base model, which reflects the upper limit of the “approximate” ability of the given base model.
-This showcases the “approximate” ability of PLAYER* in refining information and dynamically narrowing down the search domain to achieve the target objective.
+This showcases the “approximate” ability of Questum  in refining information and dynamically narrowing down the search domain to achieve the target objective.
 Since the PP and OP settings do not involve active participation in the game, we consider them as indicators of the starting point and endpoint that can be achieved through search.
-The results demonstrate that PLAYER* outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability.
+The results demonstrate that Questum  outperforms all baseline agents in reasoning and relations, showcasing its superior “search” ability.
 The larger performance gap in objective questions for other baselines reveals a critical limitation: an inability to effectively utilise the collected information to reach correct conclusions or achieve game-specific goals.
 This highlights the importance of not only gathering relevant information but also efficiently processing and applying it to attain the desired objectives.
-Our findings reveal a significant discrepancy between the win rate and the objective metric (the average of non-murderer players correctly identifying the murderer) for agents, with OP exhibiting a small difference and PP showing the largest difference.
-The case analysis unveils that in OP games, agents make similar decisions, resulting in voting outcomes that are either mostly correct or mostly wrong, leading to consistency between individual and collective voting results.
-Conversely, in PP games, agents base their decisions on their individual scripts, leading to independent actions and difficulty in reaching a consensus to vote out the murderer.
-The performance of the agents falls between that of OP and PP, with vote distributions more dispersed than OP but more concentrated than PP.
-O-CoT suffered from unfortunate circumstances, falling short by a single vote in numerous games.
-This resulted in O-CoT being on the verge of victory in many instances but ultimately yielding poor overall results.
-We also evaluate the performance of our framework on English scripts in Appendix C.3 and open-source LLMs in Appendix C.4.
+We also evaluate the performance of our framework on English scripts and open-source LLMs in Appendix D.
 The detailed dialogue history and evaluation records are available in the GitHub link provided previously.
-Efficiency and Cost Comparison
-As shown in Figure 6, we delve into the efficiency and cost analysis across various methodologies implemented for agent interaction in MMGs, as detailed in our results table (Table A5).
-The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense incurred during both gameplay and evaluation stages.
+4.5 Agent-vs-Human Evaluation
+In the Agent-vs-Human Evaluation, we identified significant limitations of traditional evaluation methods used in agent for games, which were primarily designed for Agent-vs-Agent Evaluation and focused solely on achieving victory.
+While effective in assessing agents’ strategic and reasoning abilities, these methods fall short when considering human players as the end-users of the system.
+For instance, previous approaches often resulted in repetitive and meaningless conversations, which human players found unengaging and boring.
+To illustrate these limitations, we provide specific examples in Appendix E, drawn from dialogues that occurred during games involving human participants.
+To address this, we extended our evaluation by incorporating a player-centric perspective.
+In addition to the performance metrics used in Agent-vs-Agent Evaluation, we distributed a survey to human participants to capture their gameplay experience and satisfaction levels.
+This allowed us to assess Questum’s suitability not only as a strategic agent but also as a companion for human players in interactive settings.
+The metrics include:
+Story Advancement: Measures the agent’s effectiveness in gathering relevant information and advancing the game’s plot in a meaningful way.
+Higher scores indicate that the interactions were productive and kept the game engaging.
+Question Quality: Evaluates the relevance and depth of the questions posed by the agent.
+This metric ensures that the questions contribute to the narrative and help uncover critical information, avoiding redundancy or triviality.
+Response Quality: Assesses the accuracy and informativeness of the agent’s responses. Higher scores indicate that the agent provided clear and valuable information to the human players.
+Response Speed: Represents how quickly the agent interacts, ensuring timely response for a smooth and engaging gameplay experience.
+Role Immersion: Assesses how well the agent embodies its character’s personality, emotions, and background to create a believable experience.
+Evaluation Werewolf O-CoT ThinkThrice Questum
+Win Rate .$111_{\pm.157}.$$333_{\pm.132}.$$333_{\pm.048}.$$667_{\pm.085}.$
+Objective .$222_{\pm.091}.$$333_{\pm.045}.$$296_{\pm.052}.$$556_{\pm.091}.$
+Reasoning .$678_{\pm.028}.$$644_{\pm.028}.$$684_{\pm.008}.$$723_{\pm.035}.$
+Relations .$745_{\pm.028}.$$797_{\pm.040}.$$745_{\pm.028}.$$843_{\pm.042}.$
+Overall .$608_{\pm.006}.$$619_{\pm.023}.$$625_{\pm.009}.$$717_{\pm.038}.$
+Table 3: Performance comparison of different algorithms against human players.
+Figure 4: Results of Agent-vs-Human Evaluation using Human-Centric metrics. The detailed data can be found in Table 6.
+All scores are based on discrete 1-5 scales, where higher scores are better.
+The specific scoring criteria can be found in Table 7.
+Details about the setups (e.g., recruitment of players, gameplay conditions, and tasks) can be found in Appendix B.5.
+As shown in Table 3 (results against human players) and Figure 4 (Human-Centric metrics), Questum  not only surpassed the baseline models in strategic and reasoning-oriented evaluation but also demonstrated superior performance in human-centric dimensions.
+These results indicate that Questum  effectively balances tactical prowess with narrative engagement, delivering more satisfying gameplay experiences for human participants.
+4.6 Efficiency and Cost Analysis
+As shown in Figure 5, we delve into the efficiency and cost analysis across various methodologies for agent interaction in MMGs.
+The costs are presented in actual monetary values (US Dollars) associated with the use of Azure API, providing a direct measure of the computational expense.
 [^3] Billing method details are available on the website https://azure.microsoft.com/en-gb/pricing/details/cognitive-services/openai-service/.
-Figure 6: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
-Costs are reported for one complete gameplay and one evaluation process for each script.
-The costs are divided into Gameplay and Evaluation phases.
-Notably, PP and OP methodologies do not incur costs during the Gameplay phase, as they were tested in a direct, non-interactive manner.
+Figure 5: Comparison of time (hours) and costs (\$) for calling OpenAI API across multi-agent algorithms in MMG settings.
 We can see that Werewolf is relatively efficient to run, as it is partly based on preset instructions, which saves steps.
 In comparison, O-CoT is the most expensive one, as it needs to think about what to do step by step based on its objectives.
-PLAYER* stands out for its cost-efficiency and performance.
-By employing the Action Space Refinement strategy, PLAYER* minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters.
+Questum  stands out for its cost-efficiency and performance.
+By employing the Action Space Refinement strategy, Questum  minimises unnecessary API calls, concentrating its investigative efforts on the most suspicious characters.
 This focus significantly reduces the computational resources required, thereby lowering the overall cost of operations.
-This strategic optimisation is evident across all scripts, with PLAYER* consistently registering lower costs in comparison to its counterparts during the Gameplay phase.
-In the Evaluation phase, the performance of these agents is assessed based on the WellPlay Dataset.
-Evaluation costs are similar across all methods due to a shared evaluation strategy, with variations in PP and OP costs arising from their differing levels of script access.
+This strategic optimisation is evident across all scripts, with Questum  consistently registering lower costs in comparison to its counterparts during the Gameplay phase.
+The trend in time efficiency mirrors that of the monetary cost, with similar patterns observed across the methodologies.
 4.7 Ablation Studies
-We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7.
-Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions.
-This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of information gained from conversations tends to converge.
-These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases.
-For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
-Figure 7: Comparison of agents’ behaviour across different numbers of rounds, where each agent can ask a specific number of questions (denoted as $m$).
+We conducted comprehensive ablation studies to evaluate the effectiveness of the key components of our approach.
+Removing the Pruner or Sensor modules adversely affected both performance and cost reduction, confirming their utility; results are provided in Table 12.
+Additionally, we experimented with different sensor configurations and numbers of rounds/questions, results are shown in Figure 7.
+Our analyses revealed that a carefully chosen subset of three sensors strikes a balance between performance gains and computational costs and that around three rounds of interaction with one question per round is near-optimal.
+These experimental setups, alongside full performance comparisons and in-depth evaluations, are presented in the Appendix D.3.
 5 Related Works
@@
 Multi-agent reinforcement learning marking significant progress in complex games (Lanctot et al., 2017; Perolat et al., 2022; Bakhtin et al., 2023).
-However, these methods often require extensive time and computational resources and lack linguistic communication capabilities.
-With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024).
+However, these methods often require extensive computational resources and lack linguistic communication capabilities.
+With the emergence of LLMs, there’s a shift of focus towards improving multi-agent language communication, evidenced by advancements in various games and scenarios, such as werewolf (Xu et al., 2023a), avalon (Wang et al., 2023; Shi et al., 2024), interactive narrative (Zhao et al., 2024b), MMGs (Wu et al., 2024), and survival games (Toy et al., 2024).
 Exemplified by AlphaGo (Silver et al., 2017), self-play learning frameworks (Fu et al., 2023; Chen et al., 2024) are proposed to improve LLMs’ performance.
-Compared to classic methods (Wang and Shen, 2024), agents based on LLMs are capable of inferencing across a broader range of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states.
-However, they have also been found to inherit biases that limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024).
+Compared to classic methods (Wang and Shen, 2024), LLMs based methods can perform inference across a wider variety of scenarios (Lin et al., 2023), even with some ability of theory of mind (Zhou et al., 2023) to infer other agent’s mental states.
+However, the inherit biases that can potentially limit their inferential abilities (Xie et al., 2023; Chuang et al., 2024; Wang et al., 2024).
 Works have also explored utilising LLMs as the environment (Zhang et al., 2024) or update actions (Zhao et al., 2024a) for agents.
@@
 Alignment through human feedback offers more consistent training compared to reinforcement learning (Liang et al., 2024), but obtaining this feedback can be expensive.
-To address this, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023b), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced.
-These approach was also adapted to search for optimal tools (Du et al., 2024), interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024).
-We were also inspired by stochastic search methods utilised for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
+Therefore, approaches like self-instruct (Wang et al., 2022; Liu et al., 2023), self-reflect (Yao et al., 2023), self-alignment (Sun et al., 2023; Li et al., 2024b), and few-shot planning (Song et al., 2023), have been introduced.
+These approaches was also adapted to search for optimal tools (Du et al., 2024), and interact with grounded environments (Ouyang and Li, 2023; Ismail et al., 2024).
+We were also inspired by stochastic search methods for robots in planning optimal strategies in complex environments (Gammell et al., 2020; Liang et al., 2024), shares many similarities with optimisation tasks for agents (Singh et al., 2023).
 6 Conclusion
-PLAYER* addresses limitations of LLM-based agents in complex reasoning using an anytime sampling-based planner with sensors and pruners.
-Our approach enables efficient, question-driven searching.
-We propose a quantifiable evaluation method, contribute the WellPlay dataset, and demonstrate PLAYER*’s superiority, advancing effective reasoning agents in complex environments.
-Limitations
-We identify some of the limitations below:
-Limited Annotation Scope
-Due to the cost of manual annotation and running experiments, we were unable to annotate all 100 games in the Conan dataset.
-However, we believe that the current scope of our study is sufficient to demonstrate the efficiency of our methodology.
-Future research can expand upon our work by annotating additional games and applying our approach to MMGs on a larger scale.
-Limited Evaluation of LLMs
-Due to the high cost associated with running experiments on LLMs, our experiments were primarily conducted using GPT-3.5.
-To further support our findings, we also performed smaller-scale experiments on LLaMa, which yielded results consistent with those obtained from GPT-3.5.
-However, we acknowledge that our inability to run more extensive evaluations across a wider range of popular LLMs, such as GPT-4 or lesser-known models, may have limited our ability to uncover additional insights.
-The cost of API usage for models like GPT-4 and the substantial running time required for models like LLaMa posed significant constraints on our experimental scope.
-These limitations highlight the need for future research to explore more cost-effective and efficient methods for evaluating and comparing the performance of various LLMs in the context of complex social deduction games.
+This study introduces Questum, a framework that tackles the challenges of LLM-driven agents in MMGs through sensor-based modeling, information gain-driven questioning, and suspect pruning.
+With the WellPlay dataset enabling robust evaluation, Questum  demonstrates superior reasoning, efficiency, and human-like interaction quality.
+These advancements pave the way for more effective and engaging AI systems in dynamic multi-agent scenarios.
+Limitation
+Annotation Scope
+Murder Mystery Games are built on intricately written human-authored stories, featuring dense plots, complex character relationships, and unexpected twists to enhance gameplay.
+These characteristics make annotation particularly challenging, explaining the scarcity of labeled datasets for MMGs.
+Existing resources include only 56 inferential questions across 4 selected MMGs.
+While this dataset is still insufficient for training due to the task’s inherent complexity—and the impracticality of creating a large dataset for such scenarios—we believe it adequately demonstrates the effectiveness of our methodology.
+Future research could explore synthetic data generation with enhanced quality control to further expand the dataset.
+Generalisation Across Domains
+While the proposed framework demonstrates strong performance for MMGs, its generalisability to other multi-agent or language-driven tasks has not been fully explored.
+Applying this approach to domains with different interaction dynamics or less structured narratives may require substantial modifications.
+Future research could extend the framework to other multi-agent scenarios, such as collaborative storytelling, negotiation games, or real-world applications like customer service.
+Domain adaptation techniques could help refine the framework for broader use cases.
+Resource Dependence
+The framework depends on powerful LLMs for sensor-based reasoning and question generation, which can be resource-intensive.
+Scaling this approach to larger or real-time systems may require additional computational and memory resources, limiting accessibility for smaller-scale applications.
+Efforts to optimise computational efficiency, such as caching mechanisms, or lightweight models for real-time applications, could make the framework more accessible and scalable.
 Ethics Statement
@@
 In the voting phase, each player needs to cast their vote for who they think is the murderer in each case.
-If the player with the most votes is the murderer, the civilian players win.
+If the player with over 50% of the votes is the murderer, the civilian players win.
 Otherwise, the murderer players win.
@@
 Stage 4: Voting
-In this stage, players vote anonymously to determine their suspicious regarding the identity of the murderer. Each player has one vote.
+In this stage, players vote anonymously to determine their suspicions regarding the identity of the murderer. Each player has one vote.
 Stage 5: Outcome Reveal
@@
 We model these characters through a set of agents, denoted as $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, where $N_{a}=6$ corresponds to the number of players.
-In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
+In parallel with the four victims, we define a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{v}=4$, representing the total number of victims in the scenario.
 Taking a closer look at “Tian Chou", depicted as the murderer player $a_{1}$, responsible for the demise of two among the four victims, her characterization unfolds as follows:
@@
 Your birth year was the twelfth of Guangxu’s reign during the Qing Dynasty (1886), marking the beginning of a life filled with extraordinary episodes…"
-  • Suspicion State ($s_{1}$): The script features four victims, each harboring suspicions towards the remaining five players.
-Consequently, $s_{1}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players.
+  • Suspicion State ($s_{i}jk$): The script features four victims, each harbouring suspicions towards the remaining five players.
+Consequently, $s_{i}$ is represented as a $4\times 5$ matrix, with each row corresponding to a victim and each column reflecting the suspicions they hold against other players.
 For example, a first row of $[1,0,0,1,0]$ signifies that victim “Cui Shouheng" suspects both “Zhou Lianyi" and “Yannan".
@@
     5. 5.
-During the game, the dialogue history is recorded in $D$, and as the game progresses, inquiries and responses are conducted in accordance with Algorithm 1, leading to updates in $s_{i}$.
 Appendix B The WellPlay Dataset
@@
     Background Script.
-For each character, there is corresponding background script includes all the information from their perspective.
+For each character, there is a corresponding background script includes all the information from their perspective.
 For example, for “Sylvia Costa” in the script “Death Wears White”, it is:
@@
 D. Father Tom is hate of Drake Li Morette
-Table A1: Examples of Each Type of Our Evaluation Questions
+Table 4: Examples of Each Type of Our Evaluation Questions
 Index Question
@@
 10 Do you have any personal opinions or theories about the case?
-Table A2: Predefined questions for Werewolf method.
+Table 5: Predefined questions for Werewolf method.
 B.2 Evaluation questions
 The evaluation questions are devised based on our annotations and aim to assess the understanding of the intricate relationships and narratives present within the game scripts.
-These questions are segmented into three categories, with each category of questions and examples shown in Table A1.
+These questions are segmented into three categories, with each category of questions and examples shown in Table 4.
   1. 1.
@@
 Specifically, Type A questions are valued at 10 points, Type B questions at 5 points, and Type C questions at 2 points.
-Detailed script statistics and evaluation question metrics are presented in Table1.
-B.3 Overall Performance Computing
+B.3 Annotation Details
+We recruited four human experts to read all the scripts and write the questions and answers.
+We conducted training sessions for them and organised game sessions to let them play and become familiar with the game flow.
+They were compensated at an hourly rate of \$31.92, with each narrative estimated to take about 5-20 hours to complete, depending on the complexity of the game.
+Due to the cost of human annotation on such a long and complicated corpus, we believe our current dataset is sufficient for evaluation purposes.
+Our dataset can serve as a starting point to evaluate various strategies and boost investment in larger-scale labelling and training for agents with enhanced inferential capabilities.
+B.4 Disclaimers of Risks
+Consider that the detective narrative dataset may contain descriptions of violent events, actions, or characters, we warn annotators of such content before the annotation process:
+Please be advised that the detective narrative dataset may contain descriptions of violent events, actions, or characters.
+This content is included solely for academic, research, and narrative analysis purposes.
+It is not meant to glorify or trivialise violence in any form.
+Annotators should be aware of potential exposure to violent content and engage with the dataset professionally and responsibly.
+B.5 Human Subject Evaluation Setup
+Since the same player would know the killer’s identity if they tried different models, we recruited four players to act as the murderer, each playing through three scripts (Sin, Deadly Fountain, Riverside Inn) across four methods, totalling 49 hours of gameplay.
+Each human player assumed the role of one character in the game, while the other characters were controlled by Questum -powered agents.
+Evaluation Werewolf O-CoT ThinkThrice Questum
+Story Advancement $1.50_{\pm 0.37}$ $2.08_{\pm 0.14}$ $2.67_{\pm 0.00}$ $3.50_{\pm 0.17}$
+Question Quality $1.67_{\pm 0.41}$ $1.58_{\pm 0.36}$ $2.08_{\pm 0.28}$ $3.25_{\pm 0.28}$
+Response Quality $2.83_{\pm 0.17}$ $2.92_{\pm 0.36}$ $3.00_{\pm 0.33}$ $3.50_{\pm 0.29}$
+Response Speed $1.17_{\pm 0.17}$ $1.33_{\pm 0.41}$ $1.17_{\pm 0.17}$ $2.50_{\pm 0.17}$
+Role Immersion $2.25_{\pm 0.36}$ $2.17_{\pm 0.5}$ $2.33_{\pm 0.24}$ $2.83_{\pm 0.29}$
+overall $1.88_{\pm 0.59}$ $2.01_{\pm 0.55}$ $2.25_{\pm 0.62}$ $3.11_{\pm 0.39}$
+Table 6: Results of Agent-vs-Human Evaluation using Human-Centric metrics.
+To ensure fairness, the human players were given the same information and rules as the agents, maintaining consistent conditions across both evaluation scenarios.
+Both the human players and the agents shared the same objectives: to achieve victory in the game by fulfilling their individual objectives.
+This setup allowed for a direct comparison between human-agent and agent-agent interactions, providing insights into Questum ’s ability to adapt and perform in mixed environments.
+Score Story Advancement Question Quality Response Quality Role Immersion Response Speed
+5 points Actively gathers information and conducts in-depth analyses.
+Reasoning is rigorous, offering key insights that significantly advance the narrative.
+Poses highly relevant, in-depth questions that directly propel the plot and uncover hidden clues.
+Provides accurate and comprehensive responses fully aligned with the character’s background and narrative needs.
+Offers clear, valuable information.
+Fully embodies the character.
+Actions, dialogue, and emotional expressions are authentically aligned with the character’s traits, leaving a strong impression.
+Responds promptly with no perceptible delay.
+The thought process is fluid and maintains a brisk, engaging pace.
+4 points Proactively seeks information and performs reasonable analysis.
+Reasoning is generally accurate and supports narrative progression.
+Most questions are relevant and yield useful information.
+Occasionally introduces innovative queries.
+Generally provides accurate and sufficiently complete answers.
+Minor omissions occur but do not hinder overall understanding.
+Frequently demonstrates character traits.
+Occasionally shows minor deviations, but emotional expression is largely appropriate.
+Responds relatively quickly with brief, occasional delays.
+The thought process is mostly smooth and well-paced.
+3 points Participates in information gathering but lacks analytical depth.
+Reasoning may be somewhat biased, offering moderate narrative progression.
+Some questions are relevant, but lack depth.
+Occasional repetition or irrelevant queries occur.
+Responses are basically accurate but sometimes incomplete or vague.
+May require further clarification by others.
+Occasionally exhibits character traits, but inconsistently.
+Emotional expression is average, and immersion is limited.
+Moderate response speed with some noticeable pauses.
+The reasoning process is acceptable but not seamless.
+2 points Shows limited initiative in gathering information, with superficial analysis.
+Reasoning is frequently flawed, hindering effective plot advancement.
+Most questions bear little relevance, often fail to obtain valuable clues, and frequently repeat similar inquiries.
+Responses are often inaccurate, incomplete, or misleading.
+Frequent clarification requests from others are necessary.
+Characterization is unclear; actions and dialogue often stray significantly from the intended character profile.
+Emotional expression is awkward or absent.
+Responses are slow with frequent delays.
+The reasoning process is disjointed, disturbing the overall gameplay flow.
+1 point Does not gather information or attempt analysis.
+Unable to perform meaningful reasoning, leaving the narrative stagnant.
+Poses irrelevant or no questions, providing no support for narrative progression.
+Responses are unrelated or outright refusals.
+Offers no assistance in understanding the narrative.
+Demonstrates no character immersion; actions and dialogue contradict character traits.
+No emotional expression is evident.
+Extremely slow responses or prolonged silence.
+The reasoning process halts, severely impeding the game’s progress.
+Table 7: Criteria for each evaluation dimension (Story Advancement, Question Quality, Response Quality, Role Immersion, and Response Speed).
+Appendix C Implementation
+C.1 Implementation Details
+For retrieval enhancement, we utilized the FAISS [^4] https://github.com/facebookresearch/faiss library to construct a vector database, creating FAISS indices using the L2 distance metric.
+Embeddings were generated using the Azure API’s text-embedding-ada-002 service.
+Scripts were divided into segments, with each segment having a maximum length of 50 tokens.
+For dialogue records, a question-and-answer pair was stored as a single segment.
+During retrieval, the maximum script and dialogue lengths included in the prompt were set to 4000 tokens.
+For evaluation, these maximum lengths were increased to 5000 tokens.
+Additionally, the hyperparameters were set with $\epsilon$ equal to 0.1 and $\beta$ equal to 0.2.
+Experiment
+Following the results of our ablation studies, the gameplay phase was structured to ask one question per round over three rounds.
+After the game concluded, the evaluation phase consisted of three separate evaluations, with the final results being the average of these evaluations.
+C.2 Overall Performance Computing
 In calculating the overall score for performance, we have employed both the weighted mean and the weighted standard deviation.
@@
 $s_{w}=\sqrt{s_{w}^{2}}$
-B.4 Annotation Details
-We recruited four human experts to read all the scripts and write the questions and answers.
-We conducted training sessions for them and organised game sessions to let them play and become familiar with the game flow.
-They were compensated at an hourly rate of \$31.92, with each narrative estimated to take about 5-20 hours to complete, depending on the complexity of the game.
-B.5 Disclaimers of Risks
-Consider that the detective narrative dataset may contain descriptions of violent events, actions, or characters, we warn annotators of such content before the annotation process:
-Please be advised that the detective narrative dataset may contain descriptions of violent events, actions, or characters.
-This content is included solely for academic, research, and narrative analysis purposes.
-It is not meant to glorify or trivialise violence in any form.
-Annotators should be aware of potential exposure to violent content and engage with the dataset professionally and responsibly.
-Appendix C Implementation
-C.1 Implementation Details
-RAG (Retrieval-Augmented Generation)
-For retrieval enhancement, we utilised the FAISS [^4] https://github.com/facebookresearch/faiss library to build a vector database, creating FAISS indices using the L2 distance.
-Embeddings were obtained via the Azure API’s text-embedding-ada-002 service.
-Scripts were stored in segments, with each segment having a maximum length of 50 tokens.
-For dialog records, a question-and-answer pair was stored as a single segment.
-During retrieval, the maximum script length and dialog length inserted into the prompt were both set to 4000 tokens.
-For evaluation, these maximum lengths were increased to 5000 tokens.
-Experiment
-Following the results of our ablation studies, the gameplay phase was structured to ask one question per round over three rounds.
-After the game concluded, the evaluation phase consisted of three separate evaluations, with the final results being the average of these evaluations.
-C.2 Comparison Models
+C.3 Comparison Models
 In this section, we provide a detailed overview of the methodologies compared in our study.
@@
   1. 1.
-    Werewolf(Xu et al., 2023)
+    Werewolf(Xu et al., 2023a)
     Question Generation:
     1. a.
-      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table A2.
+      Selecting questions from a predefined question written by the human specialist, based on the current dialogue and script. The list of expertly devised questions is presented in Table 5.
     2. b.
@@
     The OP evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
-C.3 Performance Evaluation Results
-Script Evaluation #QA PP OP Agent’s Response After Playing the Game
-Werewolf O-CoT ThinkThrice PLAYER*
-Death Wears White (9 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
-Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
-Reasoning $102$$454_{\pm.012}.$$503_{\pm.015}.$$350_{\pm.021}.$$395_{\pm.031}.$$392_{\pm.010}.$$408_{\pm.030}.$
-Relations $72$$431_{\pm.036}.$$425_{\pm.031}.$$454_{\pm.035}.$$472_{\pm.037}.$$398_{\pm.032}.$$500_{\pm.048}.$
-Overall $184$$420_{\pm.007}.$$480_{\pm.011}.$$367_{\pm.015}.$$393_{\pm.009}.$$377_{\pm.013}.$$407_{\pm.007}.$
-Ghost Revenge (7 players, 3 victims) Voting -- .$222_{\pm.416}.$$333_{\pm.471}.$$222_{\pm.416}.$$111_{\pm.314}.$$333_{\pm.471}.$$333_{\pm.471}.$
-Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
-Reasoning $152$$390_{\pm.020}.$$533_{\pm.030}.$$417_{\pm.023}.$$410_{\pm.019}.$$450_{\pm.056}.$$399_{\pm.040}.$
-Relations $69$$246_{\pm.014}.$$411_{\pm.103}.$$348_{\pm.014}.$$309_{\pm.047}.$$420_{\pm.038}.$$386_{\pm.030}.$
-Overall $240$$362_{\pm.016}.$$483_{\pm.020}.$$399_{\pm.009}.$$381_{\pm.049}.$$438_{\pm.030}.$$417_{\pm.032}.$
-Danshui Villa (7 players, 2 victims) Voting -- .$167_{\pm.373}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
-Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
-Reasoning $128$$315_{\pm.020}.$$422_{\pm.016}.$$307_{\pm.044}.$$352_{\pm.008}.$$347_{\pm.005}.$$367_{\pm.008}.$
-Relations $63$$323_{\pm.040}.$$476_{\pm.028}.$$344_{\pm.051}.$$381_{\pm.069}.$$333_{\pm.016}.$$370_{\pm.056}.$
-Overall $203$$326_{\pm.014}.$$403_{\pm.028}.$$293_{\pm.040}.$$349_{\pm.012}.$$339_{\pm.013}.$$367_{\pm.011}.$
-Unfinished Love (7 players, 2 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
-Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
-Reasoning $61$$552_{\pm.025}.$$563_{\pm.025}.$$563_{\pm.010}.$$486_{\pm.025}.$$552_{\pm.019}.$$601_{\pm.019}.$
-Relations $72$$500_{\pm.028}.$$546_{\pm.021}.$$523_{\pm.071}.$$583_{\pm.028}.$$556_{\pm.042}.$$616_{\pm.016}.$
-Overall $145$$457_{\pm.017}.$$475_{\pm.033}.$$469_{\pm.018}.$$467_{\pm.004}.$$460_{\pm.030}.$$525_{\pm.029}.$
-Cruise Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$333_{\pm.471}.$$333_{\pm.471}.$
-Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
-Reasoning $24$$653_{\pm.064}.$$472_{\pm.064}.$$542_{\pm.042}.$$528_{\pm.024}.$$528_{\pm.105}.$$514_{\pm.024}.$
-Relations $30$$267_{\pm.058}.$$411_{\pm.102}.$$456_{\pm.020}.$$456_{\pm.051}.$$422_{\pm.039}.$$434_{\pm.058}.$
-Overall $58$$459_{\pm.012}.$$415_{\pm.079}.$$510_{\pm.037}.$$503_{\pm.017}.$$509_{\pm.107}.$$520_{\pm.025}.$
-Sin (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$333_{\pm.471}.$
-Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
-Reasoning $20$$483_{\pm.058}.$$567_{\pm.029}.$$767_{\pm.058}.$$600_{\pm.100}.$$717_{\pm.058}.$$600_{\pm.000}.$
-Relations $21$$476_{\pm.082}.$$587_{\pm.099}.$$429_{\pm.000}.$$699_{\pm.055}.$$524_{\pm.048}.$$698_{\pm.027}.$
-Overall $44$$397_{\pm.029}.$$531_{\pm.053}.$$570_{\pm.000}.$$539_{\pm.029}.$$564_{\pm.012}.$$577_{\pm.053}.$
-Deadly Fountain (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
-Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
-Reasoning $21$$413_{\pm.153}.$$667_{\pm.126}.$$460_{\pm.027}.$$476_{\pm.082}.$$539_{\pm.055}.$$587_{\pm.028}.$
-Relations $12$$305_{\pm.048}.$$389_{\pm.048}.$$333_{\pm.084}.$$222_{\pm.096}.$$222_{\pm.048}.$$361_{\pm.127}.$
-Overall $36$$319_{\pm.097}.$$604_{\pm.104}.$$354_{\pm.022}.$$369_{\pm.053}.$$390_{\pm.043}.$$463_{\pm.041}.$
-Unbelievable Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
-Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
-Reasoning $24$$389_{\pm.105}.$$528_{\pm.064}.$$431_{\pm.064}.$$361_{\pm.064}.$$264_{\pm.127}.$$583_{\pm.042}.$
-Relations $15$$422_{\pm.102}.$$822_{\pm.102}.$$644_{\pm.077}.$$756_{\pm.102}.$$289_{\pm.154}.$$556_{\pm.077}.$
-Overall $43$$330_{\pm.081}.$$481_{\pm.020}.$$444_{\pm.037}.$$382_{\pm.035}.$$230_{\pm.034}.$$509_{\pm.038}.$
-Desperate Sunshine (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$667_{\pm.471}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
-Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
-Reasoning $18$$556_{\pm.056}.$$778_{\pm.096}.$$648_{\pm.032}.$$704_{\pm.032}.$$759_{\pm.064}.$$741_{\pm.085}.$
-Relations $36$$537_{\pm.064}.$$611_{\pm.048}.$$556_{\pm.028}.$$630_{\pm.043}.$$500_{\pm.056}.$$565_{\pm.070}.$
-Overall $57$$514_{\pm.013}.$$680_{\pm.047}.$$599_{\pm.023}.$$618_{\pm.020}.$$647_{\pm.075}.$$680_{\pm.047}.$
-Riverside Inn (4 players, 1 victim) Voting -- .$333_{\pm.471}.$ $1.000_{\pm.000}$$333_{\pm.471}.$$000_{\pm.000}.$ $1.000_{\pm.000}$$333_{\pm.471}.$
-Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
-Reasoning $18$$519_{\pm.032}.$$667_{\pm.056}.$$463_{\pm.064}.$$500_{\pm.056}.$$685_{\pm.032}.$$667_{\pm.056}.$
-Relations $18$$407_{\pm.085}.$$500_{\pm.096}.$$463_{\pm.085}.$$426_{\pm.064}.$$389_{\pm.000}.$$426_{\pm.032}.$
-Overall $39$$479_{\pm.032}.$$671_{\pm.036}.$$459_{\pm.073}.$$387_{\pm.020}.$$614_{\pm.018}.$$590_{\pm.023}.$
-Solitary Boat Firefly (6 players, 4 victims) Voting -- .$167_{\pm.373}.$$583_{\pm.493}.$$000_{\pm.000}.$$250_{\pm.433}.$$167_{\pm.373}.$$500_{\pm.500}.$
-Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
-Reasoning $109$$318_{\pm.014}.$$569_{\pm.018}.$$370_{\pm.014}.$$425_{\pm.046}.$$401_{\pm.014}.$$379_{\pm.005}.$
-Relations $69$$483_{\pm.017}.$$507_{\pm.014}.$$536_{\pm.014}.$$565_{\pm.038}.$$503_{\pm.044}.$$594_{\pm.087}.$
-Overall $198$$332_{\pm.015}.$$544_{\pm.024}.$$354_{\pm.023}.$$430_{\pm.016}.$$375_{\pm.015}.$$444_{\pm.030}.$
-Manna (6 players, 3 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$222_{\pm.416}.$$333_{\pm.471}.$$444_{\pm.497}.$$556_{\pm.497}.$
-Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
-Reasoning $123$$458_{\pm.005}.$$393_{\pm.040}.$$569_{\pm.028}.$$499_{\pm.033}.$$526_{\pm.017}.$$539_{\pm.021}.$
-Relations $88$$447_{\pm.058}.$$640_{\pm.036}.$$538_{\pm.043}.$$614_{\pm.030}.$$511_{\pm.034}.$$580_{\pm.046}.$
-Overall $235$$408_{\pm.017}.$$434_{\pm.002}.$$525_{\pm.017}.$$505_{\pm.004}.$$501_{\pm.025}.$$553_{\pm.045}.$
-Overall Voting -- .$074_{\pm.117}.$$326_{\pm.342}.$$176_{\pm.257}.$$058_{\pm.115}.$$245_{\pm.326}.$$282_{\pm.308}.$
-Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
-Reasoning $800$$417_{\pm.095}.$$508_{\pm.094}.$$437_{\pm.104}.$$430_{\pm.077}.$$449_{\pm.113}.$$465_{\pm.106}.$
-Relations $565$$411_{\pm.102}.$$513_{\pm.110}.$$466_{\pm.090}.$$495_{\pm.133}.$$444_{\pm.089}.$$500_{\pm.109}.$
-Overall $1482$$386_{\pm.057}.$$484_{\pm.072}.$$415_{\pm.074}.$$424_{\pm.067}.$$426_{\pm.087}.$$461_{\pm.086}.$
-Table A3: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
-PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
-For English Dataset
-Due to budgetary constraints, we only evaluated four scripts from the English dataset, with the performance results reported in Table A4.
-For the same scripts, we found that the experiments conducted on the English corpus corroborate the results obtained from the Chinese corpus.
-Additionally, we observed that the performance of agents based on the English corpus surpassed those based on the Chinese corpus, showing the differential inferencing abilities of LLMs across languages.
-This discrepancy could be attributed to language biases inherent in the training data utilised for these models.
-Script Evaluation #QA PP OP Agent’s Response After Playing the Game
-Werewolf O-CoT ThinkThrice PLAYER*
-Death Wears White (9 players, 1 victim) Objective $10$$200_{\pm.173}.$$900_{\pm.100}.$$300_{\pm.100}.$$267_{\pm.058}.$$300_{\pm.000}.$$267_{\pm.058}.$
-Reasoning $102$$350_{\pm.044}.$$520_{\pm.026}.$$356_{\pm.037}.$$363_{\pm.049}.$$399_{\pm.006}.$$441_{\pm.030}.$
-Relations $72$$547_{\pm.008}.$$445_{\pm.026}.$$495_{\pm.016}.$$495_{\pm.057}.$$398_{\pm.032}.$$491_{\pm.021}.$
-Overall $184$$367_{\pm.040}.$$549_{\pm.012}.$$375_{\pm.033}.$$375_{\pm.026}.$$385_{\pm.010}.$$427_{\pm.029}.$
-Ghost Revenge (7 players, 3 victims) Objective $19$$403_{\pm.061}.$$561_{\pm.132}.$$439_{\pm.109}.$$211_{\pm.106}.$$211_{\pm.106}.$$526_{\pm.106}.$
-Reasoning $152$$456_{\pm.004}.$$507_{\pm.007}.$$439_{\pm.031}.$$539_{\pm.023}.$$441_{\pm.017}.$$423_{\pm.031}.$
-Relations $69$$309_{\pm.009}.$$261_{\pm.025}.$$328_{\pm.059}.$$304_{\pm.015}.$$280_{\pm.017}.$$353_{\pm.059}.$
-Overall $240$$428_{\pm.010}.$$485_{\pm.024}.$$425_{\pm.002}.$$452_{\pm.023}.$$380_{\pm.010}.$$432_{\pm.013}.$
-Danshui Villa (7 players, 2 victims) Objective $12$$167_{\pm.083}.$$278_{\pm.048}.$$472_{\pm.048}.$$333_{\pm.000}.$$305_{\pm.048}.$$389_{\pm.096}.$
-Reasoning $128$$325_{\pm.031}.$$427_{\pm.009}.$$401_{\pm.020}.$$373_{\pm.020}.$$344_{\pm.014}.$$357_{\pm.033}.$
-Relations $63$$328_{\pm.056}.$$460_{\pm.028}.$$391_{\pm.066}.$$312_{\pm.106}.$$376_{\pm.046}.$$407_{\pm.075}.$
-Overall $203$$292_{\pm.012}.$$412_{\pm.017}.$$409_{\pm.009}.$$359_{\pm.010}.$$343_{\pm.010}.$$369_{\pm.033}.$
-Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$195_{\pm.048}.$$389_{\pm.048}.$$417_{\pm.000}.$$361_{\pm.048}.$$528_{\pm.127}.$
-Reasoning $61$$536_{\pm.018}.$$557_{\pm.028}.$$650_{\pm.038}.$$612_{\pm.009}.$$634_{\pm.038}.$$656_{\pm.000}.$
-Relations $72$$491_{\pm.016}.$$551_{\pm.016}.$$509_{\pm.032}.$$481_{\pm.021}.$$592_{\pm.008}.$$500_{\pm.037}.$
-Overall $145$$446_{\pm.006}.$$479_{\pm.029}.$$560_{\pm.011}.$$538_{\pm.010}.$$566_{\pm.009}.$$589_{\pm.018}.$
-Table A4: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
-PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
-Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved.
-This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
-Script #Tokens Stage PP OP Agent’s Response After Playing the Game
-Werewolf O-CoT ThinkThrice PLAYER*
-Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $3.363$
-Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
-Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $5.947$
-Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
-Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $4.603$
-Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
-Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $3.505$
-Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
-Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.846$
-Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
-Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
-Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
-Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.596$
-Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
-Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $1.150$
-Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
-Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.701$
-Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
-Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.561$
-Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
-Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $6.800$
-Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
-Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $5.028$
-Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
-Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
-Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
-Table A5: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
-#$\#Tokens$ represent the average length of each character’s script.
-Costs are reported for one complete gameplay and one evaluation process for each script.
-C.4 Experiments with Open-Source LLMs
-Models Llama2 70b Llama2 13b Llama2 7b gemma 7b
-Overall Score 0.312 0.281 0.267 0.273
-Table A6: Compare the performance of PLAYER* method with different open-source LLMS.
-In addition to experimenting with GPT-3.5-turbo-16k 0613, we also explored Llama2 (70b, 13b, 7b) and Gemma 7b.
-These models were tested using the scenario “Solitary Boat Firefly", and the overall results are reported in Table A6.
-The findings indicate that all four models scored significantly lower than GPT-3.5, presumably due to the limitations imposed by a 4k context window.
-This limitation likely hindered the models’ ability to encapsulate sufficient relevant information within such a constrained window.
-After thorough testing, we observed that despite being evaluated on a Chinese dataset, Llama2 70b primarily resorted to English conversations due to its limited proficiency in Chinese.
-It responded in English even to Chinese prompts.
-Other models struggled even more with executing the prompts as required.
-This limitation greatly hindered their performance in complex gaming situations, such as MMGs.
-Therefore, we decided to exclusively use GPT-3.5 for future experiments, given its ability to navigate these intricate scenarios.
-C.5 Sensors
-This section provides a detailed explanation of the sensors employed in the 2.3, which are essential for both the Search by Questioning and Action Space Refinement components.
+C.4 Sensors
+This section provides a detailed explanation of the sensors, which are essential for both the Search by Questioning and Action Space Refinement components.
   • Emotion Sensor: Assesses emotional inclination towards a character.
@@
 "choices": ["High", "Medium", "Low"],
-C.6 Prompt
-C.6.1 System Prompt
+C.5 Prompt
+C.5.1 System Prompt
 System Prompt
@@
 Please hide the fact that you committed the murder by fabricating lies and other information, and accomplish your personal objective!
-C.6.2 Gameplay Prompt
+C.5.2 Gameplay Prompt
 Self-Introduction Prompt
@@
 Sensor Prompt
-As outlined in sections 2.3 and C.5, the Sensor Prompt is crafted to collect a wide array of crucial information.
+As outlined in sections 2.2 and C.4, the Sensor Prompt is crafted to collect a wide array of crucial information.
 Within this framework, the placeholders {victim} and {character} are employed to reference the victim’s name and the agent being questioned, respectively.
 Meanwhile, {current_script} and {dialog_history} retrieve the script and dialog logs relevant to both the deceased and the agent under scrutiny, utilising the RAG technique.
-The constructs {sensor_Prompt} and {choices} have been formulated in C.5 within the Sensor Prompt segment.
+The constructs {sensor_Prompt} and {choices} have been formulated in C.4 within the Sensor Prompt segment.
 For Civilian Players:
@@
 Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data.
-The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 4.3 chapter.
+The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the D.3 chapter.
 For Civilian Players:
@@
 Please answer the question from a first-person perspective, rather than saying what someone else said.
-C.6.3 Evaluation Prompt
+C.5.3 Evaluation Prompt
 Single-Choice Question Template for the Evaluation Stage Prompt
@@
 "answer": "a,b"
+Appendix D Result
+D.1 Performance Evaluation Results
+Script Evaluation #QA PP OP Agent’s Response After Playing the Game
+Werewolf O-CoT ThinkThrice Questum
+Death Wears White (9 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective 10 .$067_{\pm.047}.$$467_{\pm.094}.$$033_{\pm.047}.$$067_{\pm.047}.$$067_{\pm.047}.$$033_{\pm.047}.$
+Reasoning 102 .$310_{\pm.026}.$$503_{\pm.012}.$$258_{\pm.024}.$$324_{\pm.014}.$$327_{\pm.012}.$$356_{\pm.030}.$
+Relations 72 .$356_{\pm.040}.$$425_{\pm.025}.$$421_{\pm.024}.$$458_{\pm.041}.$$384_{\pm.026}.$$435_{\pm.017}.$
+Overall 184 .$287_{\pm.017}.$$480_{\pm.009}.$$260_{\pm.007}.$$315_{\pm.006}.$$303_{\pm.014}.$$328_{\pm.026}.$
+Ghost Revenge (7 players, 3 victims) Win Rate - .$000_{\pm.000}.$$333_{\pm.000}.$$222_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$333_{\pm.000}.$
+Objective 19 .$158_{\pm.043}.$$333_{\pm.108}.$$193_{\pm.066}.$$158_{\pm.043}.$$193_{\pm.025}.$$333_{\pm.066}.$
+Reasoning 152 .$300_{\pm.011}.$$533_{\pm.025}.$$307_{\pm.008}.$$322_{\pm.019}.$$377_{\pm.022}.$$353_{\pm.016}.$
+Relations 69 .$222_{\pm.025}.$$411_{\pm.084}.$$314_{\pm.018}.$$295_{\pm.018}.$$353_{\pm.025}.$$353_{\pm.018}.$
+Overall 240 .$266_{\pm.016}.$$483_{\pm.016}.$$288_{\pm.009}.$$290_{\pm.016}.$$342_{\pm.015}.$$350_{\pm.002}.$
+Danshui Villa (7 players, 2 victims) Win Rate - .$000_{\pm.000}.$$167_{\pm.236}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective 12 .$056_{\pm.039}.$$222_{\pm.104}.$$083_{\pm.000}.$$111_{\pm.039}.$$194_{\pm.039}.$$111_{\pm.079}.$
+Reasoning 128 .$247_{\pm.004}.$$422_{\pm.013}.$$286_{\pm.027}.$$286_{\pm.019}.$$310_{\pm.004}.$$286_{\pm.004}.$
+Relations 63 .$254_{\pm.047}.$$476_{\pm.022}.$$312_{\pm.027}.$$365_{\pm.045}.$$259_{\pm.015}.$$296_{\pm.075}.$
+Overall 203 .$222_{\pm.014}.$$403_{\pm.023}.$$263_{\pm.020}.$$274_{\pm.015}.$$287_{\pm.008}.$$264_{\pm.024}.$
+Unfinished Love (7 players, 2 victims) Win Rate - .$000_{\pm.000}.$$167_{\pm.236}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$500_{\pm.000}.$
+Objective 12 .$139_{\pm.039}.$$167_{\pm.068}.$$083_{\pm.118}.$$028_{\pm.039}.$$000_{\pm.000}.$$333_{\pm.068}.$
+Reasoning 61 .$475_{\pm.013}.$$563_{\pm.020}.$$443_{\pm.023}.$$426_{\pm.035}.$$481_{\pm.039}.$$536_{\pm.028}.$
+Relations 72 .$421_{\pm.026}.$$546_{\pm.017}.$$472_{\pm.034}.$$519_{\pm.036}.$$514_{\pm.020}.$$560_{\pm.007}.$
+Overall 145 .$391_{\pm.006}.$$475_{\pm.027}.$$374_{\pm.019}.$$366_{\pm.009}.$$388_{\pm.026}.$$499_{\pm.019}.$
+Cruise Incident (5 players, 1 victim) Win Rate - .$000_{\pm.000}.$$333_{\pm.471}.$$667_{\pm.471}.$ $1.000_{\pm.000}$$667_{\pm.471}.$ $1.000_{\pm.000}$
+Objective 4 .$000_{\pm.000}.$$250_{\pm.204}.$$417_{\pm.118}.$$500_{\pm.000}.$$583_{\pm.312}.$$667_{\pm.236}.$
+Reasoning 24 .$417_{\pm.059}.$$472_{\pm.052}.$$458_{\pm.068}.$$444_{\pm.086}.$$458_{\pm.059}.$$528_{\pm.052}.$
+Relations 30 .$211_{\pm.042}.$$411_{\pm.083}.$$367_{\pm.082}.$$422_{\pm.042}.$$411_{\pm.016}.$$422_{\pm.016}.$
+Overall 58 .$285_{\pm.022}.$$415_{\pm.065}.$$426_{\pm.031}.$$448_{\pm.050}.$$468_{\pm.068}.$$524_{\pm.015}.$
+Sin (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$333_{\pm.471}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$
+Objective 3 .$000_{\pm.000}.$$333_{\pm.272}.$$333_{\pm.272}.$$000_{\pm.000}.$$000_{\pm.000}.$$444_{\pm.314}.$
+Reasoning 20 .$433_{\pm.062}.$$567_{\pm.024}.$$650_{\pm.108}.$$467_{\pm.024}.$$533_{\pm.047}.$$550_{\pm.041}.$
+Relations 21 .$349_{\pm.059}.$$587_{\pm.081}.$$333_{\pm.067}.$$571_{\pm.067}.$$413_{\pm.090}.$$492_{\pm.022}.$
+Overall 44 .$337_{\pm.037}.$$531_{\pm.043}.$$517_{\pm.087}.$$411_{\pm.015}.$$411_{\pm.043}.$$517_{\pm.064}.$
+Deadly Fountain (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective 3 .$000_{\pm.000}.$$556_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Reasoning 21 .$317_{\pm.192}.$$667_{\pm.103}.$$381_{\pm.000}.$$444_{\pm.098}.$$508_{\pm.022}.$$587_{\pm.022}.$
+Relations 12 .$250_{\pm.068}.$$389_{\pm.039}.$$250_{\pm.068}.$$194_{\pm.039}.$$222_{\pm.039}.$$333_{\pm.068}.$
+Overall 36 .$247_{\pm.126}.$$604_{\pm.085}.$$289_{\pm.010}.$$323_{\pm.060}.$$369_{\pm.013}.$$438_{\pm.008}.$
+Unbelievable Incident (5 players, 1 victim) Win Rate - .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective 4 .$083_{\pm.118}.$$083_{\pm.118}.$$083_{\pm.118}.$$000_{\pm.000}.$$000_{\pm.000}.$$083_{\pm.118}.$
+Reasoning 24 .$375_{\pm.068}.$$528_{\pm.052}.$$431_{\pm.052}.$$278_{\pm.071}.$$194_{\pm.071}.$$472_{\pm.020}.$
+Relations 15 .$400_{\pm.054}.$$822_{\pm.083}.$$622_{\pm.083}.$$733_{\pm.109}.$$289_{\pm.126}.$$533_{\pm.054}.$
+Overall 43 .$318_{\pm.057}.$$481_{\pm.016}.$$388_{\pm.032}.$$291_{\pm.061}.$$168_{\pm.034}.$$400_{\pm.019}.$
+Desperate Sunshine (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
+Objective 3 .$333_{\pm.000}.$$556_{\pm.157}.$$333_{\pm.000}.$$111_{\pm.157}.$$556_{\pm.157}.$$778_{\pm.157}.$
+Reasoning 18 .$519_{\pm.052}.$$778_{\pm.079}.$$537_{\pm.026}.$$630_{\pm.094}.$$759_{\pm.052}.$$741_{\pm.069}.$
+Relations 36 .$500_{\pm.023}.$$611_{\pm.039}.$$491_{\pm.094}.$$574_{\pm.035}.$$491_{\pm.047}.$$556_{\pm.068}.$
+Overall 57 .$483_{\pm.018}.$$681_{\pm.038}.$$488_{\pm.034}.$$528_{\pm.058}.$$627_{\pm.032}.$$677_{\pm.041}.$
+Riverside Inn (4 players, 1 victim) Win Rate - .$000_{\pm.000}.$$667_{\pm.118}.$$083_{\pm.118}.$$000_{\pm.000}.$$083_{\pm.118}.$$250_{\pm.204}.$
+Objective 3 .$067_{\pm.024}.$$500_{\pm.041}.$$117_{\pm.024}.$$117_{\pm.062}.$$117_{\pm.094}.$$383_{\pm.062}.$
+Reasoning 18 .$248_{\pm.020}.$$569_{\pm.015}.$$248_{\pm.013}.$$312_{\pm.022}.$$281_{\pm.011}.$$373_{\pm.017}.$
+Relations 18 .$454_{\pm.018}.$$507_{\pm.012}.$$531_{\pm.014}.$$541_{\pm.038}.$$483_{\pm.030}.$$589_{\pm.071}.$
+Overall 39 .$239_{\pm.018}.$$544_{\pm.019}.$$262_{\pm.008}.$$304_{\pm.025}.$$276_{\pm.024}.$$409_{\pm.009}.$
+Solitary Boat Firefly (6 players, 4 victims) Win Rate - .$222_{\pm.157}.$$222_{\pm.157}.$$222_{\pm.157}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
+Objective 20 .$167_{\pm.000}.$$389_{\pm.071}.$$250_{\pm.090}.$$181_{\pm.039}.$$181_{\pm.039}.$$250_{\pm.034}.$
+Reasoning 109 .$415_{\pm.011}.$$393_{\pm.033}.$$409_{\pm.004}.$$369_{\pm.044}.$$453_{\pm.008}.$$539_{\pm.027}.$
+Relations 69 .$424_{\pm.042}.$$640_{\pm.030}.$$473_{\pm.028}.$$568_{\pm.048}.$$439_{\pm.033}.$$542_{\pm.027}.$
+Overall 198 .$359_{\pm.013}.$$434_{\pm.002}.$$383_{\pm.024}.$$359_{\pm.027}.$$387_{\pm.010}.$$472_{\pm.028}.$
+Manna (6 players, 3 victims) Win Rate - .$333_{\pm.471}.$ $1.000_{\pm.000}$$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$333_{\pm.471}.$
+Objective 24 .$444_{\pm.157}.$$889_{\pm.157}.$$111_{\pm.157}.$$000_{\pm.000}.$$556_{\pm.157}.$$444_{\pm.157}.$
+Reasoning 123 .$519_{\pm.026}.$$667_{\pm.045}.$$463_{\pm.052}.$$426_{\pm.026}.$$593_{\pm.026}.$$648_{\pm.069}.$
+Relations 88 .$370_{\pm.094}.$$500_{\pm.079}.$$444_{\pm.045}.$$407_{\pm.069}.$$333_{\pm.045}.$$444_{\pm.045}.$
+Overall 235 .$470_{\pm.030}.$$671_{\pm.029}.$$391_{\pm.058}.$$340_{\pm.005}.$$526_{\pm.055}.$$562_{\pm.076}.$
+Overall Win Rate - .$016_{\pm.022}.$$300_{\pm.041}.$$127_{\pm.059}.$$063_{\pm.022}.$$111_{\pm.045}.$$222_{\pm.081}.$
+Objective 117 .$120_{\pm.007}.$$370_{\pm.029}.$$160_{\pm.020}.$$123_{\pm.008}.$$162_{\pm.024}.$$288_{\pm.021}.$
+Reasoning 800 .$336_{\pm.007}.$$508_{\pm.009}.$$343_{\pm.006}.$$349_{\pm.012}.$$384_{\pm.002}.$$423_{\pm.009}.$
+Relations 565 .$360_{\pm.005}.$$513_{\pm.007}.$$425_{\pm.008}.$$473_{\pm.010}.$$405_{\pm.006}.$$471_{\pm.009}.$
+Overall 1482 .$300_{\pm.006}.$$484_{\pm.012}.$$324_{\pm.004}.$$329_{\pm.008}.$$347_{\pm.006}.$$407_{\pm.006}.$
+Table 8: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
+PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
+For English Dataset
+Due to budgetary constraints, we only evaluated four scripts from the English dataset, with the performance results reported in Table 9.
+For the same scripts, we found that the experiments conducted on the English corpus corroborate the results obtained from the Chinese corpus.
+Additionally, we observed that the performance of agents based on the English corpus surpassed those based on the Chinese corpus, showing the differential inferencing abilities of LLMs across languages.
+This discrepancy could be attributed to language biases inherent in the training data utilised for these models.
+Script Evaluation #QA PP OP Agent’s Response After Playing the Game
+Werewolf O-CoT ThinkThrice Questum
+Death Wears White (9 players, 1 victim) Objective $10$$200_{\pm.173}.$$900_{\pm.100}.$$300_{\pm.100}.$$267_{\pm.058}.$$300_{\pm.000}.$$267_{\pm.058}.$
+Reasoning $102$$350_{\pm.044}.$$520_{\pm.026}.$$356_{\pm.037}.$$363_{\pm.049}.$$399_{\pm.006}.$$441_{\pm.030}.$
+Relations $72$$547_{\pm.008}.$$445_{\pm.026}.$$495_{\pm.016}.$$495_{\pm.057}.$$398_{\pm.032}.$$491_{\pm.021}.$
+Overall $184$$367_{\pm.040}.$$549_{\pm.012}.$$375_{\pm.033}.$$375_{\pm.026}.$$385_{\pm.010}.$$427_{\pm.029}.$
+Ghost Revenge (7 players, 3 victims) Objective $19$$403_{\pm.061}.$$561_{\pm.132}.$$439_{\pm.109}.$$211_{\pm.106}.$$211_{\pm.106}.$$526_{\pm.106}.$
+Reasoning $152$$456_{\pm.004}.$$507_{\pm.007}.$$439_{\pm.031}.$$539_{\pm.023}.$$441_{\pm.017}.$$423_{\pm.031}.$
+Relations $69$$309_{\pm.009}.$$261_{\pm.025}.$$328_{\pm.059}.$$304_{\pm.015}.$$280_{\pm.017}.$$353_{\pm.059}.$
+Overall $240$$428_{\pm.010}.$$485_{\pm.024}.$$425_{\pm.002}.$$452_{\pm.023}.$$380_{\pm.010}.$$432_{\pm.013}.$
+Danshui Villa (7 players, 2 victims) Objective $12$$167_{\pm.083}.$$278_{\pm.048}.$$472_{\pm.048}.$$333_{\pm.000}.$$305_{\pm.048}.$$389_{\pm.096}.$
+Reasoning $128$$325_{\pm.031}.$$427_{\pm.009}.$$401_{\pm.020}.$$373_{\pm.020}.$$344_{\pm.014}.$$357_{\pm.033}.$
+Relations $63$$328_{\pm.056}.$$460_{\pm.028}.$$391_{\pm.066}.$$312_{\pm.106}.$$376_{\pm.046}.$$407_{\pm.075}.$
+Overall $203$$292_{\pm.012}.$$412_{\pm.017}.$$409_{\pm.009}.$$359_{\pm.010}.$$343_{\pm.010}.$$369_{\pm.033}.$
+Unfinished Love (7 players, 2 victims) Objective $12$$167_{\pm.000}.$$195_{\pm.048}.$$389_{\pm.048}.$$417_{\pm.000}.$$361_{\pm.048}.$$528_{\pm.127}.$
+Reasoning $61$$536_{\pm.018}.$$557_{\pm.028}.$$650_{\pm.038}.$$612_{\pm.009}.$$634_{\pm.038}.$$656_{\pm.000}.$
+Relations $72$$491_{\pm.016}.$$551_{\pm.016}.$$509_{\pm.032}.$$481_{\pm.021}.$$592_{\pm.008}.$$500_{\pm.037}.$
+Overall $145$$446_{\pm.006}.$$479_{\pm.029}.$$560_{\pm.011}.$$538_{\pm.010}.$$566_{\pm.009}.$$589_{\pm.018}.$
+Table 9: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games.
+PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
+D.2 Experiments with Open-Source LLMs
+Models Llama2 70b Llama2 13b Llama2 7b gemma 7b
+Overall Score 0.312 0.281 0.267 0.273
+Table 10: Compare the performance of Questum  method with different open-source LLMS.
+Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved.
+This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
+Script #Tokens Stage PP OP Agent’s Response After Playing the Game
+Werewolf O-CoT ThinkThrice Questum
+Death Wears White (9 players, 1 victim) 3,190 Gameplay -- -- $3.797$ $4.643$ $3.992$ $1.782$
+Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
+Ghost Revenge (7 players, 3 victims) 5,487 Gameplay -- -- $6.702$ $8.231$ $7.056$ $3.152$
+Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
+Danshui Villa (7 players, 2 victims) 5,111 Gameplay -- -- $5.215$ $6.262$ $5.449$ $2.440$
+Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
+Unfinished Love (7 players, 2 victims) 2,501 Gameplay -- -- $3.945$ $4.754$ $4.237$ $1.858$
+Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
+Cruise Incident (5 players, 1 victim) 1,262 Gameplay -- -- $0.975$ $1.146$ $1.021$ $0.448$
+Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
+Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.320$
+Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
+Deadly Fountain (4 players, 1 victim) 1,852 Gameplay -- -- $0.671$ $0.810$ $0.724$ $0.316$
+Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
+Unbelievable Incident (5 players, 1 victim) 3,182 Gameplay -- -- $1.304$ $1.567$ $1.367$ $0.610$
+Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
+Desperate Sunshine (4 players, 1 victim) 3,370 Gameplay -- -- $0.803$ $0.972$ $0.847$ $0.372$
+Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
+Riverside Inn (4 players, 1 victim) 1,909 Gameplay -- -- $0.633$ $0.762$ $0.682$ $0.297$
+Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
+Solitary Boat Firefly (6 players, 4 victims) 8,893 Gameplay -- -- $7.799$ $9.257$ $8.252$ $3.604$
+Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
+Manna (6 players, 3 victims) 9,028 Gameplay -- -- $5.805$ $6.925$ $6.108$ $2.665$
+Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
+Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $\bm{17.862}$
+Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
+Table 11: Compare the costs in US dollars(\$) of calling OpenAI API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage.
+#$\#Tokens$ represent the average length of each character’s script.
+Costs are reported for one complete gameplay and one evaluation process for each script.
+In addition to experimenting with GPT-3.5-turbo-16k 0613, we also explored Llama2 (70b, 13b, 7b) and Gemma 7b.
+These models were tested using the scenario “Solitary Boat Firefly", and the overall results are reported in Table 10.
+The findings indicate that all four models scored significantly lower than GPT-3.5, presumably due to the limitations imposed by a 4k context window.
+This limitation likely hindered the models’ ability to encapsulate sufficient relevant information within such a constrained window.
+After thorough testing, we observed that despite being evaluated on a Chinese dataset, Llama2 70b primarily resorted to English conversations due to its limited proficiency in Chinese.
+It responded in English even to Chinese prompts.
+Other models struggled even more with executing the prompts as required.
+This limitation greatly hindered their performance in complex gaming situations, such as MMGs.
+Therefore, we decided to exclusively use GPT-3.5 for future experiments, given its ability to navigate these intricate scenarios.
+D.3 Ablation Studies
+As shown in Table 12, we can observe that the two main modules “Search with Sensors" and “Approximate with Pruner" in our algorithm both reduced the cost and improved the performance as expected.
+Questum  equipped with both modules gives the best performance.
+Score Gameplay Costs (\$) Gameplay Times (h)
+Questum 0.407 17.861 5.740
+w/o Pruner 0.396 28.731 9.932
+w/o Sensor 0.379 15.057 4.457
+w/o Sensor & Pruner 0.357 19.144 6.577
+Table 12: Ablation study of removing the Pruner and Sensor on performance, gameplay costs, and times.
+Sensor Selection
+Figure 6: Compare the number of sensors, score, and cost.
+The chart displays the range of scores for different sensor combinations, with the I-shaped markers representing the score range and the · indicating the average score.
+In practice, sensors may exhibit non-linear relationships, which necessitates careful evaluation to avoid redundancy.
+So we assess the unique contributions of each new sensor, ensuring that every selected sensor introduces meaningful and distinct information.
+For MMGs, we initialise 5 sensors with domain knowledge: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible.
+Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator.
+Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime.
+Evidence: This sensor evaluates whether there is evidence that directly links the mentioned character to the crime scene.
+Background: This sensor evaluates whether there is any history of conflict, rivalry, or enmity between the mentioned character and the rests.
+As shown in Figure 6, we evaluate the performance gain provided by each sensor.
+The benefits diminish as additional sensors contribute less new information, while the negative impact from longer input sequences grows.
+Moreover, adding more sensors increases computational costs.
+To balance sensor effectiveness with task complexity, we selected three sensors.
+The final combination—Emotion, Motivation, and Opportunity Assessment—offered the best performance in our experiments.
+The experiment was run with different combinations of sensors.
+For example, we ran the experiment $\binom{5}{2}$ times to cover all possible combinations of sensors when the number of sensors is $2$.
+Optimal Number of Rounds and Questum  plannings
+We compared agents’ behavior across different numbers of rounds and with varying numbers of questions agents can ask per round, as shown in Figure 7.
+Performance peaks around round 3, after which it shows variability, with some rounds experiencing slight declines or plateauing in scores despite more rounds or questions.
+This indicates that after a rapid initial learning or adaptation phase, where agents effectively use additional questions to enhance their understanding and strategies, the value of IGed from conversations tends to converge.
+These results also provide empirical support for the assumption made in our methodology that the more inquiries we pose to an agent, the expected reward associated with questioning the same agent decreases.
+For the main results we reported, we use the original setting for number of rounds in MMG, which is 3, and based on the outcomes of the ablation studies, we chose the most effective number of questions to ask each round, which is $1$ question.
+Figure 7: Comparison of agents’ behaviour across different numbers of rounds, where each agent can ask a specific number of questions (denoted as $m$).
+Appendix E Agent Interaction Examples
+To illustrate the effectiveness of our proposed agent, Questum, we present examples of dialogue records showcasing interactions between human players and agent players.
+Below are four versions in which the agent-character Inspector Howard interrogates Farmer Jack about eyewitness details concerning the murder of Merchant Green.
+These versions correspond to different models: Werewolf, ThinkThrice, O-CoT, and Questum.
+E.1 Dialogue Versions
+E.1.1 Werewolf
+Inspector Howard $\rightarrow$ Farmer Jack:
+“Did you notice any unusual interactions between Merchant Green and other villagers?”
+Farmer Jack replies:
+“No, nothing in particular.”
+E.1.2 O-CoT
+Inspector Howard $\rightarrow$ Farmer Jack:
+“Jack, that night when you went to Merchant Green’s house, did you hear any voices or special sounds?”
+Farmer Jack replies:
+“I didn’t hear any voices, but I did hear the sound of a door being kicked, which was very loud and abrupt.”
+E.1.3 ThinkThrice
+Inspector Howard $\rightarrow$ Farmer Jack:
+“That night, did you notice any unusual situations around Merchant Green’s house, like arguments or someone leaving?”
+Farmer Jack replies:
+“I heard some sounds, but couldn’t see who it was.”
+E.1.4 Questum
+Inspector Howard $\rightarrow$ Farmer Jack:
+“Jack, that night you saw someone leaving Merchant Green’s house, can you recall what that person was wearing, their actions, or their general direction?”
+Farmer Jack replies:
+“Inspector, that night it was very dark, I could only faintly see that the person was wearing dark clothes, moving very quickly towards the back mountain path.
+I didn’t hear any sounds from him, but his footsteps were very hurried.”
+E.2 Analysis of Models
+Story Progression: Questum significantly enhances story progression by providing key clues such as the perpetrator’s dark clothing, rapid movement, and the direction towards the back mountain path.
+This enriches the narrative by introducing specific elements that can be further investigated.
+In contrast, Werewolf offers minimal advancement with a vague response denying any unusual interactions.
+ThinkThrice slightly improves by mentioning heard sounds but lacks actionable details, while O-CoT introduces the sound of a door being kicked but does not provide additional directions for the investigation.
+Questioning Ability: Questum demonstrates superior questioning ability by targeting specific details—clothing, actions, and direction—thereby guiding Farmer Jack to provide more focused and relevant information.
+Werewolf poses an overly broad question that fails to elicit detailed responses.
+ThinkThrice narrows the focus slightly but remains too general, and O-CoT asks a more specific question but does not explore as many facets as Questum.
+Answering Ability: Responses in Questum are notably detailed, offering observations about the perpetrator’s appearance and behavior, which are crucial for the investigation.
+This contrasts with Werewolf, where the answer is too brief and lacks specific observations.
+ThinkThrice provides a partial detail about hearing sounds but fails to elaborate further, while O-CoT mentions the sound of a door being kicked but does not add more context or clues.
+Role Immersion: Questum excels in role immersion by portraying Farmer Jack as a tense and observant witness, recalling specific details under pressure.
+This adds depth to his character and makes the interaction more engaging.
+Werewolf’s response lacks emotional depth, presenting a mechanical denial.
+ThinkThrice, while slightly more descriptive, still does not convey significant emotional or psychological engagement.
+O-CoT includes an abrupt sound but does not enhance the character’s emotional state effectively.
+Questum outperforms Werewolf, ThinkThrice, and O-CoT across all evaluated dimensions.
+Its precise and guided questioning prompts comprehensive and relevant answers, thereby advancing the story effectively.
+Additionally, the detailed responses enhance character immersion, making the narrative more compelling and believable.
+The comparison underscores the importance of strategic dialogue design in storytelling, where targeted interactions can significantly enrich the narrative experience.